### v1

# Tradeoffs among Action Taking Policies Matter in Active Sequential Multi-Hypothesis Testing: the Optimal Error Exponent Region

Abstract
Reliability of sequential hypothesis testing can be greatly improved when decision maker is given the freedom to adaptively take an action that determines the distribution of the current collected sample. Such advantage of sampling adaptivity has been realized since Chernoff’s seminal paper in 1959 [1]. While a large body of works have explored and investigated the gain of adaptivity, in the general multiple-hypothesis setting, the fundamental limits of individual error probabilities have not been fully understood. In particular, in the asymptotic regime as the expected stopping time tends to infinity, the error exponents are only characterized in specific cases, such as that of the total error probability. In this paper, we consider a general setup of active sequential multiple-hypothesis testing where at each time slot, a temporally varying subset of data sources (out of a known set) emerges from which the decision maker can select to collect samples, subject to a family of expected selection budget constraints. The selection of sources, understood as the “action” at each time slot, is constrained in a predefined action space. At the end of each time slot, the decision maker either decides to make the inference on the $M$ hypotheses, or continues to observe the data sources for the next time slot. The optimal tradeoffs among $M(M-1)$ types of error exponents are characterized, and the achievable region is shown to be a convex polytope. A companion asymptotically optimal test that strikes the balance between exploration and exploitation is proposed to achieve any target error exponents within the region. To the best of our knowledge, this is the first time in the literature to identify such tradeoffs among error exponents, and it uncovers the tension among different action taking policies even in the basic setting of Chernoff [1].
I Introduction
Adaptivity in taking samples from different distributions is known to greatly enhance the performance of sequential statistical inference. The intuition is that different distributions have different levels of informativeness of the inference target, and the most informative one may vary with the ground truth. The decision maker builds up more confidence towards the ground truth as it collects more samples, so it should choose the most informative distribution at the moment accordingly. As for hypothesis testing, the paradigm is called active sequential hypothesis testing [1, 2, 3], where the decision maker has the freedom to adaptively take an action that determines the distribution of the sample on the fly. It can be understood as the setting where the decision maker selects a particular data source out of multiple ones to collect a sample, each of which corresponds to a possible action. This paradigm finds extensive application in various fields such as spectrum sensing, crowdsourced data analytics, etc., and motivated from these applications, there have been extensions that incorporate a variety of expected selection budge constraints [4, 5, 6, 7] and temporally changing available data sources [8].
In existing works about active sequential multi-hypothesis testing, however, the fundamental limits of individual error probabilities have not been fully understood. In the related literature [1, 2, 3, 4, 6, 5, 7], the main focus is on minimizing various costs related to the expected stopping time subject to constraints such as a constant total error probability [2], a constant maximal error probability [1, 7], constant type-I and type-II error probabilities [5, 6, 4], and constant decision risks [3]. When the treatment pertains to the general non-asymptotic regime, it is usually difficult to characterize the optimal performance analytically, and hence the goal becomes developing efficient algorithms for finding the optimal selection policy numerically under a given framework. As for the asymptotic regime, one may turn the problems in [1, 2, 3, 4, 6, 5, 7] around and study the optimal error probabilities subject to expected stopping time and budget constraints. It is then not difficult to characterize the exponentially decaying rates (error exponents) in the respective settings. However, in the general $M$-ary multi-hypothesis setting, there are in total $M(M-1)$ types of error probabilities, and it remains unclear how the asymptotically optimal policies and inference strategies change when trading off one error probability with others.
Towards finer resolution of the fundamental performance limits in active sequential $M$-ary multi-hypothesis testing problems, in this work we focus on $M(M-1)$ types of individual error probabilities in a general framework that covers those in the literature [1, 2, 3, 4, 6, 5, 7, 8] as special cases. In this framework, out of $n$ data sources in total, at each time slot a random subset of them is revealed to the decision maker as the available sources. The distribution of this random subset is i.i.d. over time. The decision maker then takes an action (may be randomized) to select data sources from this available subset and collects samples from the selected ones. At the end of each time slot, according to the samples accumulated so far, the decision maker either determines to make an inference among the $M$ hypotheses or advances to the next time slot. Under a family of expected selection budget constraints including the expected stopping time, the goal is to study the asymptote of individual error probabilities, in particular the exponentially decaying rates (error exponents), as all budget constraints scale to infinity with fixed ratios.
Our main contribution is the characterization of the optimal error exponent region. We show that the achievable error exponents can be represented as linear combinations of Kullback-Leibler (KL) divergences, each of which corresponds to a particular set of data sources’ joint level of informativeness for distinguishing one hypothesis from another. The coefficients in the linear combinations correspond to the expected selection frequencies of data sources under the given policy and turn out to be the key design choices. In our characterization, these selection frequencies are indexed naturally by the taken action and the available subset of data sources. The optimal error exponent region is the union of these $M(M-1)$-tuples of error exponents over all possible such selection frequencies that meet the normalized expected budget constraints (i.e., the ratios to the expected stopping time) and the distribution of the random available subset. Leveraging the fact that these constraints are all affine, we further show that the optimal error exponent region is a convex polytope if there are just a finite number of expected selection budget constraints. The characterized region turns out to be a direct product of $M$ sub-regions each of which is in an $(M-1)$-dimensional space, and a tradeoff only appears among the exponents that correspond to the same declared hypothesis against all other $M-1$ possible ground truth hypotheses. To the best of our knowledge, such tradeoffs among $M(M-1)$ types of error exponents have not been identified before, and it uncovers the tension among different action taking policies even in the basic setting of Chernoff [1].
To characterize the optimal error exponent region, we divide the proof of our main theorem into two parts: the converse part and the achievability part. The converse part of our main theorem is proved by standard arguments that involves data processing inequality and Doob’s Optional Stopping Theorem, similar to that in [8]. The more interesting part is achievability. The stopping and inference rules of our test mainly follow a family of multi-hypothesis sequential probability ratio tests (MSPRT) proposed in [9], with thresholds designed judiciously to achieve the target error exponents and meet the expected stopping time constraint. For the action taking policy, since the constraints are all in expectation, it is natural to use a randomized one, and hence the design boils down to choosing the action taking probabilites. To achieve a given tuple of error exponents in the region, the coefficients in the linear combinations of the KL divergences serve as the guidance for setting these action taking probabilities in the randomized policy. As suggested in the literature [1, 2, 3, 8], one way is to exploit the accumulated information and choose the action taking probabilities derived from the linear combinations that correspond to the hypothesis with the maximum likelihood. It turns out that similar to [1], such a “pure-exploitation” approach requires stronger assumptions on the data sources so that it can be shown to be asymptotically optimal. Inspired by [7], in our proposed policy, on top of the “pure-exploitation” part, we further mix it with a carefully designed probability for “exploration” so that more relaxed assumptions can be set.
The conference version of this work [10] considered a restricted setup where there are just two simple expected budget constraints: individual and total budget constraints, and the the collected samples at each time slot are assumed to be conditionally independent random variables given the current action. In contrast, in this work, we consider a family of expected selection budget constraints, namely linear expected budget constraints in the remaining context, and relax the conditional independence assumption.
The remainder of this paper is organized as follows. In Section II, we give the problem formulation, including the statistical model of collected samples and available subset of data sources, the family of active sequential tests, and the expected budget constraints. Our main theorem is presented in Section III along with an illustrative example scenario in the basic setting of Chernoff [1], demonstrating the tradeoffs among error exponents. The proposed asymptotically optimal test along with the achievability proof of our main theorem are given in Section IV. The details of the converse proof are provided in Section V. Section VI contains the proof that the optimal error exponent region is a convex polytope. Finally, in Section VII, we conclude this paper.
II Problem Formulation
There are $M$ hypotheses and $n$ data sources labeled by integers in $\left[M\right]\triangleq\{0,1,2,...,M-1\}$ and $\left[n\right]\triangleq\{1,...,n\}$, respectively, where $2\leq M<\infty$, $1\leq n<\infty$. If not specified, in this paper upper and lower case letters are used to denote random variables and the values of random variables, respectively.
II-A Statistical model
The statistical model of the general framework considered in this work comprises two parts: the available subset of data sources and the samples drawn from the data sources.
At each time instance $t\geq 1$, the temporally available subset of data source $Z_{t}$ (a random variable), following a probability mass function (pmf) $P_{Z}$ (known to the decision maker) in an i.i.d. manner over time, is revealed to the decision maker. Let $\mathcal{Z}=\{z_{1},z_{2},...,z_{\lvert\mathcal{Z}\rvert}\}$ denote the support of $P_{Z}$, $\lvert\mathcal{Z}\rvert<\infty$. Each $z\in\mathcal{Z}$ consists of a subset of data sources, i.e., $z_{i}\subseteq\left[n\right]\ \forall\,i\in\{1,2,...,\lvert\mathcal{Z}\rvert\}$.
As for the collected samples, at each time slot, all samples drawn from selected data sources jointly follow a corresponding joint distribution. Specifically, at each time slot $t\geq 1$, a sample drawn from a selected data source $j\in\left[n\right]$ is denoted as $X_{t,j}$ while all samples drawn from selected data sources are collected as a set $\mathbf{X}_{t}\triangleq\{X_{t,j}\,|\,j\in A_{t}\cap Z_{t}\}$, where $A_{t}$ is the taken action at time slot $t$ formulated as a subset of data sources in Section II-B. Furthermore, at each time slot $t\geq 1$, the joint distribution of samples in the set $\mathbf{X}_{t}$ is denoted as $P^{A_{t},Z_{t}}_{\theta}$. For notational convenience, let us denote $\mathbb{P}_{\theta}\{\cdot\}$ and $\mathbb{E}_{\theta}[\cdot]$ as the short-hand notations for “the probability” and “the expectation” under hypothesis $\theta\in\left[M\right]$.
II-B Action taking, stopping time, and inference
At each time instance $t\geq 1$, after the available subset $Z_{t}$ is revealed, the decision maker first takes an action denoted as $A_{t}$ drawn from an action space $\mathcal{A}\triangleq\{a_{1},a_{2},...,a_{\lvert\mathcal{A}\rvert}\}$, where $\lvert\mathcal{A}\rvert<\infty$ and each $a\in\mathcal{A}$ is a subset of data sources, i.e., $a_{i}\subseteq\left[n\right]\ \forall\,i\in\{1,2,...,\lvert\mathcal{A}\rvert\}$. Furthermore, the decision maker is allowed to skip any action, i.e., $\emptyset\in\mathcal{A}$, indicating the option to not select any sources at a time slot. The action taking process can be randomized, that is, $A_{t}\in\mathcal{A}$ is a random variable, following a distribution determined by the decision maker. This distribution is the action taking policy and is adaptive in general, that is, at each time instance $t\geq 1$, the probability of taking $A_{t}\in\mathcal{A}$ is a function of the data (comprising samples, actions, and available sets) observed until time slot $t-1$.
After taking an action at each time slot $t$, the decision maker collects a set of samples generated by the data sources that are selected, i.e., in the intersection of the action $A_{t}$ and the temporary available set $Z_{t}$. In short, $\forall\,t\geq 1$, the collected samples are in the set $\mathbf{X}_{t}=\{X_{t,j}\,|\,j\in A_{t}\cap Z_{t}\}$ where
$\mathbf{X}_{t}\sim P^{A_{t},Z_{t}}_{\theta}\quad\forall\,t\geq 1$
under hypothesis $\theta\in\left[M\right]$. We assume given the actions, the random sample set $\mathbf{X}_{t}$ is conditionally independent across time slots. Moreover, we use $\mathcal{X}^{a,z}_{\theta}$ to denote the support of the joint distribution $P^{a,z}_{\theta}$ $\forall\,a\in\mathcal{A},\ \forall\,z\in\mathcal{Z},\ \forall\,\theta\in\left[M\right]$. After receiving the set $\mathbf{X}_{t}$ at time slot $t\geq 1$, the decision maker either stops and makes an inference or proceeds to the next time slot. The time that the decision maker stops is the stopping time, denoted as $\tau\in\mathbb{N}$, and the inferred result is denoted as $\phi\in\left[M\right]$.
The procedures at the decision maker is summarized in the definition below.
Definition 1 (Active sequential test)
An active sequential test in the proposed framework comprises the following parts: at each time instance $t\geq 1$,
  • The distribution that $A_{t}$ follows is a function of $F_{t-1}$ and $Z_{t}$, where
    $F_{t}\triangleq\{\{\mathbf{X}_{s}\}_{s=1}^{t},\{A_{s}\}_{s=1}^{t},\{Z_{s}\}_{s=1}^{t}\}$
    $\forall\,t\geq 1$. By convention, $F_{0}=\emptyset$.
  • $\tau$ and $\phi$ are both functions of $F_{t}$, $\forall\,t\geq 1$.
As for the performance, this work focuses on the individual error probability: $\forall\,\theta\in\left[M\right]\setminus\{m\},\forall\,m\in\left[M\right]$,
$\pi_{m|\theta}\triangleq\mathbb{P}_{\theta}\{\phi=m\}.$ (1)
II-C Constraints and error exponents
For describing the budget constraints, we first define a set $\mathcal{C}_{B}$:
$\mathcal{C}_{B}\triangleq\left\{\underline{x}\triangleq\left\{x_{1},x_{2},...,x_{n}\right\}\in\mathbb{R}^{n}_{+}\,\middle|\,b_{i}(\underline{x})\leq U_{i}\ \forall\,i\in\mathcal{I}\right\},$ (2)
where $\mathbb{R}_{+}$ denotes the set of non-negative reals, $\mathcal{I}$ is an index set, each $b_{i}(\cdot)$ is a linear function satisfying $b_{i}(0)=0$ and $b_{i}(\underline{x})\geq 0\ \forall\,\underline{x}\in\mathbb{R}^{n}_{+}\ \forall\,i\in\mathcal{I}$, and each $U_{i}$ is namely a budget satisfying $U_{i}>0\ \forall\,i\in\mathcal{I}$. Note that $\mathcal{I}$ is allowed to be countable or uncountable and it is given to the decision maker along with $b_{i}(\cdot)$ and $U_{i}\ \forall\,i\in\mathcal{I}$.
Then, the budget constraints including expected stopping time constraints are considered as follows: For some $T\geq 1$ and a given set $\mathcal{C}_{B}$:
$\mathbb{E}_{\theta}\left[\tau\right]\leq T,\quad\underline{B}^{\theta}\triangleq\left\{B^{\theta}_{1},B^{\theta}_{2},...,B^{\theta}_{n}\right\}\in\mathcal{C}_{B}\quad\forall\,\theta\in\left[M\right],$ (3)
where $\mathcal{C}_{B}$ is defined in (2), each $B^{\theta}_{j}\triangleq\mathbb{E}_{\theta}\left[\sum_{t=1}^{\tau}\mathbbm{1}{\left\{j\in A_{t}\cap Z_{t}\right\}}\right]$ denotes the expected total number of selecting source $j\in\left[n\right]$ under hypothesis $\theta\in\left[M\right]$. Hence, the constraints (3) are also called linear expected budget constraints including expected stopping time constraints. Note that the constraints (3) should be satisfied for all hypothesis $\theta\in\left[M\right]$.
Our focus is on the fundamental limits of the performance in an asymptotic regime as $U_{i}\ \forall\,i\in\mathcal{I}$ and $T$ all tend to infinity with fixed ratios
$r_{i}\triangleq\frac{U_{i}}{T}\quad\forall\,i\in\mathcal{I},$
where $r_{i}>0\ \forall\,i\in\mathcal{I}$.
An illustrative example of the budget constraints defined in (3) is provided in Example 1, showing the budget constraints considered in the preliminary version of this work [10].
Example 1 (Individual and total budget constraints)
The budget constraints defined in [10] are individual and total budget constraints, a special case of $\mathcal{C}_{B}$ by identifying the following equivalences: =$\mathcal{I}=\left\{0,1,...,n\right\}$, for the individual budget constraints, $U_{i}=B_{i}\ \forall\,i\in\mathcal{I}\setminus\{0\}$, $b_{i}(\underline{x})=x_{i}\ \forall\,i\in\mathcal{I}\setminus\{0\}$, and for the total budget constraints, $U_{0}=B$, $b_{0}(\underline{x})=\sum_{j=1}^{n}x_{j}$, where $B_{i},B$ are defined in [10] as the individual budget for each data source $i\in\left[n\right]$ and the total budget, respectively. As a result, the constraints (3) turn to the individual and total budget constraints including expected stopping time considered in [10, Equation (3)]. For the asymptotic analysis, [10] also regarded $\frac{B_{i}}{T}\ \forall\,i\in\left[n\right]$ and $\frac{B}{T}$ as fixed ratios.
The achievable individual error exponent $e_{m|\theta},\forall\,\theta\in\left[M\right]\setminus\{m\},\forall\,m\in\left[M\right]$ is defined as the follows.
Definition 2 (Error exponents)
A given $M(M-1)$-tuple
$\Big{\{}e_{m|\theta}\,\Big{|}\,m\in\left[M\right],\theta\in\left[M\right]\setminus\{m\}\Big{\}}$ (4)
is achievable if there exists an active sequential test (comprising action taking, stopping time, and inference) satisfying the constraints (3) such that each of $M(M-1)$ individual error probabilities vanishes and satisfies
$\liminf_{T\to\infty}\left\{\!\frac{1}{T}\log\frac{1}{\pi_{m|\theta}}\!\right\}\geq e_{m|\theta}\ \forall\,\theta\in\left[M\right]\setminus\{m\},\forall\,m\in\left[M\right].$ (5)
Our goal is to characterize the optimal error exponent region, the collection of all achievable tuples.
III Main Results
In this section, we first establish the optimal region of the achievable individual error exponents, showing that the region is a convex polytope. Then we specialize our result under Chernoff’s setting [1, 2, 3] to clearly demonstrate that there may exist tradeoffs among the individual error exponents. Finally, we show that there is no tradeoff in the asymptotic regime when it comes to the performance metrics considered in prior works [1, 2, 3], such as the decision risks.
Notations
×$\bigtimes_{m=0}^{M-1}A_{m}$ denotes the direct product over sets, i.e., $A_{0}\times A_{2}\times...\times A_{M-1}$. $\text{Conv}(A)$ and $\text{extr}(A)$ denote as the convex hull of set $A$ and the collection of extreme points of set $A$, respectively. $\mathbb{R}_{+}$ denotes the set of non-negative reals. When the context is clear, we drop the indices and use $\{e_{m|\theta}\}$ to denote the $M(M-1)$-tuple in (4). $\{x_{j}\}_{j=1}^{k}$ denotes the $k$-tuple $\{x_{1},x_{2},...,x_{k}\}$ where $x_{j}$ is any value indexed by $j\in\{1,2,...,k\}$ for any integer $k\geq 2$.
III-A Main theorem
For establishing the characterization of the error exponents, we need to impose the following two assumptions.
Assumption 1
There exists $L<\infty$ such that
$\forall\,a,z\in\mathcal{A}\times\mathcal{Z},\,\sup_{m\in\left[M\right]}\sup_{\theta\in\left[M\right]\setminus\{m\}}\sup_{\mathbf{x}\in\mathcal{X}^{a,z}_{\theta}}\left\lvert\log\frac{P^{a,z}_{\theta}(\mathbf{x})}{P^{a,z}_{m}(\mathbf{x})}\right\rvert\leq L.$ (6)
where if $a=\emptyset$ (or $\log\frac{0}{0}$), the corresponding log-likelihood ratio is set to zero by convention.
Assumption 2
For any two hypotheses $m,\theta\in\left[M\right]$, $\theta\neq m$, $\exists z\in\mathcal{Z}$, $\exists a\in\mathcal{A}$, such that ${D}\left(P^{a,z}_{m}\,\middle\|P^{a,z}_{\theta}\right)>0$.
Assumption 1 is used in our achievability and converse proof and commonly made in the literature [2, 1, 3, 7, 8]. Furthermore, Assumption 2 is needed for proving the achievability, and further discussion will be made in Section IV. Roughly speaking, it implies that for any two different hypotheses, there exists a feasible set of data sources to be selected at some time slots, ensuring the possibility of distinguishing between any two different hypotheses.
We are ready to state our main result.
Theorem 1 (The optimal error exponent region)
Suppose both Assumption 1 and 2 hold. Let
$\bm{\beta}^{m}\triangleq\Big{\{}\beta^{m}_{a,z}\,\Big{|}\,a\in\mathcal{A},z\in\mathcal{Z}\Big{\}}\in\mathbb{R}^{|\mathcal{A}||\mathcal{Z}|}_{+}$
denote a tuple of coefficients, indexed by the hypothesis $m$ and the action $a$ and the available subset $z$, and $\mathcal{C}$ be the set of all $\bm{\beta}$’s (here we remove the index $m$ for notational convenience) satisfying the constraints below:
$b_{i}\left(\left\{\sum_{a\in\mathcal{A}}\sum_{z\in\mathcal{Z}}\beta_{a,z}\mathbbm{1}{\left\{j\in a\cap z\right\}}\right\}_{j=1}^{n}\right)\leq r_{i}\quad\forall\,i\in\mathcal{I}$ (7)
$\sum_{a\in\mathcal{A}}\beta_{a,z}\leq\alpha_{z}\triangleq P_{Z}(z)\quad\forall\,z\in\mathcal{Z}$ (8)
Define
$\zeta\triangleq\bigtimes\limits_{m=0}^{M-1}\,\left\{\bigcup\limits_{\bm{\beta}^{m}\in\mathcal{C}}\zeta_{m}(\bm{\beta}^{m})\right\}$ (9)
where for each $m\in\left[M\right]$, $\zeta_{m}(\bm{\beta}^{m})$ denotes the collection of $(M-1)$-tuples $\big{\{}e_{m|\theta}\,\big{|}\,\theta\in\left[M\right]\setminus\{m\}\big{\}}$ satisfying
$e_{m|\theta}\leq e^{*}_{m|\theta}(\bm{\beta}^{m})\triangleq\sum_{a\in\mathcal{A}}\sum_{z\in\mathcal{Z}}\beta^{m}_{a,z}{D}\left(P^{a,z}_{m}\,\middle\|P^{a,z}_{\theta}\right)$
for every $\theta\in\left[M\right]\setminus\{m\}$. Then, any $M(M-1)$-tuple $\{e_{m|\theta}\}$ inside $\zeta$ can be achieved by our proposed test with a corresponding adaptive selection policy. Conversely, an achievable $M(M-1)$-tuple $\{e_{m|\theta}\}$ must lie inside the region $\zeta$.
Proof:
The proof of the converse part of Theorem 1 primarily relies on the data processing inequality and Doob’s Optional Stopping Theorem, similar to that in [8]. Details can be found in Section V. Proof of the achieva The proposed test and the achievability proof are provided in Section IV. ∎
Before we proceed, let us make three remarks based on the characterization in our main theorem.
Remark 1 (Tradeoffs among error exponents)
From the characterization of the error exponent region as a direct product of $M$ components as shown in (9), it can be readily seen that the tradeoff only appears among the $(M-1)$-tuples in the component $\bigcup_{\bm{\beta}^{m}\in\mathcal{C}}\zeta_{m}(\bm{\beta}^{m})$ for each $m\in\left[M\right]$, that is, in each region $\zeta_{m}(\bm{\beta}^{m})$, one may not find an optimal $\bm{\beta}^{m}\in\mathcal{C}$ to maximize all bounds $\big{\{}e^{*}_{m|\theta}(\bm{\beta}^{m})\,\big{|}\,\theta\in\left[M\right]\setminus\{m\}\big{\}}$ simultaneously. Later in Section III-B, the tradeoff is illustrated in a simple yet classical setting originally considered by Chernoff [1].
Remark 2 (No tradeoff when $M=2$)
When there are only two hypotheses ($M=2$), each $\zeta_{m}(\bm{\beta}^{m})$ becomes a segment, i.e. $\zeta_{0}(\bm{\beta}^{0})=\left\{e_{0|1}\,\big{|}\,e_{0|1}\leq e^{*}_{0|1}(\bm{\beta}^{0})\right\}$ and $\zeta_{1}(\bm{\beta}^{1})=\left\{e_{1|0}\,\big{|}\,e_{1|0}\leq e^{*}_{1|0}(\bm{\beta}^{1})\right\}$, and it is clear that there is no tradeoff at all.
Remark 3 (Specialization for [10])
In the conference version [10], a special case of the problem formulation was considered. Specifically, [10] first introduced the expected individual and total budget constraints (see Example 1). Furthermore, [10] assumed at each time slot $t\geq 1$, given the current action, the collected samples in the set $\mathbf{X}_{t}$ are conditionally independent of one another, that is, $\forall\,t\geq 1$, if $A_{t}=a\in\mathcal{A},Z_{t}=z\in\mathcal{Z}$, the joint distribution $P^{a,z}_{\theta}=\prod_{j\in a\cap z}P_{\theta,j}$, where $P_{\theta,j}$ is the distribution that the sample drawn from the selected data source $j$ follows under hypothesis $\theta$, leading to ${D}\left(P^{a,z}_{m}\,\middle\|P^{a,z}_{\theta}\right)=\sum_{j\in a\cap z}{D}\left(P_{m,j}\,\middle\|P_{\theta,j}\right)\ \forall\,a\in\mathcal{A},\ z\in\mathcal{Z},\ \theta\in\left[M\right]$. Through this specialization (consists of the individual and total budget constraints and the conditional independence assumption), Theorem 1 easily turns to [10, Theorem 1] following the proof steps similar to that of Theorem 1 with this specialization.
From the form of the region $\zeta$ in Theorem 1, one cannot clearly understand the structure of the region $\zeta$ due to the different realizations of the set $\mathcal{C}$. Nevertheless, under a non-trivial scenario, we provide a more structural understanding of the region $\zeta$. In numerous applications of active sequential hypothesis testing such as sensor networks considered in [10], there are only finite numbers of linear budget constraints, e.g., individual and total budgets constraints [10], that is, $\lvert\mathcal{I}\rvert<\infty$. In this case, we can show that the region $\zeta$ is a convex polytope.
Proposition 1 (Convex polyhedral region)
The error exponent region in (9) can be equivalently written as
$\zeta=\bigtimes\limits_{m=0}^{M-1}\,\text{Conv}\left\{\text{extr}\bigl{(}\bigcup_{\bm{\beta}^{m}\in V_{c}}\zeta_{m}(\bm{\beta}^{m})\bigr{)}\right\}$
where $V_{c}$ denotes the collection of all extreme points of the constraint set $\mathcal{C}$ with $\lvert\mathcal{I}\rvert<\infty$.
Proof:
The proof relies on the facts that the set $\mathcal{C}$ with $\lvert\mathcal{I}\rvert<\infty$ is a convex polytope and, for each $m\in\left[M\right]$, the region $\bigcup_{\bm{\beta}^{m}\in\mathcal{C}}\zeta_{m}(\bm{\beta}^{m})$ is a compact convex set. The detailed proof is provided in Section VI (the definition of the extreme points is also given in Section VI). ∎
III-B Discussion
To highlight the tradeoffs among individual error exponents, we specialize Theorem 1 under Chernoff’s setting [1, 2, 3], where all data sources are available at each time instance ($\mathcal{Z}=\left\{\left[n\right]\right\}$), but only one data source can be chosen per time ($\mathcal{A}=\left\{\{1\},\{2\},...\{n\}\right\}$) without the budget constraints ($\mathcal{C}_{B}=\mathbb{R}^{n}_{+}$). The following straightforward corollary summarizes the error exponent region.
Corollary 1 (The optimal error exponent region in Chernoff’s setting)
The optimal error exponent region for individual error probabilities in Chernoff’s setting [1] is given by
$\zeta=\bigtimes\limits_{m=0}^{M-1}\,\left\{\bigcup\limits_{\bm{\beta}^{m}\in\mathcal{P}^{n}}\zeta_{m}(\bm{\beta}^{m})\right\}$ (10)
where $\mathcal{P}^{n}$ is the $n$-dimensional probability simplex, and for each $m\in\left[M\right]$, $\zeta_{m}(\bm{\beta}^{m})$ is defined as in Theorem 1 with the following specialization:
$e^{*}_{m|\theta}(\bm{\beta}^{m})=\sum_{j=1}^{n}\beta^{m}_{j}{D}\left(P_{m,j}\,\middle\|P_{\theta,j}\right)\quad\forall\,\theta\in\left[M\right]\setminus\{m\}.$
For illustration of the tradeoffs, let us further consider a special case when there are just three hypotheses and two data sources in Corollary 1. In this case, for any $\theta,m\in[3]$, $\theta\neq m$, the bound $e^{*}_{m|\theta}(\bm{\beta}^{m})$ is simplified as
$e^{*}_{m|\theta}(\bm{\beta}^{m})=\beta^{m}_{1}{D}\left(P_{m,1}\,\middle\|P_{\theta,1}\right)+\beta^{m}_{2}{D}\left(P_{m,2}\,\middle\|P_{\theta,2}\right).$
To see the tradeoffs, we may focus on a particular $m$, say $m=0$, since tradeoffs only happen within each component in the direct product in (9). Subsequently, let us investigate the tradeoff between $e^{*}_{0|1}(\bm{\beta}^{0})$ and $e^{*}_{0|2}(\bm{\beta}^{0})$, where
$e^{*}_{0|1}(\bm{\beta}^{0})=\beta^{0}_{1}{D}\left(P_{0,1}\,\middle\|P_{1,1}\right)+\beta^{0}_{2}{D}\left(P_{0,2}\,\middle\|P_{1,2}\right),$
$e^{*}_{0|2}(\bm{\beta}^{0})=\beta^{0}_{1}{D}\left(P_{0,1}\,\middle\|P_{2,1}\right)+\beta^{0}_{2}{D}\left(P_{0,2}\,\middle\|P_{2,2}\right).$
With $\bm{\beta}^{0}\in\mathcal{P}^{2}$, i.e., $\beta^{0}_{1}+\beta^{0}_{2}=1$ and $\beta^{0}_{1},\beta^{0}_{2}\geq 0$, the tradeoff emerges when each source is the most informative in distinguishing hypothesis 00 from exactly one hypothesis, e.g.,
${D}\left(P_{0,1}\,\middle\|P_{1,1}\right)>{D}\left(P_{0,2}\,\middle\|P_{1,2}\right)$
but ${D}\left(P_{0,1}\,\middle\|P_{2,1}\right)<{D}\left(P_{0,2}\,\middle\|P_{2,2}\right)$
As a result, $e^{*}_{0|1}(\bm{\beta}^{0})$ and $e^{*}_{0|2}(\bm{\beta}^{0})$ cannot be maximized simultaneously, resulting the tradeoff. Conversely, if one of the sources (say source $1$) is the most informative in distinguishing hypothesis 00 from both $1$ and $2$, one can set $\beta^{0}_{1}=1$ to maximize $e^{*}_{0|1}(\bm{\beta}^{0})$ and $e^{*}_{0|2}(\bm{\beta}^{0})$ simultaneously.
From the above illustrative example, it can be seen that in Chernoff’s setting, the tradeoffs among the individual error exponents depend on the heterogeneous level of informativeness across data sources in distinguishing a declared hypothesis again other ground truth hypotheses. This insight also carries to the general case considered in our main result.
As for why the tradeoff only appears when $M\geq 3$, from a high-level viewpoint, for each $m\in\left[M\right]$, $\bm{\beta}^{m}$ for $\zeta_{m}(\bm{\beta}^{m})$ can be interpreted as a “guidance of selection” for distinguishing hypotheses, and when $M\geq 3$, $\bm{\beta}^{m}$ is responsible for distinguishing hypothesis $m$ from more than just one hypotheses (i.e., $\lvert\left[M\right]\setminus\{m\}\rvert\geq 2$), making it challenging to find a guidance simultaneously optimal under the heterogeneous distinguishability among data sources. However, when $M=2$, each $\bm{\beta}^{m}$ for $\zeta_{m}(\bm{\beta}^{m})$ just needs to take care of distinguishing one hypothesis from the other, and hence there is no tension in picking the guidance of selection.
III-C Performance metrics without tradeoff
In the literature of active sequential hypothesis testing, such tradeoffs in the asymptotic regime have not been observed. The reason is that prior studies have not delved into individual error probabilities for $M\geq 3$. For example, [1, 3, 2] focused on a coarser performance metric, the decision risk of declaring the decision as hypothesis $m$
$\mathrm{R}_{m}\triangleq\sum_{\theta\in\left[M\right]\setminus\{m\}}\rho_{\theta}\pi_{m|\theta},\quad m\in\left[M\right].$
Here $\rho_{\theta}$ is the prior probability for hypothesis $\theta\in\left[M\right]$.
Fundamental limits of the exponentially decaying rates of the decision risks, denoted as $\{\gamma_{m}\mid m\in\left[M\right]\}$, can then be characterized from our main result as a simple corollary.
Corollary 2 (Exponents of decision risks)
The optimal exponent region consists of non-negative $(\gamma_{1},...,\gamma_{M})$ satisfying
$\gamma_{m}\leq\max_{\bm{\beta}\in\mathcal{P}^{n}}\min_{\theta\in\left[M\right]\setminus\{m\}}e^{*}_{m|\theta}(\bm{\beta})\quad\forall\,m\in\left[M\right].$
Note that Corollary 2 can also be derived from the respective results in [1, 3, 2]. Clearly there are no tradeoffs among these $\gamma_{m}$’s, $m\in\left[M\right]$ due to the product form shown in (10).
Remark 4
Suppose we take the error probabilities of individual hypotheses
$\pi_{\theta}\triangleq\sum_{m\in\left[M\right]\setminus\{\theta\}}\pi_{m|\theta},\quad\theta\in\left[M\right]$
as performance metrics. In contrast to the decision risks, the tradeoff among the corresponding error exponents exists. The asymptote of error probabilities of individual hypotheses, however, was not studied in existing works [1, 3, 2].
IV Proof of Achievability
IV-A Overview
For the achievability part of Theorem 1, from the expression in (9), the goal is to show that for any $\bm{\beta}^{0},\bm{\beta}^{1},\ldots,\bm{\beta}^{M-1}\in\mathcal{C}$, the tuple of $M(M-1)$ individual error exponents
$\Big{\{}e_{m|\theta}\,\Big{|}\,m\in\left[M\right],\theta\in\left[M\right]\setminus\{m\}\Big{\}}$
satisfying
$\big{\{}e_{m|\theta}\,\big{|}\,\theta\in\left[M\right]\setminus\{m\}\big{\}}\in\zeta_{m}(\bm{\beta}^{m})\quad\forall\,m\in\left[M\right]$
is achievable. Recall that each $\bm{\beta}^{m}$, $m\in\left[M\right]$, comprises coefficients indexed by the action $a$ and the available subset $z$, which hints the probability of taking action $a$ when the available subset is $z$ under a randomized policy.
Note that $\bm{\beta}^{\theta}$ (here we change the indexing convention from $m$ to $\theta$ for notational convenience) corresponds to the action taking policy under hypothesis $\theta$, and they are in general different across all $\theta\in\left[M\right]$. One strategy that exploits the information accumulated so far is to first find a maximum likelihood estimate (MLE) $\hat{\theta}_{\text{MLE}}\in\left[M\right]$ and employ the policy corresponding to $\bm{\beta}^{\hat{\theta}_{\text{MLE}}}$ [1]. In our design, in addition to exploitation, we further mix it with a small (vanishing-in-$T$) probability of exploration inspired by [7], which is a common idea in the literature of sequential decision making problems such as bandits. This probability of exploration and the rate of uniformly exploring all sources are carefully chosen to meet the constraints and achieve the target error exponents.
As for the stopping rule and the inference strategy, it follows those in the multiple-hypothesis sequential probability ratio tests (MSPRT) [9] with thresholds designed judiciously to achieve the target error exponents and meet the constraints. A key statistic here is the cumulative log-likelihood ratio
$S_{t,\theta,m}\triangleq\log\frac{\mathsf{P}_{\theta}^{(t)}(F_{t})}{\mathsf{P}_{m}^{(t)}(F_{t})},$
where $\mathsf{P}_{\theta}^{(t)}(\cdot)$ represents the probability distribution of $F_{t}$, the accumulated information up to the end of time instance $t$ (see Definition 1), under hypothesis $\theta$. the explicit form of $S_{t,\theta,m}$ is provided in Lemma 2 in Appendix.
To this end, let us sketch the proposed test, with more details in Section IV. For a given $(\bm{\beta}^{0},\bm{\beta}^{1},\ldots,\bm{\beta}^{M-1})$, at each time instance $t\geq 1$, upon $Z_{t}$ being revealed, the decision maker carries out the following:
  • Action taking and sample collection: Find the MLE $\hat{\theta}_{\text{MLE}}$ at time slot $t-1$, that is, $\theta$ such that $\min_{m\neq\theta}S_{t-1,\theta,m}\geq 0$. Based on this $\hat{\theta}_{\text{MLE}}$ and the corresponding $\bm{\beta}^{\hat{\theta}_{\text{MLE}}}$ together with a judiciously designed exploration rate, we then set up a conditional pmf $P_{A_{t}|Z_{t},F_{t-1}}(\cdot|z,f_{t-1})$ for randomized action taking. How to design this conditional pmf is elaborated later. Subsequently, a random action $a$ is drawn from the conditional pmf, and the decision maker collects a set of samples $\mathbf{X}_{t}$ accordingly.
  • Stopping rule: If exist a hypothesis $\theta$ such that $S_{t,\theta,m}$ exceeds a set of designed thresholds $\forall\,m\neq\theta$, then stop.
  • Inference: Infer the hypothesis $\theta$ that causes stopping.
As for the conditional pmf used for action taking, if hypothesis $\theta$ has the largest cumulative log-likelihood ratio $(\hat{\theta}_{\text{MLE}}=\theta)$ at time slot $t-1$, then we take
$P_{A_{t}|Z_{t},F_{t-1}}(a|z,f_{t-1})=\begin{cases}(1-\epsilon(T))\frac{\beta^{\theta}_{a,z}}{\alpha_{z}}+\epsilon(T)\frac{1}{h\alpha_{z}},&a\in\mathcal{A}\setminus\{\emptyset\}\\[4.30554pt] 1-\sum_{a^{\prime}\in\mathcal{A}\setminus\{\emptyset\}}P_{A_{t}|Z_{t},F_{t-1}}(a^{\prime}|z,f_{t-1}),&a=\emptyset\end{cases}$
where $\epsilon(T)$ is the probability of conducting exploration, vanishing as $T\rightarrow\infty$. The decaying rate is judiciously designed to ensure conducting the exploration and retain the asymptotic optimality of the test. Additionally, the probability of set $z$ being available ($\alpha_{z}$) is in the denominator since it is a conditional probability. Furthermore, the exploration rate $\frac{1}{h}>0$ is chosen to be bounded from above as
$\frac{1}{h}\leq\min\left\{\frac{\min_{z\in\mathcal{Z}}\alpha_{z}}{\lvert\mathcal{A}\rvert},\ \inf_{i\in\mathcal{I}}\left\{\frac{r_{i}}{b_{i}\left(\left\{\sum_{a\in\mathcal{A}}\sum_{z\in\mathcal{Z}}\mathbbm{1}{\left\{j\in a\cap z\right\}}\right\}_{j=1}^{n}\right)}\right\}\right\}$
where each $b_{i}(\cdot)$ is defined in (2). The choice is to satisfy the expected budget constraints and that the total probability of action taking sums up to be $1$.
Before we move on to the details of the proposed test and the analysis, we make the following remark on the exploration part in the action taking policy.
Remark 5 (On the necessity of exploration)
If we consider “pure-exploitation”, the result presented in Theorem 1 remains the same. However, for the achievability, a stronger assumption than Assumption 2 is needed, namely, $\forall\,z\in\mathcal{Z}$, $\forall\,a\in\mathcal{A}\setminus\{\emptyset\}$,
${D}\left(P^{a,z}_{\theta}\,\middle\|P^{a,z}_{m}\right)>0\quad\forall\theta,m\in\left[M\right],\theta\neq m.$
This stronger assumption can be traced back to Chernoff [1]. Some prior works [2, 3] also provide modified tests (with a cold-start phase for example in [2]) to solely consider the weaker assumption (Assumption 2). Nevertheless, we fail to extend the analysis of these exploitation-based tests to establish the achievability part of the optimal error exponent region under Assumption 2.
IV-B Details of the proposed test
Notations
The summation $\sum_{a,z}$ without its index set, is a short-hand notation for $\sum_{a\in\mathcal{A}\setminus\{\emptyset\}}\sum_{z\in\mathcal{Z}}$. Note that we exclude the empty set because when $a=\emptyset$, ${D}\left(P^{a,z}_{\theta}\,\middle\|P^{a,z}_{m}\right)$ is defined as zero by convention for any $z\in\mathcal{Z}$, $\theta,m\in\left[M\right]$. The notations used in our proposed test are summarized in Table I.
Notations Descriptions
$F_{t}$$\left\{\left\{\mathbf{X}_{s}\right\}_{s=1}^{t},\left\{A_{s}\right\}_{s=1}^{t},\left\{Z_{s}\right\}_{s=1}^{t}\right\}$
$S_{t,\theta,m}$ $\log\frac{\mathsf{P}_{\theta}^{(t)}(F_{t})}{\mathsf{P}_{m}^{(t)}(F_{t})}$
$l$ $\frac{1}{T^{1/6}}\left(\frac{\min_{\theta\in\left[M\right]}\min_{m\neq\theta}A_{\theta,m}}{\min_{\theta\in\left[M\right]}\min_{m\neq\theta}\tilde{A}_{\theta,m}}\right)^{2}$
$\epsilon(T)$ $\left(\log(T)\right)^{-1/4}$
$\tilde{A}_{\theta,m}$ $\frac{1}{h}\sum_{a,z}{D}\left(P^{a,z}_{\theta}\,\middle\|P^{a,z}_{m}\right)$
$A_{\theta,m}$ $\frac{\left(1-\epsilon(T)\right)e^{*}_{\theta|m}\left(\bm{\beta}^{\theta}\right)+\epsilon(T)\tilde{A}_{\theta,m}}{1+l}$
$I(T)$ $\epsilon(T)\min_{\theta\in\left[M\right]}\min_{m\neq\theta}\tilde{A}_{\theta,m}$
$b(l)$ $\frac{l^{3}}{4(1+l)^{3}}\left(\frac{I(T)}{4L}\right)^{2}$
$B(l)$ $2+\frac{2(1+l)^{2}}{l^{2}}\left(\frac{4L}{I(T)}\right)^{2}$
$q(l,n,I(T))$ $1+\frac{1}{b(l)}\left\{1+\log\left(MB(l)\left(1+b(l)\right)\right)\right\}$
$A^{\text{max}}_{\theta}$ $\max_{m\neq\theta}A_{\theta,m}$
$\epsilon_{\theta}$ $A^{\text{max}}_{\theta}\left\{1-\frac{K\left(-MB(l)\left(1+b(l)\right)e^{-b(l)(T-1)}\right)}{b(l)}\right\}$
$K(x)$ $\sqrt{xe+1}-1$ with domain:$[-e^{-1},0]$; range:$[-1,0]$
TABLE I: Summary of notations in our proposed test
We consider two cases in our proposed test as follows:
  1. 1.
    If $T<\max\{e,q\left(l,n,I(T)\right)\}$, then stop at time instance $t=1$, do not select any sources, and randomly guess a hypothesis.
  2. 2.
    If $T\geq\max\{e,q\left(l,n,I(T)\right)\}$, then at each time instance $t\geq 1$, do the asymptotically optimal test.
where $e$ is Euler’s number ($e\cong 2.71828...$) and $q\left(l,n,I(T)\right)$ is defined in Table I, $n$ is the total number of data sources, $l$ and $I(T)$ are two specific sequences defined in Table I. Note that the error performance of the random guess in the first case does not dominate the asymptotic error performance (when $T\to\infty$) since by Lemma 1 in Appendix, $q\left(l,n,I(T)\right)$ is sublinear in $T$. Furthermore, the first case naturally satisfies the constraints (3). Therefore, we only consider the second case. Technically, our analysis of expected stopping time relies on $T\geq\max\{e,q\left(l,n,I(T)\right)\}$.
Before we start showing the achievability of our main theorem, we completely provide the procedures of the asymptotically optimal test (for the second case) described in the following: Given any $\bm{\beta}^{0},\bm{\beta}^{1},\ldots,\bm{\beta}^{M-1}\in\mathcal{C}$, at each time instance $t\geq 1$:
Firstly, the decision maker conduct the randomized action taking following the pmf $P_{A_{t}|Z_{t},F_{t-1}}(\cdot|z,f_{t-1})$ with the MLE $\hat{\theta}_{\text{MLE}}$ at time slot $t-1$, under a revealed $Z_{t}=z\in\mathcal{Z}$. Specifically, if $\hat{\theta}_{\text{MLE}}=\theta$ at time slot $t-1$,
$P_{A_{t}|Z_{t},F_{t-1}}(a|z,f_{t-1})=\begin{cases}(1-\epsilon(T))\frac{\beta^{\theta}_{a,z}}{\alpha_{z}}+\epsilon(T)\frac{1}{h\alpha_{z}},&a\in\mathcal{A}\setminus\{\emptyset\}\\[4.30554pt] 1-\sum_{a^{\prime}\in\mathcal{A}\setminus\{\emptyset\}}P_{A_{t}|Z_{t},F_{t-1}}(a^{\prime}|z,f_{t-1}),&a=\emptyset\end{cases}$
where $\frac{1}{h}$ is designed as follows:
$0<\frac{1}{h}\leq\min\left\{\frac{\min_{z\in\mathcal{Z}}\alpha_{z}}{\lvert\mathcal{A}\rvert},\ \inf_{i\in\mathcal{I}}\left\{\frac{r_{i}}{b_{i}\left(\left\{\sum_{a\in\mathcal{A}}\sum_{z\in\mathcal{Z}}\mathbbm{1}{\left\{j\in a\cap z\right\}}\right\}_{j=1}^{n}\right)}\right\}\right\}$
Then, the decision maker collects a set of samples drawn from the selected data sources, namely, $\mathbf{X}_{t}\triangleq\{X_{t,j}\,|\,j\in A_{t}\cap Z_{t}\}$.
Secondly, after collecting $\mathbf{X}_{t}$, the decision maker decides to either stop or proceed to the next time slot. The stopping time is defined as follows,
$\tau\triangleq\inf\left\{t\geq 1:\exists\theta\in\left[M\right],\min_{m\neq\theta}\left\{S_{t,\theta,m}-\left(TA_{\theta,m}-\epsilon_{\theta}\right)\right\}\geq 0\right\}$
Finally, if $\tau=t$, we claim $\phi$ is the hypothesis that causing the stopping.
Note that $F_{t}$, $S_{t,\theta,m}$, $A_{\theta,m}$, and $\epsilon_{\theta}$ are defined in Table I and the explicit form of $S_{t,\theta,m}$ is provided in Lemma 2 in Appendix.
IV-C Analysis of the proposed test
The proof of achievability involves three parts: 1) Demonstrating that our proposed test satisfies the expected stopping time constraints, where the proof steps are slightly modified the approaches proposed in the achievability proofs of [2, Proposition 2] and [7, Theorem 4], which originate from [1, Theorem 1]. 2) Showing that our proposed test satisfies the expected budget constraints, i.e., $\forall\,\theta\in\left[M\right]$, $\underline{B}^{\theta}\in\mathcal{C}_{B}$ where $\underline{B}^{\theta}\ \forall\,\theta\in\left[M\right]$ and $\mathcal{C}_{B}$ are defined in Section II-C. 3) Verifying that, with a given $\{\bm{\beta}^{i}\,|\,\bm{\beta}^{i}\in\mathcal{C},\,i\in\left[M\right]\}$, our proposed test achieves the tuples of $M(M-1)$ individual error exponents
$\Big{\{}e_{m|\theta}\,\Big{|}\,m\in\left[M\right],\theta\in\left[M\right]\setminus\{m\}\Big{\}}$
satisfying
$\big{\{}e_{m|\theta}\,\big{|}\,\theta\in\left[M\right]\setminus\{m\}\big{\}}\in\zeta_{m}(\bm{\beta}^{m})\quad\forall\,m\in\left[M\right].$
Note that the last two parts are similar to the achievability proof of [8, Theorem 1].
Throughout the proof of achievability, we are given a set of $\bm{\beta}^{0},\bm{\beta}^{1},\ldots,\bm{\beta}^{M-1}\in\mathcal{C}$.
IV-C1 Expected stopping time constraints
In this section, the goal is to show, for some $T\geq 1$:
$\mathbb{E}_{\theta}[\tau]\leq T\quad\forall\,\theta\in\left[M\right]$
Given that the above constraints are trivially satisfied in the first case of our proposed test, we solely illustrate the satisfaction of the above constraints in the second case, namely, the case with $T\geq\max\{e,q\left(l,n,I(T){}\right)\}$.
We first define a special stopping time $\tau_{\theta}$ for any $\theta\in\left[M\right]$ as follows:
$\tau_{\theta}\triangleq\inf\left\{t\,:\,\min_{m\neq\theta}\left\{S_{t,\theta,m}-TA_{\theta,m}+\epsilon_{\theta}\right\}\geq 0\right\}$
Therefore, by the definition of the stopping time $\tau$ in our proposed test, we have
$\tau=\min_{\theta\in\left[M\right]}\tau_{\theta}$
Hence, $\mathbb{E}_{\theta}\left[\tau\right]\leq\mathbb{E}_{\theta}\left[\tau_{\theta}\right],\forall\,\theta\in\left[M\right]$.
Then, the goal turns to show, in the second case of our proposed test,
$\mathbb{E}_{\theta}\left[\tau_{\theta}\right]\leq T\quad\forall\,\theta\in\left[M\right]$
To summary the proof steps, in the beginning, we decompose $\mathbb{E}_{\theta}\left[\tau_{\theta}\right]$ into two terms by tail sum theorem,
$\mathbb{E}_{\theta}\left[\tau_{\theta}\right]\leq\kappa_{\theta,T,l}+\sum_{k=\kappa_{\theta,T,l}}^{\infty}\mathbb{P}_{\theta}\left(\tau_{\theta}>k\right)\quad\forall\,\theta\in\left[M\right]$
where $\kappa_{\theta,T,l}\triangleq\lceil T-\frac{\epsilon_{\theta}}{A^{\text{max}}_{\theta}}\rceil$, $\forall\,\theta\in\left[M\right]$, $A^{\text{max}}_{\theta}\triangleq\max_{m\neq\theta}A_{\theta,m}$, and $\epsilon_{\theta}$ and $A_{\theta,m}$ are defined in Table I. Then, we show that the probability $\mathbb{P}_{\theta}\left(\tau_{\theta}>k\right)$ is exponential decay as $k\rightarrow\infty$ $\forall\,\theta\in\left[M\right]$ by Azuma’s inequality. Moreover, the upper bound for $\sum_{k=\kappa_{\theta,T,l}}^{\infty}\mathbb{P}_{\theta}\left(\tau_{\theta}>k\right)$ eliminates the term $\left(\frac{\epsilon_{\theta}}{A^{\text{max}}_{\theta}}+1\right)$ for any $T\geq\max\{e,q\left(l,n,I(T)\right)\}$, achieving the expected stopping time constraints for $\tau_{\theta}$, namely, the expected stopping time constraints for our stopping time $\tau$.
Hence, we follow the decomposition and provide the desired upper bound for $\mathbb{P}_{\theta}\left(\tau_{\theta}>k\right)$ $\forall\,k\geq\kappa_{\theta,T,l}$ $\forall\,\theta\in\left[M\right]$. To begin with, let us focus on one hypothesis as the ground truth hypothesis (say hypothesis $\theta$) without loss of generality and define the following two sets for any time slot $t\geq 1$ as in the proof of [2, Proposition 2]:
$\textbf{B}_{\theta}(t)\triangleq\left\{F_{t}\,:\,\min_{m\neq\theta}\left\{S_{t,\theta,m}-\left(TA_{\theta,m}-\epsilon_{\theta}\right)\right\}<0\right\}$
$\tilde{\textbf{B}}_{\theta}(t)\triangleq\left\{F_{t}\,:\,\min_{m\neq\theta}S_{t,\theta,m}<0\right\}$
where $F_{t}$ is defined in Table I. Additionally, we define $\tilde{\tau}_{\theta}$ as follows,
$\tilde{\tau}_{\theta}\triangleq\inf\left\{t\,:\,\min_{m\neq\theta}S_{t^{\prime},\theta,m}\geq 0\,,\forall\,t^{\prime}\geq t\right\}$
where $\tilde{\tau}_{\theta}$ means that for any time slot $t\geq\tilde{\tau}_{\theta}$, the maximum likelihood estimate $\hat{\theta}_{\text{MLE}}=\theta$.
Then, we derive the upper bound for $\mathbb{P}_{\theta}(\tau_{\theta}>k)$. By the definitions of stopping time $\tau_{\theta}$ and the region $\textbf{B}_{\theta}(k)$, $\forall\,k\geq\kappa_{\theta,T,l}$, we have
$\mathbb{P}_{\theta}\left(\tau_{\theta}>k\right)$ $\leq\mathbb{P}_{\theta}\left(\textbf{B}_{\theta}(k)\right)$
$=\mathbb{P}_{\theta}\left(\textbf{B}_{\theta}(k)\cap\left\{\tilde{\tau}_{\theta}\leq g_{k,l}\right\}\right)+\mathbb{P}_{\theta}\left(\textbf{B}_{\theta}(k)\cap\left\{\tilde{\tau}_{\theta}>g_{k,l}\right\}\right)$
$\leq\mathbb{P}_{\theta}\left(\textbf{B}_{\theta}(k)\cap\left\{\tilde{\tau}_{\theta}\leq g_{k,l}\right\}\right)+\mathbb{P}_{\theta}\left(\tilde{\tau}_{\theta}>g_{k,l}\right)$
$\leq\underbrace{\mathbb{P}_{\theta}\left(\textbf{B}_{\theta}(k)\cap\left\{\tilde{\tau}_{\theta}\leq g_{k,l}\right\}\right)}_{(i)}+\sum_{t^{\prime}=g_{k,l}}^{\infty}\underbrace{\mathbb{P}_{\theta}\left(\tilde{\textbf{B}}_{\theta}(t^{\prime})\right)}_{(ii)}$
where $g_{k,l}\triangleq k\left(\frac{l/2}{1+l}\right),\forall\,k\geq 1$. Note that $l$ is a specific sequence defined in Table I. Next, we derive upper bounds for the term $(i)$ and the term $(ii)$.
For the upper bound of the term $(i)$, we slightly modify the achievability proof proposed in [7, Theorem 4]. Specifically, by union bound and the definition of $\textbf{B}_{\theta}(k)$, we have
$(i)$ $=\mathbb{P}_{\theta}\left(\left\{\min_{m\neq\theta}\left\{S_{k,\theta,m}-\left(TA_{\theta,m}-\epsilon_{\theta}\right)\right\}<0\right\}\cap\left\{\tilde{\tau}_{\theta}\leq g_{k,l}\right\}\right)$
$\leq\sum_{m\neq\theta}\mathbb{P}_{\theta}\left(\left\{S_{k,\theta,m}<TA_{\theta,m}-\epsilon_{\theta}\right\}\cap\left\{\tilde{\tau}_{\theta}\leq g_{k,l}\right\}\right)$
$=\sum_{m\neq\theta}\mathbb{P}_{\theta}\left(\left\{U_{k}<TA_{\theta,m}-\epsilon_{\theta}-\sum_{s=1}^{k}\mathbb{E}_{\theta}\left[\text{LLR}_{\theta,m}(s)\,\middle|\,\mathcal{F}_{s-1}\right]\right\}\cap\left\{\tilde{\tau}_{\theta}\leq g_{k,l}\right\}\right)$
where $\{\mathcal{F}_{s}\}_{s\geq 0}$ is the natural filtration defined in Lemma 2 in Appendix and
$\mathrm{LLR}_{\theta,m}(s)\triangleq\log\frac{P^{A_{s},Z_{s}}_{\theta}(\mathbf{X}_{s})}{P^{A_{s},Z_{s}}_{m}(\mathbf{X}_{s})}\quad\forall\,s\geq 1,$
$U_{k}\triangleq S_{k,\theta,m}-\sum_{s=1}^{k}\mathbb{E}_{\theta}\left[\text{LLR}_{\theta,m}(s)\,\middle|\,\mathcal{F}_{s-1}\right].$ (11)
To further bound the term $(i)$, we proceed by deriving a lower bound for $\sum_{s=1}^{k}\mathbb{E}_{\theta}\left[\text{LLR}_{\theta,m}(s)\,\middle|\,\mathcal{F}_{s-1}\right]$,
$\sum_{s=1}^{k}\mathbb{E}_{\theta}\left[\text{LLR}_{\theta,m}(s)\,\middle|\,\mathcal{F}_{s-1}\right]$ $=\sum_{s=1}^{\tilde{\tau}_{\theta}}\mathbb{E}_{\theta}\left[\text{LLR}_{\theta,m}(s)\,\middle|\,\mathcal{F}_{s-1}\right]+\sum_{s=\tilde{\tau}_{\theta}+1}^{k}\mathbb{E}_{\theta}\left[\text{LLR}_{\theta,m}(s)\,\middle|\,\mathcal{F}_{s-1}\right]$
$\geq\sum_{s=\tilde{\tau}_{\theta}+1}^{k}\mathbb{E}_{\theta}\left[\text{LLR}_{\theta,m}(s)\,\middle|\,\mathcal{F}_{s-1}\right]$ (12)
$=\sum_{s=\tilde{\tau}_{\theta}+1}^{k}\sum_{a,z}P_{A_{s},Z_{s}|\mathcal{F}_{s-1}}\left(a,z\,\middle|\,\mathcal{F}_{s-1}\right){D}\left(P^{a,z}_{\theta}\,\middle\|P^{a,z}_{m}\right)$ (13)
$=\sum_{s=\tilde{\tau}_{\theta}+1}^{k}\sum_{a,z}P_{A_{s}|Z_{s},F_{s-1}}\left(a\,\middle|\,z,f_{s-1}\right)\alpha_{z}{D}\left(P^{a,z}_{\theta}\,\middle\|P^{a,z}_{m}\right)$ (14)
$=\sum_{s=\tilde{\tau}_{\theta}+1}^{k}\sum_{a,z}\left(\left(1-\epsilon(T)\right)\beta^{\theta}_{a,z}+\epsilon(T)\frac{1}{h}\right){D}\left(P^{a,z}_{\theta}\,\middle\|P^{a,z}_{m}\right)$ (15)
where (12) and (13) are because $\forall\,s\geq 1$, we have
$\mathbb{E}_{\theta}\left[\text{LLR}_{\theta,m}(s)\,\middle|\,\mathcal{F}_{s-1}\right]$ $=\sum_{a,z}P_{A_{s},Z_{s}|\mathcal{F}_{s-1}}\left(a,z\,\middle|\,\mathcal{F}_{s-1}\right)\mathbb{E}_{\theta}\left[\log\frac{P^{a,z}_{\theta}(\mathbf{X}_{s})}{P^{a,z}_{m}(\mathbf{X}_{s})}\,\middle|\,A_{s}=a,Z_{s}=z,\mathcal{F}_{s-1}\right]$ (16)
$=\sum_{a,z}P_{A_{s},Z_{s}|\mathcal{F}_{s-1}}\left(a,z\,\middle|\,\mathcal{F}_{s-1}\right){D}\left(P^{a,z}_{\theta}\,\middle\|P^{a,z}_{m}\right)\geq 0$ (17)
(16) is from law of total expectation and (17) is because given the current action $A_{s}=a$ and available set $Z_{s}=z$, the random sample set $\mathbf{X}_{s}$ follows the joint distribution $P^{a,z}_{\theta}$ and is conditionally independent of the past information $\mathcal{F}_{s-1}$ for any time slot $s\geq 1$, (14) holds since the current available set $Z_{s}$ and the past information $F_{s-1}$ are independent and we switch from conditioning on $\mathcal{F}_{s-1}$ to merely conditioning on $F_{s-1}$ without loss of generality by Definition 1, (15) is from our proposed test as $\hat{\theta}_{\text{MLE}}=\theta$. Specifically, by the definition of $\tilde{\tau}_{\theta}$, the maximum likelihood estimate $\hat{\theta}_{\text{MLE}}=\theta$ for all time slots after $\tilde{\tau}_{\theta}$, that is, $\forall\,a\in\mathcal{A}\setminus\{\emptyset\},z\in\mathcal{Z}$,
$P_{A_{s}|Z_{s},F_{s-1}}\left(a\,\middle|\,z,f_{s-1}\right)=\left(1-\epsilon(T)\right)\frac{\beta^{\theta}_{a,z}}{\alpha_{z}}+\epsilon(T)\frac{1}{h\alpha_{z}}$
Then, plugging (15) and the definition of $A_{\theta,m}$ (see Table I), the lower bound of $\sum_{s=1}^{k}\mathbb{E}_{\theta}\left[\text{LLR}_{\theta,m}(s)\,\middle|\,\mathcal{F}_{s-1}\right]$ turns to
$\sum_{s=1}^{k}\mathbb{E}_{\theta}\left[\text{LLR}_{\theta,m}(s)\,\middle|\,\mathcal{F}_{s-1}\right]\geq A_{\theta,m}\left(1+l\right)\left(k-\tilde{\tau}_{\theta}\right)$
Note that $l$ is a specific sequence defined in Table I.
Therefore, the upper bound of the term $(i)$ turns to
$(i)$ $=\mathbb{P}_{\theta}\left(\textbf{B}_{\theta}(k)\cap\left\{\tilde{\tau}_{\theta}\leq g_{k,l}\right\}\right)$
$\leq\sum_{m\neq\theta}\mathbb{P}_{\theta}\left(\left\{U_{k}<TA_{\theta,m}-\epsilon_{\theta}-A_{\theta,m}\left(1+l\right)\left(k-\tilde{\tau}_{\theta}\right)\right\}\cap\left\{\tilde{\tau}_{\theta}\leq g_{k,l}\right\}\right)$
$\leq\sum_{m\neq\theta}\mathbb{P}_{\theta}\left(\left\{U_{k}<TA_{\theta,m}-\epsilon_{\theta}-k\left(1+\frac{l}{2}\right)A_{\theta,m}\right\}\cap\left\{\tilde{\tau}_{\theta}\leq g_{k,l}\right\}\right)$ (18)
$\leq\sum_{m\neq\theta}\mathbb{P}_{\theta}\left(U_{k}<A_{\theta,m}\underbrace{\left(T-\frac{\epsilon_{\theta}}{A^{\text{max}}_{\theta}}-k\left(1+\frac{l}{2}\right)\right)}_{(*)}\right)$ (19)
where (18) is because $\tilde{\tau}_{\theta}\leq g_{k,l}=k\left(\frac{l/2}{1+l}\right)$ and $A^{\text{max}}_{\theta}\triangleq\max_{m\neq\theta}A_{\theta,m}$ (see Table I).
Next, we apply Azuma’s inequality (see Fact 1) to further derive the upper bound for the term $(i)$.
Fact 1 (Azuma’s inequality)
Let $\{U_{i}\}_{i\geq 0}\triangleq\left\{U_{0},U_{1},...\right\}$ is a martingale process. Furthermore, for any $i\geq 1$, if $\underline{U}_{i}\leq U_{i}-U_{i-1}\leq\overline{U}_{i}$ and there exists a finite sequence $\{C_{i}\}_{i\geq 1}\triangleq\left\{C_{1},C_{2},...\right\}$ such that $\overline{U}_{i}-\underline{U}_{i}\leq C_{i}$ for any $i\geq 1$, then we have for any $v>0$,
$P\left(U_{i}-U_{0}\leq-v\right)\leq\exp{\left\{\frac{-2v^{2}}{\sum_{s=1}^{i}C^{2}_{s}}\right\}}$
Specifically, through mapping the parameters $v$, $\{U_{i}\}_{i\geq 0}$, and $\{C_{i}\}_{i\geq 1}$ in Fact 1 to the corresponding values in (19), we can provide an upper bound for the term $(i)$ by exploiting Fact 1.
We start by showing that $\{U_{i}\}_{i\geq 0}$ defined as follows is a martingale process under hypothesis $\theta$ and then find the parameter $\{C_{i}\}_{i\geq 1}$.
$U_{i}=\begin{cases}\sum_{s=1}^{i}\left(\text{LLR}_{\theta,m}(s)-\mathbb{E}_{\theta}\left[\text{LLR}_{\theta,m}(s)\,\middle|\,\mathcal{F}_{s-1}\right]\right),&i\geq 1\\[4.30554pt] 0,&i=0\end{cases}$ (20)
Note that, as $i=k$, $U_{k}$ is in fact (11) by plugging the expression of $S_{k,\theta,m}$ shown in Lemma 2 in Appendix.
The martingale property of the stochastic process $\{U_{i}\}_{i\geq 0}$ is shown as follows
$\mathbb{E}_{\theta}\left[U_{i}-U_{i-1}\,\middle|\,\mathcal{F}_{i-1}\right]$ $=\mathbb{E}_{\theta}\left[\text{LLR}_{\theta,m}(i)-\mathbb{E}_{\theta}\left[\text{LLR}_{\theta,m}(i)\,\middle|\,\mathcal{F}_{i-1}\right]\,\middle|\,\mathcal{F}_{i-1}\right]$
$=\mathbb{E}_{\theta}\left[\text{LLR}_{\theta,m}(i)\,\middle|\,\mathcal{F}_{i-1}\right]-\mathbb{E}_{\theta}\left[\mathbb{E}_{\theta}\left[\text{LLR}_{\theta,m}(i)\,\middle|\,\mathcal{F}_{i-1}\right]\,\middle|\,\mathcal{F}_{i-1}\right]$
$=\mathbb{E}_{\theta}\left[\text{LLR}_{\theta,m}(i)\,\middle|\,\mathcal{F}_{i-1}\right]-\mathbb{E}_{\theta}\left[\text{LLR}_{\theta,m}(i)\,\middle|\,\mathcal{F}_{i-1}\right]=0$
Subsequently, for the parameters $\{C_{i}\}_{i\geq 1}$, by Assumption 1,we have $\forall\,i\geq 1$, $\forall\,A_{i}\in\mathcal{A}$, $\forall\,Z_{i}\in\mathcal{Z}$,
$\lvert\text{LLR}_{\theta,m}(i)\rvert=\lvert\log\frac{P^{A_{i},Z_{i}}_{\theta}(\mathbf{X}_{i})}{P^{A_{i},Z_{i}}_{m}(\mathbf{X}_{i})}\rvert\leq L\quad\forall\,\mathbf{X}_{i}\in\mathcal{X}_{\theta}^{A_{i},Z_{i}},\ \forall\,\theta,m\in\left[M\right],\ \theta\neq m.$
Hence, it is straightforward that $\forall\,i\geq 1$,
$-2L\leq U_{i}-U_{i-1}\leq 2L$
since $U_{i}-U_{i-1}=\text{LLR}_{\theta,m}(i)-\mathbb{E}_{\theta}\left[\text{LLR}_{\theta,m}(i)\,\middle|\,\mathcal{F}_{i-1}\right]$. As a result, by assigning $\overline{U}_{i}=2L$ and $\underline{U}_{i}=-2L$, the $\{C_{i}\}_{i\geq 1}$ is derived as follows
$\overline{U}_{i}-\underline{U}_{i}\leq 4L\triangleq C_{i}\quad\forall\,i\geq 1$
Finally, we proceed to find the parameter $v$ in our scenario. First of all, it is evident that, by Assumption 2, we have $\tilde{A}_{\theta,m}>0$ (see Table I), namely, $A_{\theta,m}>0$, $\forall\,\theta,m\in\left[M\right]$, $\theta\neq m$. Then, we turn to show that the term ()(*)( ∗ ) defined in (19) is smaller than zero.
The term $(*)<0$ holds true since when $k\geq\kappa_{\theta,T,l}=\lceil T-\frac{\epsilon_{\theta}}{A^{\text{max}}_{\theta}}\rceil$, we have $\forall\,T\geq q\left(l,n,I(T)\right)$,
((*)( ∗ ) $\leq T-\frac{\epsilon_{\theta}}{A^{\text{max}}_{\theta}}-\left(1+\frac{l}{2}\right)\kappa_{\theta,T,l}$
$\leq T-\frac{\epsilon_{\theta}}{A^{\text{max}}_{\theta}}-\left(1+\frac{l}{2}\right)\left(T-\frac{\epsilon_{\theta}}{A^{\text{max}}_{\theta}}\right)$
$=-\frac{l}{2}\left(T-\frac{\epsilon_{\theta}}{A^{\text{max}}_{\theta}}\right)<0$ (21)
where (21) is because $l\geq 0$ and plugging the definitions of $\epsilon_{\theta}$, $q\left(l,n,I(T)\right)$ (see Table I), and $\Delta\triangleq-MB(l)(1+b(l))e^{-b(l)(T-1)}$, we have
$T-\frac{\epsilon_{\theta}}{A^{\text{max}}_{\theta}}$ $=T-(1-\frac{K(\Delta)}{b(l)})$
$\geq T-1-\frac{1}{b(l)}$ (22)
$\geq q\left(l,n,I(T)\right)-1-\frac{1}{b(l)}=\frac{\log\left(MB(l)(1+b(l))\right)}{b(l)}>0\quad\forall\,T\geq q\left(l,n,I(T)\right)$
Notably, the function $b(l),B(l),K(\cdot)$ are defined in Table I and (22) is because as $T\geq q(l,n,I(T))$, we have $\Delta\geq-e^{-1}$, namely, $-1\leq K(\Delta)\leq 0$ by the range of the function $K(\cdot)$.
Consequently, we assign
$v=A_{\theta,m}\left(k\left(1+\frac{l}{2}\right)+\frac{\epsilon_{\theta}}{A^{\text{max}}_{\theta}}-T\right)$
After mapping to Azuma’s inequality, we are ready to bound the term $(i)$. By plugging $C_{i}=4L$ for all $i\geq 1$ and $v=A_{\theta,m}\left(k\left(1+\frac{l}{2}\right)+\frac{\epsilon_{\theta}}{A^{\text{max}}_{\theta}}-T\right)$ in Fact 1, we have
$(i)$ $\leq\sum_{m\neq\theta}\exp{\left\{\frac{-2A^{2}_{\theta,m}\left(k\left(1+\frac{l}{2}\right)+\frac{\epsilon_{\theta}}{A^{\text{max}}_{\theta}}-T\right)^{2}}{k\left(4L\right)^{2}}\right\}}$
$\leq\sum_{m\neq\theta}\exp{\left\{\frac{-A^{2}_{\theta,m}k\left(\left(1+\frac{l}{2}\right)+\frac{1}{k}\left(\frac{\epsilon_{\theta}}{A^{\text{max}}_{\theta}}-T\right)\right)^{2}}{8L^{2}}\right\}}$
$\leq\sum_{m\neq\theta}\exp{\{-\frac{A^{2}_{\theta,m}k(\frac{l}{2})^{2}}{8L^{2}}\}}$ (23)
$\leq M\exp{\{-k\frac{l^{2}}{2}(\frac{\min_{m\neq\theta}A_{\theta,m}}{4L})^{2}\}}.$
where (23) is because
$\forall\,k\geq\kappa_{\theta,T,l}\triangleq\lceil T-\frac{\epsilon_{\theta}}{A^{\text{max}}_{\theta}}\rceil,\,1+\frac{1}{k}\left(\frac{\epsilon_{\theta}}{A^{\text{max}}_{\theta}}-T\right)\geq 0$
Next, carrying the steps of showing the upper bound for the term $(i)$, we proceed to derive the upper bound for the term $(ii)$,
$(ii)$ $=\mathbb{P}_{\theta}\left(\tilde{\textbf{B}}_{\theta}(t^{\prime})\right)\leq\sum_{m\neq\theta}\mathbb{P}_{\theta}\left(S_{t^{\prime},\theta,m}<0\right)$ (24)
$=\sum_{m\neq\theta}\mathbb{P}_{\theta}\left(U_{t^{\prime}}<-\sum_{s=1}^{t^{\prime}}\mathbb{E}_{\theta}\left[\text{LLR}_{\theta,m}(s)\,\middle|\,\mathcal{F}_{s-1}\right]\right)$ (25)
where (24) holds true by union bound with the definition of $\tilde{\textbf{B}}_{\theta}(t^{\prime})$ and in (25), $U_{t^{\prime}}$ is defined in (20) as $i=t^{\prime}$.
To further bound the term $(ii)$, plugging the definition of $I(T)$ (see Table I), the lower bound of $\mathbb{E}_{\theta}\left[\text{LLR}_{\theta,m}(s)\,\middle|\,\mathcal{F}_{s-1}\right]$ for any $s\geq 1$ is derived as follows
$\mathbb{E}_{\theta}\left[\text{LLR}_{\theta,m}(s)\,\middle|\,\mathcal{F}_{s-1}\right]$ $=\sum_{a,z}P_{A_{s}|Z_{s},F_{s-1}}\left(a\,\middle|\,z,f_{s-1}\right)\alpha_{z}{D}\left(P^{a,z}_{\theta}\,\middle\|P^{a,z}_{m}\right)$
$\geq\sum_{a,z}\epsilon(T)\frac{1}{h}{D}\left(P^{a,z}_{\theta}\,\middle\|P^{a,z}_{m}\right)$ (26)
$\geq I(T)$
where (26) is directly from our proposed test, that is, $\forall\,a\in\mathcal{A}\setminus\{\emptyset\}$, $\forall\,z\in\mathcal{Z}$, $\forall\,f_{s-1}$, we have
$P_{A_{s}|Z_{s},F_{s-1}}\left(a\,\middle|\,z,f_{s-1}\right)\geq\epsilon(T)\frac{1}{h\alpha_{z}}$
Hence, the term $(ii)$ is upper bounded as follows
$(ii)=\mathbb{P}_{\theta}\left(\tilde{\textbf{B}}_{\theta}(t^{\prime})\right)\leq\sum_{m\neq\theta}\mathbb{P}_{\theta}\left(U_{t^{\prime}}<-t^{\prime}I(T)\right)$
Then, similarly, applying Fact 1 with $v=t^{\prime}I(T)$ and $C_{i}=4L$ for all $i\geq 1$, we have
$(ii)$ $\leq M\exp{\{-t^{\prime}\frac{I^{2}(T)}{8L^{2}}\}}$
$\leq M\exp{\{-t^{\prime}\frac{l^{2}}{2}(\frac{I(T)/(1+l/2)}{4L})^{2}\}}$ (27)
where (27) is because
$\frac{I^{2}\left(T\right)}{8L^{2}}$ $\geq\frac{I^{2}\left(T\right)}{8L^{2}}\left(\frac{l}{2\left(1+\frac{l}{2}\right)}\right)^{2}=\frac{l^{2}}{2}\left(\frac{I\left(T\right)/\left(1+l/2\right)}{4L}\right)^{2}$
As a result, the upper bound of $\mathbb{P}_{\theta}(\tau_{\theta}>k)$ can be obtained by combining the upper bounds of the term $(i)$ and the term $(ii)$ as follows,
$\mathbb{P}_{\theta}\left(\tau_{\theta}>k\right)$ $\leq M\exp{\{-k\frac{l^{2}}{2}(\frac{\min_{m\neq\theta}A_{\theta,m}}{4L})^{2}\}}+M\frac{\exp{\{-k\frac{l^{3}}{4(1+l)}(\frac{I(T)/(1+l/2)}{4L})^{2}\}}}{1-\exp{\{-\frac{l^{2}}{2}(\frac{I(T)/(1+l/2)}{4L})^{2}\}}}$
$\leq M\exp{\{-k\frac{l^{2}}{2(1+l)^{2}}(\frac{D}{4L})^{2}\}}+M\frac{\exp{\{-k\frac{l^{3}}{4(1+l)^{3}}(\frac{I(T)}{4L})^{2}\}}}{1-\exp{\{-\frac{l^{2}}{2(1+l)^{2}}(\frac{I(T)}{4L})^{2}\}}}$
where $D\triangleq\min_{\theta\in\left[M\right]}\min_{m\neq\theta}A_{\theta,m}$.
Subsequently, plugging the definitions of $l$, $I(T)$, and $\epsilon(T)$ (see Table I), we have $\forall\,T\geq e$,
$D^{2}$ $=lT^{\frac{1}{6}}\left(\frac{I(T)}{\epsilon(T)}\right)^{2}=T^{\frac{1}{6}}l\left(\log(T)\right)^{\frac{1}{2}}I^{2}(T)>I^{2}(T)l>I^{2}(T)\frac{l}{2(1+l)}$
where $T^{1/6}\left(\log\left(T\right)\right)^{1/2}>1$ for all $T\geq e$. Hence, it is straightforward that
$k\frac{l^{2}}{2\left(1+l\right)^{2}}\left(\frac{D}{4L}\right)^{2}>k\frac{l^{3}}{4(1+l)^{3}}\left(\frac{I(T)}{4L}\right)^{2}$ (28)
Then, plugging (28) and the definitions of $b(l)$ and $B(l)$ (see Table I), we have
$\mathbb{P}_{\theta}\left(\tau_{\theta}>k\right)$ $\leq M\exp{\left\{-k\frac{l^{3}}{4(1+l)^{3}}\left(\frac{I(T)}{4L}\right)^{2}\right\}}\left(1+\frac{1}{1-\exp{\left\{-\frac{l^{2}}{2(1+l)^{2}}\left(\frac{I(T)}{4L}\right)^{2}\right\}}}\right)$
$\leq M\exp{\left\{-kb(l)\right\}}\left(2+\frac{1}{\frac{l^{2}}{2(1+l)^{2}}\left(\frac{I(T)}{4L}\right)^{2}}\right)$ (29)
$=M\exp{\left\{-kb(l)\right\}}B\left(l\right)$
where (29) holds true since $\frac{1}{1-e^{-x}}\leq 1+\frac{1}{x},\forall\,x>0$.
As a result, by plugging the definitions of $\epsilon_{\theta}$, $\Delta$, and $\kappa_{\theta,T,l}$ the satisfaction of the expected stopping time constraints is ready to be shown:
$\mathbb{E}_{\theta}\left[\tau_{\theta}\right]$ $\leq\kappa_{\theta,T,l}+\sum_{k=\kappa_{\theta,T,l}}^{\infty}\mathbb{P}_{\theta}\left(\tau_{\theta}>k\right)$
$\leq T+1-\frac{\epsilon_{\theta}}{A^{\text{max}}_{\theta}}+MB(l)\frac{e^{-\kappa_{\theta,T,l}b(l)}}{1-e^{-b(l)}}$
$\leq T+\frac{1}{b(l)}K(\Delta)-\frac{1}{b(l)}\Delta e^{-K(\Delta)}$ (30)
$=T+\frac{1}{b(l)}K(\Delta)-\frac{1}{b(l)}W_{0}(\Delta)e^{W_{0}(\Delta)}e^{-K(\Delta)}$ (31)
$\leq T+\frac{1}{b(l)}K(\Delta)-\frac{1}{b(l)}K(\Delta)$ (32)
$=T\quad\forall\,T\geq q(l,n,I(T))$
where (30) is due to
$MB(l)\frac{e^{-\kappa_{\theta,T,l}b(l)}}{1-e^{-b(l)}}$ $\leq MB(l)\left(1+\frac{1}{b(l)}\right)e^{-\kappa_{\theta,T,l}b(l)}$
$=MB(l)\left(1+\frac{1}{b(l)}\right)e^{-\lceil T-\frac{\epsilon_{\theta}}{A^{\text{max}}_{\theta}}\rceil b(l)}$
$=MB(l)\left(1+\frac{1}{b(l)}\right)e^{-\lceil T-\left(1-\frac{K(\Delta)}{b(l)}\right)\rceil b(l)}$
$\leq MB(l)\left(1+\frac{1}{b(l)}\right)e^{-\left(T-\left(1-\frac{K(\Delta)}{b(l)}\right)\right)b(l)}$
$=\frac{1}{b(l)}MB(l)\left(b(l)+1\right)e^{-b(l)\left(T-1\right)}e^{-K(\Delta)}$
$=-\frac{1}{b(l)}\Delta e^{-K(\Delta)}$
(31) is due to the fact that $-e^{-1}\leq\Delta<0$, $\forall\,T\geq q(l,n,I(T))$ and $W_{0}(\cdot)$ is the principle branch of Lambert function. Moreover, (32) holds true due to the fact that $xe^{x}$ is increasing when $x\geq-1$ and by [11, Theorem 3.2],
$K(x)\leq W_{0}(x),\,-e^{-1}\leq x<0$
Specifically, by [11, Theorem 3.2] and the range of $\Delta$, we have $W_{0}(\Delta)\geq K(\Delta)\geq-1$ so that it is straightforward that $W_{0}(\Delta)e^{W_{0}(\Delta)}\geq K(\Delta)e^{K(\Delta)}$.
To end the proof of the satisfaction of the expected stopping time constraints, we have, $\forall\,T\geq\max\{e,q\left(l,n,I(T)\right)\}$,
$\mathbb{E}_{\theta}[\tau]\leq\mathbb{E}_{\theta}[\tau_{\theta}]\leq T\quad\forall\,\theta\in\left[M\right]$
We then show that our proposed test satisfies the expected budget constraints, i.e., $\forall\,\theta\in\left[M\right]$, $\underline{B}^{\theta}\in\mathcal{C}_{B}$.
IV-C2 Linear expected budget constraints
Define
$\underline{B}_{t}\triangleq\left\{B_{t,1},B_{t,2},\ldots,B_{t,n}\right\}$
where we denote $B_{t,j}\triangleq\sum_{s=1}^{t}\mathbbm{1}{\left\{j\in A_{t}\cap Z_{t}\right\}}$ as the total number of selecting source $j$ until time instance $t$, $\forall\,j\in\left[n\right],\forall\,t\geq 1$.
We then adapt a martingale argument to prove the satisfaction of the linear expected budget constraints.
First of all, we construct a stochastic process $\forall\,i\in\mathcal{I}$
$\left\{M_{t,i}\triangleq b_{i}\left(\underline{B}_{t}\right)-r_{i}t\,\middle|\,t\geq 0\right\}$
where $b_{i}(\cdot)\ \forall\,i\in\mathcal{I}$ is a linear function defined in $\mathcal{C}_{B}$. We proceed to show that this stochastic process is a super martingale adapted to the natural filtration $\{\mathcal{F}_{t}\}_{t\geq 0}$ under any given hypothesis $\theta\in\left[M\right]$. Specifically, we have
$\mathbb{E}_{\theta}\left[M_{t+1,i}\,\middle|\,\mathcal{F}_{t}\right]$ $=\mathbb{E}_{\theta}\left[b_{i}\left(\underline{B}_{t+1}\right)-\left(t+1\right)r_{i}\,\middle|\,\mathcal{F}_{t}\right]$
$=b_{i}(\underline{B}_{t})-r_{i}t+\mathbb{E}_{\theta}\left[b_{i}\left(\left\{\mathbbm{1}{\left\{j\in A_{t+1}\cap Z_{t+1}\right\}}\right\}_{j=1}^{n}\right)\,\middle|\,\mathcal{F}_{t}\right]-r_{i}$
$\leq M_{t,i}$ (33)
where (33) holds true because
$\mathbb{E}_{\theta}\left[b_{i}\left(\left\{\mathbbm{1}{\left\{j\in A_{t+1}\cap Z_{t+1}\right\}}\right\}_{j=1}^{n}\right)\,\middle|\,\mathcal{F}_{t}\right]$
$=b_{i}\left(\left\{\mathbb{E}_{\theta}\left[\mathbbm{1}{\left\{j\in A_{t+1}\cap Z_{t+1}\right\}}\,\middle|\,\mathcal{F}_{t}\right]\right\}_{j=1}^{n}\right)$
$=b_{i}\left(\left\{\sum_{a,z}P_{A_{t+1}|Z_{t+1},F_{t}}\left(a|z,f_{t}\right)\alpha_{z}\mathbbm{1}{\left\{j\in a\cap z\right\}}\right\}_{j=1}^{n}\right)$
$\leq\max_{\theta\in\left[M\right]}b_{i}\left(\left\{\sum_{a,z}\left(\left(1-\epsilon\left(T\right)\right)\beta^{\theta}_{a,z}+\epsilon(T)\frac{1}{h}\right)\mathbbm{1}{\left\{j\in a\cap z\right\}}\right\}_{j=1}^{n}\right)$ (34)
$=\max_{\theta\in\left[M\right]}\left(1-\epsilon(T)\right)b_{i}\left(\left\{\sum_{a,z}\beta^{\theta}_{a,z}\mathbbm{1}{\left\{j\in a\cap z\right\}}\right\}_{j=1}^{n}\right)+\epsilon(T)\frac{1}{h}b_{i}\left(\left\{\sum_{a,z}\mathbbm{1}{\left\{j\in a\cap z\right\}}\right\}_{j=1}^{n}\right)$
$\leq(1-\epsilon(T))r_{i}+\epsilon(T)r_{i}$ (35)
$=r_{i}$
Note that (34) is from the action taking policy defined in our proposed test, (35) is due to the linearity of $b_{i}(\cdot)$, $\bm{\beta}^{\theta}\in\mathcal{C}$ for all $\theta\in\left[M\right]$, and the design of $\frac{1}{h}$ in our proposed test.
Consequently, from the fact that $\mathbb{E}_{\theta}[\tau]\leq T$ shown in Section IV-C1 and Optional Stopping Theorem, we have
$0=\mathbb{E}_{\theta}\left[M_{0,j}\right]\geq\mathbb{E}_{\theta}\left[M_{\tau,j}\right]\geq\mathbb{E}_{\theta}\left[b_{i}\left(\underline{B}_{\tau}\right)\right]-r_{i}T=b_{i}\left(\left\{\mathbb{E}_{\theta}\left[B_{\tau,j}\right]\right\}_{j=1}^{n}\right)-r_{i}T$
Hence, plugging the definition of $\underline{B}^{\theta}$ (see Section II-C), we have
$b_{i}\left(\underline{B}^{\theta}\right)=b_{i}\left(\left\{\mathbb{E}_{\theta}\left[B_{\tau,j}\right]\right\}_{j=1}^{n}\right)\leq r_{i}T=U_{i}\quad\forall\,\theta\in\left[M\right],\forall\,i\in\mathcal{I},$
thereby $\underline{B}^{\theta}\in C_{B}$, $\forall\,\theta\in\left[M\right]$.
IV-C3 Achievable error exponents
Given any hypotheses $\theta$, $\forall\,m\neq\theta$, as $T$ is sufficiently large,
$\pi_{m|\theta}$ $=\mathbb{P}_{\theta}\left(\min_{i\neq m}\left\{S_{\tau,m,i}-TA_{m,i}+\epsilon_{m}\right\}\geq 0\right)\mathbbm{1}{\left\{T\geq q\left(l,n,I(T)\right)\right\}}+\frac{1}{M}\mathbbm{1}{\left\{T<q\left(l,n,I(T)\right)\right\}}$
$=\mathbb{P}_{\theta}\left(\min_{i\neq m}\left\{S_{\tau,m,i}-TA_{m,i}+\epsilon_{m}\right\}\geq 0\right)$ (36)
$\leq\mathbb{P}_{\theta}\left(S_{\tau,m,\theta}-TA_{m,\theta}+\epsilon_{m}\geq 0\right)$
$\leq\frac{\mathbb{E}_{\theta}\left[e^{S_{\tau,m,\theta}}\right]}{e^{TA_{m,\theta}-\epsilon_{m}}}=e^{-T\left(A_{m,\theta}-\frac{\epsilon_{m}}{T}\right)}$ (37)
where (36) holds true by Lemma 1 in Appendix, implying when $T$ is sufficiently large, $T\geq q(l,n,I(T))$. Furthermore, (37) is due to Markov’s inequality and Optional Stopping Theorem, namely, by Lemma 2 in Appendix, we have $\left\{e^{S_{t,m,\theta}}\right\}_{t\geq 1}$ is a martingale with respect to the natural filtration under hypothesis $\theta$, leading to $\mathbb{E}_{\theta}\left[e^{S_{\tau,m,\theta}}\right]=\mathbb{E}_{\theta}\left[e^{S_{0,m,\theta}}\right]=1$, $\forall\,\theta,m\in\left[M\right],\theta\neq m$.
Therefore, $\forall\,\theta,m\in\left[M\right]$, $\theta\neq m$, we can conclude
$\liminf_{T\to\infty}\left\{\frac{1}{T}\log\frac{1}{\pi_{m|\theta}}\right\}\geq\liminf_{T\to\infty}\left\{A_{m,\theta}-\frac{\epsilon_{m}}{T}\right\}=e^{*}_{m|\theta}(\bm{\beta}^{m})$
where $\epsilon_{m}\in o(T)$ defined in Table I for any $m\in\left[M\right]$. ∎
V Proof of Converse
Let us first introduce the following notation: for $p,q\in[0,1]$,
${d}\left(p\,\middle\|q\right)$ $\textstyle\triangleq p\log\frac{p}{q}+(1-p)\log\frac{1-p}{1-q}={D}\left(\mathrm{Ber}(p)\,\middle\|\mathrm{Ber}(q)\right),$
$h(p)$ $\textstyle\triangleq p\log\frac{1}{p}+(1-p)\log\frac{1}{1-p}=H(\mathrm{Ber}(p)).$
To prove the converse part of Theorem 1, it is sufficient to show that, for any achievable $M(M-1)$ types of error exponents $\Big{\{}e_{m|\theta}\,\Big{|}\,m\in\left[M\right],\theta\in\left[M\right]\setminus\{m\}\Big{\}}$, there exists $\bm{\beta}^{0},\bm{\beta}^{1},...,\bm{\beta}^{M-1}\in\mathcal{C}$ such that $\forall\,m\in\left[M\right]$ $\forall\,\theta\in\left[M\right]\setminus\{m\}$,
$e_{m|\theta}\leq\sum_{a\in\mathcal{A}}\sum_{z\in\mathcal{Z}}\beta^{m}_{a,z}{D}\left(P^{a,z}_{m}\,\middle\|P^{a,z}_{\theta}\right),$
thereby the $\Big{\{}e_{m|\theta}\,\Big{|}\,m\in\left[M\right],\theta\in\left[M\right]\setminus\{m\}\Big{\}}\in\zeta$ where $\zeta$ is the optimal error exponent region defined in Theorem 1.
Before we proceed to the proof of the converse part, it is firstly summarized as follows: First, we utilize the data processing inequality and Optional Stopping Theorem to find the upper bound of the $\frac{1}{T}\log\frac{1}{\pi_{m|\theta}}$ for any $m\in\left[M\right]$, $\theta\in\left[M\right]\setminus\{m\}$, $\forall\,T\geq 1$, where the action taking frequency shown in the upper bound plays a key role. Then, we derive the necessary conditions for the action-taking frequency to company the linear expected budget constraints including expected stopping time constraints. We can observe that the collection of action-taking frequencies satisfying the necessary conditions is the constraint set $\mathcal{C}$ defined in Theorem 1. Finally, we exploit the fact that $\mathcal{C}$ is a compact set to complete the proof.
We start the proof of the converse part with an useful decomposition of $\log\frac{1}{\pi_{m|\theta}},\ \forall\,m\in\left[M\right],\ \forall\,\theta\in\left[M\right]\setminus\{m\}$:
$\log\frac{1}{\pi_{m|\theta}}=\frac{1}{\pi_{m|m}}\left({d}\left(\pi_{m|m}\,\middle\|\pi_{m|\theta}\right)+h\left(\pi_{m|m}\right)+\mathbb{P}_{m}\left(\phi\neq m\right)\log\mathbb{P}_{\theta}\left(\phi\neq m\right)\right)$ (38)
where the decomposition is derived as follows,
${d}\left(\pi_{m|m}\,\middle\|\pi_{m|\theta}\right)$ $=\mathbb{P}_{m}\left(\phi=m\right)\log{\left(\frac{\mathbb{P}_{m}\left(\phi=m\right)}{\mathbb{P}_{\theta}\left(\phi=m\right)}\right)}+\mathbb{P}_{m}\left(\phi\neq m\right)\log{\left(\frac{\mathbb{P}_{m}\left(\phi\neq m\right)}{\mathbb{P}_{\theta}\left(\phi\neq m\right)}\right)}$
$=-\mathbb{P}_{m}\left(\phi=m\right)\log\frac{1}{\mathbb{P}_{m}\left(\phi=m\right)}-\mathbb{P}_{m}\left(\phi\neq m\right)\log\frac{1}{\mathbb{P}_{m}\left(\phi\neq m\right)}$
$\quad-\mathbb{P}_{m}\left(\phi=m\right)\log\mathbb{P}_{\theta}\left(\phi=m\right)-\mathbb{P}_{m}\left(\phi\neq m\right)\log\mathbb{P}_{\theta}\left(\phi\neq m\right)$
$=-h\left(\pi_{m|m}\right)-\pi_{m|m}\log\pi_{m|\theta}-\mathbb{P}_{m}\left(\phi\neq m\right)\log\mathbb{P}_{\theta}\left(\phi\neq m\right)$
Hence, it is clear that the decomposition holds.
Exploiting this decomposition, we firstly utilize the data processing inequality,
${d}\left(\pi_{m|m}\,\middle\|\pi_{m|\theta}\right)\leq{D}\left(\mathsf{P}_{m}^{(\tau)}\,\middle\|\mathsf{P}_{\theta}^{(\tau)}\right)=\mathbb{E}_{m}\left[S_{\tau,m,\theta}\right]$ (39)
where
$S_{\tau,m,\theta}$ $=\sum_{s=1}^{\tau}\log\frac{P^{A_{s},Z_{s}}_{m}(\mathbf{X}_{s})}{P^{A_{s},Z_{s}}_{\theta}(\mathbf{X}_{s})}$
$=\sum_{s=1}^{\tau}\sum_{a\in\mathcal{A}}\sum_{z\in\mathcal{Z}}\mathbbm{1}\{A_{s}=a,Z_{s}=z\}\log\frac{P^{a,z}_{m}(\mathbf{X}_{s})}{P^{a,z}_{\theta}(\mathbf{X}_{s})}$
Note that this explicit form of the cumulative log-likelihood ratio is derived in Lemma 2 in Appendix.
Then, by Optional Stopping Theorem, we have
$\mathbb{E}_{m}\left[S_{\tau,m,\theta}\right]=\sum_{a\in\mathcal{A}}\sum_{z\in\mathcal{Z}}\mathbb{E}_{m}\left[\sum_{s=1}^{\tau}\mathbbm{1}{\left\{A_{s}=a,Z_{s}=z\right\}}\right]{D}\left(P^{a,z}_{m}\,\middle\|P^{a,z}_{\theta}\right)$ (40)
Subsequently, by defining the following frequency for any $a\in\mathcal{A}$ and $z\in\mathcal{Z}$,
$\beta^{m,(T)}_{a,z}\triangleq\frac{\mathbb{E}_{m}\left[\sum_{s=1}^{\tau}\mathbbm{1}{\left\{A_{s}=a,Z_{s}=z\right\}}\right]}{T}\quad\forall\,T\geq 1,$ (41)
the necessary conditions that $\bm{\beta}^{m,(T)}=\left\{\beta^{m,(T)}_{a,z}\,\middle|\,a\in\mathcal{A},z\in\mathcal{Z}\right\}$ for any $m\in\left[M\right]$ should satisfy for achieving the constraints (3) are shown in the following:
Firstly, for the linear expected budget constraints, plugging the definition of $B^{m}_{j}$ (see Section II-C), we have, $\forall\,j\in\left[n\right]$,
$B^{m}_{j}$ $=\sum_{t=1}^{\infty}\mathbb{E}_{m}\left[\mathbbm{1}{\left\{j\in A_{t}\cap Z_{t}\right\}}\mathbbm{1}{\left\{t\leq\tau\right\}}\right]$
$=\sum_{t=1}^{\infty}\sum_{a,z}\mathbb{P}_{m}\left(A_{t}=a,Z_{t}=z\right)\mathbbm{1}{\left\{j\in a\cap z\right\}}\mathbb{E}_{m}\left[\mathbbm{1}{\left\{t\leq\tau\right\}}\,\middle|\,A_{t}=a,Z_{t}=z\right]$
$=\sum_{a,z}\mathbbm{1}{\left\{j\in a\cap z\right\}}\sum_{t=1}^{\infty}\mathbb{P}_{m}\left(A_{t}=a,Z_{t}=z,t\leq\tau\right)$
$=\sum_{a,z}\mathbbm{1}{\left\{j\in a\cap z\right\}}\mathbb{E}_{m}\left[\sum_{t=1}^{\tau}\mathbbm{1}{\left\{A_{t}=a,Z_{t}=z\right\}}\right].$
By $\underline{B}^{m}\in\mathcal{C}_{B}$ for all $m\in\left[M\right]$, we have $b_{i}(\underline{B}^{m})\leq U_{i}\ \forall\,i\in\mathcal{I}$, that is, plugging the definition of $\beta^{m,(T)}_{a,z}$ in (41) we have
$b_{i}\left(\left\{\sum_{a\in\mathcal{A}}\sum_{z\in\mathcal{Z}}\mathbbm{1}{\left\{j\in a\cap z\right\}}\beta^{m,(T)}_{a,z}\right\}_{j=1}^{n}\right)\leq\frac{U_{i}}{T}=r_{i}\quad\forall\,i\in\mathcal{I},$
where the both sides of the inequality $b_{i}(\underline{B}^{m})\leq U_{i}\ \forall\,i\in\mathcal{I}$ are divided by $T$ for any $T\geq 1$. Note that each $b_{i}(\cdot)$ is a linear function.
Then, for the expected stopping time constraint, the corresponding necessary condition that $\bm{\beta}^{m,(T)}$ should satisfy is expressed as follows: $\forall\,z\in\mathcal{Z}$,
$\sum_{a\in\mathcal{A}}\beta^{m,(T)}_{a,z}=\frac{\mathbb{E}_{m}\left[\sum_{s=1}^{\tau}\mathbbm{1}{\left\{Z_{s}=z\right\}}\right]}{T}=\frac{\mathbb{E}_{m}\left[\tau\right]\alpha_{z}}{T}\leq\alpha_{z}.$
Therefore, we have $\bm{\beta}^{m,(T)}\in\mathcal{C},\ \forall\,T\geq 1,m\in\left[M\right]$, where $\mathcal{C}$ is defined in Theorem 1.
After deriving the necessary conditions of each $\bm{\beta}^{m,(T)}$, we combine (38), (39) and (40) with $\bm{\beta}^{m,(T)}\in\mathcal{C}$ to get $\forall T\geq 1$, $\forall\,m\in\left[M\right]$, $\forall\,\theta\in\left[M\right]\setminus\{m\}$:
$\frac{1}{T}\log\frac{1}{\pi_{m|\theta}}$ $\leq\frac{1}{\pi_{m|m}}\left(\sum_{a\in\mathcal{A}}\sum_{z\in\mathcal{Z}}\beta^{m,(T)}_{a,z}{D}\left(P^{a,z}_{m}\,\middle\|P^{a,z}_{\theta}\right)+\frac{h\left(\pi_{m|m}\right)}{T}+\frac{\mathbb{P}_{m}\left(\phi\neq m\right)\log\mathbb{P}_{\theta}\left(\phi\neq m\right)}{T}\right)$
$\leq\frac{1}{\pi_{m|m}}\left(\sum_{a\in\mathcal{A}}\sum_{z\in\mathcal{Z}}\beta^{m,(T)}_{a,z}{D}\left(P^{a,z}_{m}\,\middle\|P^{a,z}_{\theta}\right)+\frac{h\left(\pi_{m|m}\right)}{T}\right)$ (42)
where (42) holds since $\mathbb{P}_{\theta}\left(\phi\neq m\right)\leq 1$.
We then take liminf on both sides of the inequality (42) and get:
$\liminf_{T\to\infty}\left\{\frac{1}{T}\log\frac{1}{\pi_{m|\theta}}\right\}\leq\liminf_{T\to\infty}\left\{\sum_{a\in\mathcal{A}}\sum_{z\in\mathcal{Z}}\beta^{m,(T)}_{a,z}{D}\left(P^{a,z}_{m}\,\middle\|P^{a,z}_{\theta}\right)\right\}$
where $\pi_{m|m}$ converges to one as $T$ goes to infinity and $h(\pi_{m|m})\leq\log(2)$.
Now, since $\mathcal{C}$ is a compact set, by Bolzano-Weierstrass theorem, there exists a convergent subsequence of $\left\{\bm{\beta}^{m,(T)}\,\middle|\,T\geq 1\right\}$ indexed by some $\left\{T_{k}\right\}_{k\in{\mathbb{N}_{+}}}$, i.e., $\left\{\bm{\beta}^{m,(T_{k})}\,\middle|\,k\in{\mathbb{N}_{+}}\right\}$, converging to a $\bm{\beta}^{m,*}\in\mathcal{C}$. Furthermore, because the liminf over a sequence is upper bounded by the liminf over any of its subsequences, we have $\forall\,m\in\left[M\right]$, $\forall\,\theta\in\left[M\right]\setminus\{m\}$,
$\liminf_{T\to\infty}\left\{\frac{1}{T}\log\frac{1}{\pi_{m|\theta}}\right\}$ $\leq\liminf_{k\to\infty}\left\{\sum_{a\in\mathcal{A}}\sum_{z\in\mathcal{Z}}\beta^{m,(T_{k})}_{a,z}{D}\left(P^{a,z}_{m}\,\middle\|P^{a,z}_{\theta}\right)\right\}$
$=\sum_{a\in\mathcal{A}}\sum_{z\in\mathcal{Z}}\beta^{m,*}_{a,z}{D}\left(P^{a,z}_{m}\,\middle\|P^{a,z}_{\theta}\right)$
where the equality holds true since $\bm{\beta}^{m,(T_{k})}$ converges to $\bm{\beta}^{m,*}$ and those KL-divergences are constant.
Consequently, for any achievable $M(M-1)$ types of error exponents $\Big{\{}e_{m|\theta}\,\Big{|}\,m\in\left[M\right],\theta\in\left[M\right]\setminus\{m\}\Big{\}}$, we have shown that $\forall\,m\in\left[M\right]$, there exists a $\bm{\beta}^{m}\in\mathcal{C}$ ($\bm{\beta}^{m,*}$ here) such that:
$e_{m|\theta}\leq\sum_{a\in\mathcal{A}}\sum_{z\in\mathcal{Z}}\beta^{m}_{a,z}{D}\left(P^{a,z}_{m}\,\middle\|P^{a,z}_{\theta}\right)\quad\forall\,\theta\in\left[M\right]\setminus\{m\}$
thereby completing the converse proof, i.e., $\Big{\{}e_{m|\theta}\,\Big{|}\,m\in\left[M\right],\theta\in\left[M\right]\setminus\{m\}\Big{\}}\in\zeta$. ∎
VI Proof of Proposition 1
To prove Proposition 1, due to the direct product form of the region $\zeta$, it is sufficient to show that, $\forall\,m\in\left[M\right]$,
$\cup_{\bm{\beta}^{m}\in\mathcal{C}}\zeta_{m}(\bm{\beta}^{m})=\text{Conv}\left\{\text{extr}\left(\cup_{\bm{\beta}^{m}\in V_{c}}\zeta_{m}(\bm{\beta}^{m})\right)\right\}$
Note that for notational simplicity, we let $\underline{x}\triangleq\left\{x_{i}\,\middle|\,i\in\left[M\right]\setminus\{m\}\right\}$ and in the following proof, we just consider a fixed index $m\in\left[M\right]$ without loss of generality. Additionally, the definition of an extreme point is that, for any set $A$, a point $a\in A$ is an extreme point of the set $A$ if for any $b,c\in A$, $b\neq c$ and for any $\lambda\in(0,1)$, the point $a$ satisfies $a\neq\lambda b+(1-\lambda)c$.
Direction “LHS $\subseteq$ RHS"
For any $\underline{x}\in\cup_{\bm{\beta}^{m}\in\mathcal{C}}\zeta_{m}(\bm{\beta}^{m})$, there exists a $\tilde{\bm{\beta}}^{m}\in\mathcal{C}$ and some $\eta_{1},...,\eta_{M-1}\in\left[0,1\right]$ such that
$x_{i}=\eta_{i}e^{*}_{m|i}(\tilde{\bm{\beta}}^{m})\quad\forall\,i\in\left[M\right]\setminus\{m\}$
where $e^{*}_{m|i}(\tilde{\bm{\beta}}^{m})$ is defined in Theorem 1.
Moreover, since the constraint set $\mathcal{C}$ is clearly a convex polytope as $|\mathcal{I}|<\infty$, we assume that $V_{c}=\{\tilde{\underline{v}}_{1},...,\tilde{\underline{v}}_{p}\}$ where $p<\infty$. Therefore, there exists a sequence of $\lambda_{1},..,\lambda_{p}\in\left[0,1\right]$ and $\sum_{j=1}^{p}\lambda_{j}=1$ such that $\tilde{\bm{\beta}}^{m}=\sum_{j=1}^{p}\lambda_{j}\tilde{\underline{v}}_{j}$.
As a result, by the definition of $e^{*}_{m|i}(\tilde{\bm{\beta}}^{m})$, we have, $\forall\,i\in\left[M\right]\setminus\{m\}$,
$x_{i}=\eta_{i}e^{*}_{m|i}\left(\sum_{j=1}^{p}\lambda_{j}\tilde{\underline{v}}_{j}\right)=\sum_{j=1}^{p}\lambda_{j}\eta_{i}e^{*}_{m|i}\left(\tilde{\underline{v}}_{j}\right)$
Then, due to $\left\{\eta_{i}e^{*}_{m|i}\left(\tilde{v}_{j}\right)\,\middle|\,i\in\left[M\right]\setminus\{m\}\right\}\in\zeta_{m}\left(\tilde{\underline{v}}_{j}\right)$ for any $j\in\{1,...,p\}$, it is straightforward that
$\underline{x}\in\text{Conv}\left\{\cup_{\bm{\beta}^{m}\in V_{c}}\zeta_{m}(\bm{\beta}^{m})\right\}$
Note that the index $m$ in $\bm{\beta}^{m}$ here is just for the notational consistency with respect to our main theorem.
Furthermore, because the region $\text{Conv}\left\{\cup_{\bm{\beta}^{m}\in V_{c}}\zeta_{m}(\bm{\beta}^{m})\right\}$ is a compact convex set, by Krein–Milman Theorem (a well-known result in Topology), we have
$\text{Conv}\left\{\cup_{\bm{\beta}^{m}\in V_{c}}\zeta_{m}(\bm{\beta}^{m})\right\}=\text{Conv}\left\{\text{extr}\left(\text{Conv}\left\{\cup_{\bm{\beta}^{m}\in V_{c}}\zeta_{m}(\bm{\beta}^{m})\right\}\right)\right\}$
Consequently, by the fact that the convex hull operation on the set $\cup_{\bm{\beta}^{m}\in V_{c}}\zeta_{m}(\bm{\beta}^{m})$ does not add or eliminate any extreme points of $\cup_{\bm{\beta}^{m}\in V_{c}}\zeta_{m}(\bm{\beta}^{m})$, namely,
$\text{extr}\left(\text{Conv}\left\{\cup_{\bm{\beta}^{m}\in V_{c}}\zeta_{m}(\bm{\beta}^{m})\right\}\right)=\text{extr}\left(\cup_{\bm{\beta}^{m}\in V_{c}}\zeta_{m}(\bm{\beta}^{m})\right)$
we can conclude that
$\text{Conv}\left\{\text{extr}\left(\text{Conv}\left\{\cup_{\bm{\beta}^{m}\in V_{c}}\zeta_{m}(\bm{\beta}^{m})\right\}\right)\right\}=\text{Conv}\left\{\text{extr}\left(\cup_{\bm{\beta}^{m}\in V_{c}}\zeta_{m}(\bm{\beta}^{m})\right)\right\}$
thereby
$\cup_{\bm{\beta}^{m}\in\mathcal{C}}\zeta_{m}(\bm{\beta}^{m})\subseteq\text{Conv}\left\{\cup_{\bm{\beta}^{m}\in V_{c}}\zeta_{m}(\bm{\beta}^{m})\right\}=\text{Conv}\left\{\text{extr}\left(\cup_{\bm{\beta}^{m}\in V_{c}}\zeta_{m}(\bm{\beta}^{m})\right)\right\}$
Direction “LHS $\supseteq$ RHS"
Since $\mathcal{C}$ is a convex polytope, $V_{c}$ is a finite set when $|\mathcal{I}|<\infty$. Hence, we assume there are $k<\infty$ extreme points of the region $\cup_{\bm{\beta}^{m}\in V_{c}}\zeta_{m}(\bm{\beta}^{m})$, i.e., $\text{extr}\left(\cup_{\bm{\beta}^{m}\in V_{c}}\zeta_{m}(\bm{\beta}^{m})\right)=\{\underline{v}_{1},\underline{v}_{2},...,\underline{v}_{k}\}$. Then, we have $\forall\,\underline{x}\in\text{Conv}\left\{\text{extr}\left(\cup_{\bm{\beta}^{m}\in V_{c}}\zeta_{m}(\bm{\beta}^{m})\right)\right\}$,
$\exists\lambda_{1},\lambda_{2},...,\lambda_{k}\in[0,1],\,\sum_{i=1}^{k}\lambda_{i}=1$
such that
$\underline{x}=\sum_{i=1}^{k}\lambda_{i}\underline{v}_{i}$
Furthermore, we know $\forall\,i\in\{1,...,k\}$,
$\underline{v}_{i}\in\cup_{\bm{\beta}^{m}\in V_{c}}\zeta_{m}(\bm{\beta}^{m})\subseteq\cup_{\bm{\beta}^{m}\in\mathcal{C}}\zeta_{m}(\bm{\beta}^{m}),$
and $\cup_{\bm{\beta}^{m}\in\mathcal{C}}\zeta_{m}(\bm{\beta}^{m})$ is also a convex set. Therefore, we have
$\underline{x}\in\cup_{\bm{\beta}^{m}\in\mathcal{C}}\zeta_{m}(\bm{\beta}^{m})$
As a result,
$\cup_{\bm{\beta}^{m}\in\mathcal{C}}\zeta_{m}(\bm{\beta}^{m})\supseteq\text{Conv}\left\{\text{extr}\left(\cup_{\bm{\beta}^{m}\in V_{c}}\zeta_{m}(\bm{\beta}^{m})\right)\right\}$
VII Conclusion
In this work, we have characterized the optimal individual error exponent region in active sequential multiple-hypothesis testing problems encountering temporally available data sources under a family of linear budget constraints including expected stopping time, showing that the region is a convex polytope if the numbers of budget constraints are finite. Then, through specializing our main theorem under Chernoff’s setting, we have illustrated that the tradeoffs depend on the heterogeneous level of informativeness across data sources in distinguishing a declared hypothesis against other ground truth hypotheses. To the best of our knowledge, such tradeoffs among individual error exponents have not been identified in the literature before. At the end of this work, we have shown that our main theorem covers the previous results [1, 2, 3] that do not delve into individual error probabilities for $M\geq 3$, leading to no tradeoffs.
Appendix A Appendix
Lemma 1
$q(l,n,I(T))$ is sublinear in $T$.
Proof:
By the definition of $q(l,n,I(T))$ (see Table I), it suffices to show that $\frac{1}{b(l)}\log\left(MB(l)\left(1+b(l)\right)\right)\in o(T)$. Note the following:
  1. 1.
    $\tilde{A}_{\theta,m}$, $A_{\theta,m}$ are finite for any two different hypotheses $\theta,m\in\left[M\right]$ due to Assumption 1.
  2. 2.
    $M$, $L$ are also finite.
By plugging the definitions of the respective terms, including $l$, $B(l)$, $b(l)$, etc., as shown in Table I, straightforward calculations lead to
$\frac{1}{b(l)}\log\left(MB(l)\left(1+b(l)\right)\right)\in O\left(T^{\frac{5}{6}}\left(\log(T)\right)^{\frac{1}{2}}\right)$
and the proof is complete. ∎
Lemma 2
The stochastic process
$\left\{e^{S_{t,m,\theta}}\,\middle|\,t\geq 0\right\}$
is a martingale adapted to the natural filtration $\left\{\mathcal{F}_{t}\right\}_{t\geq 0}$ under hypothesis $\theta$, $\forall\,\theta,m\in\left[M\right]$, $\theta\neq m$. $\mathcal{F}_{t}\triangleq\sigma(F_{t})$ for all $t\geq 0$.
Proof:
We first derive the explicit form of $\mathsf{P}_{m}^{(t)}$ for any $m\in\left[M\right]$ and any $t\geq 1$, demonstrating that $S_{t,m,\theta}$ can be expressed as the sum of log-likelihood ratios. Subsequently, we show the martingale property with the additivity of the sum of log-likelihood ratios.
Firstly, let us derive the explicit form of $S_{t,m,\theta}$. We start with the following explicit expression, $\forall\,t\geq 1$, $\forall\,f_{t}$, $\forall\,m\in\left[M\right]$,
$\mathsf{P}_{m}^{(t)}\left(f_{t}\right)$ $=\mathbb{P}_{m}\left(\mathbf{x}_{1},z_{1},a_{1}\right)\prod_{s=2}^{t}\mathbb{P}_{m}\left(\mathbf{x}_{s},z_{s},a_{s}\,\middle|\,f_{s-1}\right)$
$=\mathbb{P}_{m}\left(\mathbf{x}_{1},z_{1},a_{1}\right)\prod_{s=2}^{t}P_{Z}\left(z_{s}\right)\mathbb{P}_{m}\left(\mathbf{x}_{s},a_{s}\,\middle|\,f_{s-1},z_{s}\right)$ (43)
$=\mathbb{P}_{m}\left(\mathbf{x}_{1},z_{1},a_{1}\right)\prod_{s=2}^{t}P_{Z}\left(z_{s}\right)P_{A_{s}|Z_{s},F_{s-1}}\left(a_{s}\,\middle|\,z_{s},f_{s-1}\right)\mathbb{P}_{m}\left(\mathbf{x}_{s}\,\middle|\,f_{s-1},z_{s},a_{s}\right)$
$=p_{\text{init}}\prod_{s=2}^{t}P_{Z}\left(z_{s}\right)P_{A_{s}|Z_{s},F_{s-1}}\left(a_{s}\,\middle|\,z_{s},f_{s-1}\right)\prod_{s=1}^{t}P^{a_{s},z_{s}}_{m}\left(\mathbf{x}_{s}\right)$ (44)
where $p_{\text{init}}\triangleq P_{Z}(z_{1})P_{A_{1}|Z_{z}}(a_{1}\,|\,z_{1})$, $\mathbf{x}_{t}\triangleq\{x_{t,j}\,|\,j\in a_{t}\cap z_{t}\}$, and by convention, $\mathsf{P}_{m}^{(0)}\left(f_{0}\right)\triangleq 0$. In that, (43) holds true since $Z_{t}$ and $F_{t-1}$ are independent $\forall\,t\geq 1$, and (44) is due to the conditional independence of samples given the current action across time slots. Note that if there is no other definition in the context, we use upper and lower case letters to denote random variables and the values of random variables, respectively.
Hence, we have the expression of $S_{t,m,\theta}$, $\forall\,m,\theta\in\left[M\right]$, $m\neq\theta$, $\forall\,t\geq 1$,
$S_{t,m,\theta}=\log{\frac{\mathsf{P}_{m}^{(t)}(F_{t})}{\mathsf{P}_{\theta}^{(t)}(F_{t})}}=\sum_{s=1}^{t}\log{\frac{P^{A_{s},Z_{s}}_{m}(\mathbf{X}_{s})}{P^{A_{s},Z_{s}}_{\theta}(\mathbf{X}_{s})}}$
where $S_{0,m,\theta}\triangleq 0$ by convention.
Then, we are ready to show the martingale property. To begin with, $\forall\,t\geq 0$, by the expression of $S_{t,m,\theta}$, we have
$\mathbb{E}_{\theta}\left[e^{S_{t+1,m,\theta}}\,\middle|\,\mathcal{F}_{t}\right]=e^{S_{t,m,\theta}}\mathbb{E}_{\theta}\left[e^{\log{\frac{P^{A_{t+1},Z_{t+1}}_{m}(\mathbf{X}_{t+1})}{P^{A_{t+1},Z_{t+1}}_{\theta}(\mathbf{X}_{t+1})}}}\,\middle|\,\mathcal{F}_{t}\right]$
where $\left\{e^{S_{t,m,\theta}}\,\middle|\,t\geq 0\right\}$ is adapted to the natural filtration $\left\{\mathcal{F}_{t}\right\}_{t\geq 0}$. Then, the goal is turned to show for any $t\geq 0$,
$\mathbb{E}_{\theta}\left[e^{\log{\frac{P^{A_{t+1},Z_{t+1}}_{m}(\mathbf{X}_{t+1})}{P^{A_{t+1},Z_{t+1}}_{\theta}(\mathbf{X}_{t+1})}}}\,\middle|\,\mathcal{F}_{t}\right]=1$
for showing that $\left\{e^{S_{t,m,\theta}}\,\middle|\,t\geq 0\right\}$ is a martingale process under hypothesis $\theta$.
Since we define the distribution of action taking at time slot $t+1$ is a function of $F_{t}$ and $Z_{t+1}$ (see Definition 1), we simply switch from conditioning on $\mathcal{F}_{t}$ to merely conditioning on $F_{t}$ without loss of generality in the following steps. So, we have
$\mathbb{E}_{\theta}\left[e^{\log{\frac{P^{A_{t+1},Z_{t+1}}_{m}(\mathbf{X}_{t+1})}{P^{A_{t+1},Z_{t+1}}_{\theta}(\mathbf{X}_{t+1})}}}\,\middle|\,\mathcal{F}_{t}\right]$ $=\sum_{a\in\mathcal{A}}\sum_{z\in\mathcal{Z}}P_{A_{t+1},Z_{t+1}|F_{t}}\left(a,z\,\middle|\,f_{t}\right)\mathbb{E}_{\theta}\left[e^{\log{\frac{P^{a,z}_{m}(\mathbf{X}_{t+1})}{P^{a,z}_{\theta}(\mathbf{X}_{t+1})}}}\,\middle|\,A_{t+1}=a,Z_{t+1}=z,F_{t}=f_{t}\right]$
$=\sum_{a\in\mathcal{A}}\sum_{z\in\mathcal{Z}}P_{A_{t+1},Z_{t+1}|F_{t}}\left(a,z\,\middle|\,f_{t}\right)\mathbb{E}_{\theta}\left[\frac{P^{a,z}_{m}(\mathbf{X}_{t+1})}{P^{a,z}_{\theta}(\mathbf{X}_{t+1})}\,\middle|\,A_{t+1}=a,Z_{t+1}=z,F_{t}=f_{t}\right].$
Subsequently, since the random sample set $\mathbf{X}_{t+1}$ follows the joint distribution $P^{a,z}_{\theta}$ under a given $A_{t+1}=a$ and $Z_{z+1}=z$, and the samples are conditionally independent given the current action across time slots, we have
$\mathbb{E}_{\theta}\left[\frac{P^{a,z}_{m}(\mathbf{X}_{t+1})}{P^{a,z}_{\theta}(\mathbf{X}_{t+1})}\,\middle|\,A_{t+1}=a,Z_{t+1}=z,F_{t}=f_{t}\right]=1$
by Assumption 1.
To complete the proof, we can conclude
$\mathbb{E}_{\theta}\left[e^{\log{\frac{P^{A_{t+1},Z_{t+1}}_{m}(\mathbf{X}_{t+1})}{P^{A_{t+1},Z_{t+1}}_{\theta}(\mathbf{X}_{t+1})}}}\,\middle|\,\mathcal{F}_{t}\right]$ $=\sum_{a\in\mathcal{A}}\sum_{z\in\mathcal{Z}}P_{A_{t+1},Z_{t+1}|F_{t}}\left(a,z\,\middle|\,f_{t}\right)$
$=\sum_{z\in\mathcal{Z}}\alpha_{z}\sum_{a\in\mathcal{A}}P_{A_{t+1}|Z_{t+1},F_{t}}\left(a\,\middle|\,z,f_{t}\right)$
$=1$
where the last equality holds by Definition 1 and the stochastic model of temporally available data sources in Section II-A. ∎