### v2

# Tradeoffs among Action Taking Policies Matter in Active Sequential Multi-Hypothesis Testing: the Optimal Error Exponent Region

Abstract
Reliability of sequential hypothesis testing can be greatly improved when decision maker is given the freedom to adaptively take an action that determines the distribution of the current collected sample. Such advantage of sampling adaptivity has been realized since Chernoff’s seminal paper in 1959 [1]. While a large body of works have explored and investigated the gain of adaptivity, in the general multiple-hypothesis setting, the fundamental limits of individual error probabilities have not been fully understood. In particular, in the asymptotic regime as the expected stopping time tends to infinity, the error exponents are only characterized in specific cases, such as that of the total error probability. In this paper, we consider a general setup of active sequential multiple-hypothesis testing where at each time slot, a temporally varying subset of data sources (out of a known set) emerges from which the decision maker can select to collect samples, subject to a family of expected selection budget constraints. The selection of sources, understood as the “action” at each time slot, is constrained in a predefined action space. At the end of each time slot, the decision maker either decides to make the inference on the $M$ hypotheses, or continues to observe the data sources for the next time slot. The optimal tradeoffs among $M(M-1)$ types of error exponents are characterized. A companion asymptotically optimal test that strikes the balance between exploration and exploitation is proposed to achieve any target error exponents within the region. To the best of our knowledge, this is the first time in the literature to identify such tradeoffs among error exponents, and it uncovers the tension among different action taking policies even in the basic setting of Chernoff [1].
I Introduction
Adaptivity in taking samples from different distributions is known to greatly enhance the performance of sequential statistical inference. The intuition is that different distributions have different levels of informativeness of the inference target, and the most informative one may vary with the ground truth. The decision maker builds up more confidence towards the ground truth as it collects more samples, so it should choose the most informative distribution at the moment accordingly. As for hypothesis testing, the paradigm is called active sequential hypothesis testing [1, 2, 3], where the decision maker has the freedom to adaptively take an action that determines the distribution of the sample on the fly. It can be understood as the setting where the decision maker selects a particular data source out of multiple ones to collect a sample, each of which corresponds to a possible action. This paradigm finds extensive application in various fields such as spectrum sensing, crowdsourced data analytics, etc., and motivated from these applications, there have been extensions that incorporate a variety of expected selection budge constraints [4, 5, 6, 7] and temporally changing available data sources [8].
In existing works about active sequential multi-hypothesis testing, however, the fundamental limits of individual error probabilities have not been fully understood. In the related literature [1, 2, 3, 4, 6, 5, 7], the main focus is on minimizing various costs related to the expected stopping time subject to constraints such as a constant total error probability [2], a constant maximal error probability [1, 7], constant type-I and type-II error probabilities [5, 6, 4], and constant decision risks [3]. When the treatment pertains to the general non-asymptotic regime, it is usually difficult to characterize the optimal performance analytically, and hence the goal becomes developing efficient algorithms for finding the optimal selection policy numerically under a given framework. As for the asymptotic regime, one may turn the problems in [1, 2, 3, 4, 6, 5, 7] around and study the optimal error probabilities subject to expected stopping time and budget constraints. It is then not difficult to characterize the exponentially decaying rates (error exponents) in the respective settings. However, in the general $M$-ary multi-hypothesis setting, there are in total $M(M-1)$ types of error probabilities, and it remains unclear how the asymptotically optimal policies and inference strategies change when trading off one error probability with others.
Towards finer resolution of the fundamental performance limits in active sequential $M$-ary multi-hypothesis testing problems, in this work we focus on $M(M-1)$ types of individual error probabilities in a general framework that covers those in the literature [1, 2, 3, 4, 6, 5, 7, 8] as special cases. In this framework, out of $n$ data sources in total, at each time slot a random subset of them is revealed to the decision maker as the available sources. The distribution of this random subset is i.i.d. over time. The decision maker then takes an action (may be randomized) to select data sources from this available subset and collects samples from the selected ones. At the end of each time slot, according to the samples accumulated so far, the decision maker either determines to make an inference among the $M$ hypotheses or advances to the next time slot. Under a family of expected selection budget constraints including the expected stopping time, the goal is to study the asymptote of individual error probabilities, in particular the exponentially decaying rates (error exponents), as all budget constraints scale to infinity with fixed ratios.
Our main contribution is the characterization of the optimal error exponent region. We show that the achievable error exponents can be represented as linear combinations of Kullback-Leibler (KL) divergences, each of which corresponds to a particular set of data sources’ joint level of informativeness for distinguishing one hypothesis from another. The coefficients in the linear combinations correspond to the expected selection frequencies of data sources under the given policy and turn out to be the key design choices. In our characterization, these selection frequencies are indexed naturally by the taken action and the available subset of data sources. The optimal error exponent region is the union of these $M(M-1)$-tuples of error exponents over all possible such selection frequencies that meet the normalized expected budget constraints (i.e., the ratios to the expected stopping time) and the distribution of the random available subset. Leveraging the fact that these constraints are all affine, we further note that the optimal error exponent region is a convex set. The characterized region turns out to be a direct product of $M$ sub-regions each of which is in an $(M-1)$-dimensional space, and a tradeoff only appears among the exponents that correspond to the same declared hypothesis against all other $M-1$ possible ground truth hypotheses. To the best of our knowledge, such tradeoffs among $M(M-1)$ types of error exponents have not been identified before, and it uncovers the tension among different action taking policies even in the basic setting of Chernoff [1].
To characterize the optimal error exponent region, we divide the proof of our main theorem into two parts: the converse part and the achievability part. The converse part of our main theorem is proved by standard arguments that involves data processing inequality and Doob’s Optional Stopping Theorem, similar to that in [8]. The more interesting part is achievability. The stopping and inference rules of our test mainly follow a family of multi-hypothesis sequential probability ratio tests (MSPRT) proposed in [9], with thresholds designed judiciously to achieve the target error exponents and meet the expected stopping time constraint. For the action taking policy, since the constraints are all in expectation, it is natural to use a randomized one, and hence the design boils down to choosing the action taking probabilites. To achieve a given tuple of error exponents in the region, the coefficients in the linear combinations of the KL divergences serve as the guidance for setting these action taking probabilities in the randomized policy. As suggested in the literature [1, 2, 3, 8], one way is to exploit the accumulated information and choose the action taking probabilities derived from the linear combinations that correspond to the hypothesis with the maximum likelihood. It turns out that similar to [1], such a “pure-exploitation” approach requires stronger assumptions on the data sources so that it can be shown to be asymptotically optimal. Inspired by [7], in our proposed policy, on top of the “pure-exploitation” part, we further mix it with a carefully designed probability for “exploration” so that more relaxed assumptions can be set.
The conference version of this work [10] considered a restricted setup where there are just two simple expected budget constraints: individual and total budget constraints, and the collected samples at each time slot are assumed to be conditionally independent random variables given the current action. In contrast, in this work, we consider a family of expected selection budget constraints, namely linear expected budget constraints in the remaining context, and relax the conditional-independence assumption.
The remainder of this paper is organized as follows. In Section II, we give the problem formulation, including the statistical model of collected samples and available subset of data sources, the family of active sequential tests, and the expected budget constraints. Our main theorem is presented in Section III along with an illustrative example scenario in the basic setting of Chernoff [1], demonstrating the tradeoffs among error exponents. The proposed asymptotically optimal test along with the achievability proof of our main theorem are given in Section IV. The details of the converse proof are provided in Section V. Section VI details the proof of a technical lemma in the achievability part. Finally, in Section VII, we conclude this paper.
II Problem Formulation
There are $M$ hypotheses and $n$ data sources labeled by integers in $\left[M\right]\triangleq\{0,1,2,\ldots,M-1\}$ and $\left[n\right]\triangleq\{1,2,\ldots,n\}$, respectively, where $2\leq M<\infty$, $1\leq n<\infty$. If not specified, in this paper upper and lower case letters are used to denote random variables and the values of random variables, respectively.
II-A Statistical model
The statistical model of the general framework considered in this work comprises two parts: the available subset of data sources and the samples drawn from the data sources.
At each time instance $t\geq 1$, the temporally available subset of data source $Z_{t}$ (a random variable), following a probability mass function (pmf) $P_{Z}$ (known to the decision maker) in an i.i.d. manner over time, is revealed to the decision maker. Let $\mathcal{Z}=\{z_{1},z_{2},\ldots,z_{\lvert\mathcal{Z}\rvert}\}$ denote the support of $P_{Z}$, $\lvert\mathcal{Z}\rvert<\infty$. Each $z\in\mathcal{Z}$ consists of a subset of data sources, i.e., $z_{i}\subseteq\left[n\right]\ \forall\,i\in\{1,2,\ldots,\lvert\mathcal{Z}\rvert\}$.
As for the collected samples across time slots, under hypothesis $\theta\in\left[M\right]$, samples generated by $n$ data sources (if they are selected) follow a joint distribution $P_{X_{1},X_{2},\ldots,X_{n};\theta}$ where $X_{j}$ denotes the random variable sampled from the source $j\in\left[n\right]$. Moreover, samples drawn from a subset of data sources (if selected) follow the corresponding distribution marginalized from $P_{X_{1},X_{2},\ldots,X_{n};\theta}$. For example, under hypothesis $\theta\in\left[M\right]$, if $\{1,2\}\subseteq\left[n\right]$ is the selected subset of data sources, then samples drawn from these two sources follow the distribution $P_{X_{1},X_{2};\theta}$ marginalized from $P_{X_{1},X_{2},\ldots,X_{n};\theta}$. For notational convenience, $\mathbb{P}_{\theta}\left(\cdot\right)$ and $\mathbb{E}_{\theta}[\cdot]$ respectively denote the short-hand notations for “the probability” and “the expectation” given the ground truth hypothesis is $\theta\in\left[M\right]$.
II-B Action taking, stopping time, and inference
At each time instance $t\geq 1$, after the available subset $Z_{t}$ is revealed, the decision maker first takes an action denoted as $A_{t}$ drawn from an action space $\mathcal{A}\triangleq\{a_{1},a_{2},\ldots,a_{\lvert\mathcal{A}\rvert}\}$, where $\lvert\mathcal{A}\rvert<\infty$ and each $a\in\mathcal{A}$ is a subset of data sources, i.e., $a_{i}\subseteq\left[n\right]\ \forall\,i\in\{1,2,\ldots,\lvert\mathcal{A}\rvert\}$. Furthermore, the decision maker is allowed to skip any action, i.e., $\emptyset\in\mathcal{A}$, indicating the option to not select any sources at a time slot. The action taking process can be randomized, that is, $A_{t}\in\mathcal{A}$ is a random variable, following a distribution determined by the decision maker. This distribution is the action taking policy and is adaptive in general, that is, at each time instance $t\geq 1$, the probability of taking $A_{t}\in\mathcal{A}$ is a function of the data (comprising samples, actions, and available sets) observed until time slot $t-1$.
After taking an action at each time slot $t$, the decision maker collects a set of samples $\{X_{t,j}\,|\,j\in A_{t}\cap Z_{t}\}$ following the corresponding marginal distribution, where each $X_{t,j}$ is the sample generated by the data source $j$ that is selected, i.e., in the intersection of the action $A_{t}$ and the temporary available set $Z_{t}$. In short, $\forall\,t\geq 1$, given $A_{t}=a\in\mathcal{A}$ and $Z_{t}=z\in\mathcal{Z}$, the collected samples are represented as the random tuple $\mathbf{X}_{t}\triangleq\{X_{t,j}\,|\,j\in a\cap z\}$ following the distribution marginalized from $P_{X_{1},X_{2},\ldots,X_{n};\theta}$, $P_{\{X_{j}|j\in a\cap z\};\theta}$, under hypothesis $\theta\in\left[M\right]$, $\forall\,a\in\mathcal{A}$, $\forall\,z\in\mathcal{Z}$. For notational convenience, we denote
$P^{a,z}_{\theta}\triangleq P_{\{X_{j}|j\in a\cap z\};\theta}.$
We assume given the actions, the random sample set $\mathbf{X}_{t}$ is conditionally independent across time slots. After receiving the set $\mathbf{X}_{t}$ at time slot $t\geq 1$, the decision maker either stops and makes an inference or proceeds to the next time slot. The time that the decision maker stops is the stopping time, denoted as $\tau\in\mathbb{N}$, and the inferred result is denoted as $\phi\in\left[M\right]$.
The procedures at the decision maker is summarized in the definition below.
Definition 1 (Active sequential test)
An active sequential test in the proposed framework comprises the following two parts: at each time instance $t\geq 1$,
  • The distribution that the randomized action $A_{t}$ follows is a function of $F_{t-1}$ and $Z_{t}$, where
    $F_{t}\triangleq\{\{\mathbf{X}_{s}\}_{s=1}^{t},\{A_{s}\}_{s=1}^{t},\{Z_{s}\}_{s=1}^{t}\}\quad\forall\,t\geq 1.$
    By convention, $F_{0}=\emptyset$.
  • $\tau$ and $\phi$ are both functions of $F_{t}$, $\forall\,t\geq 1$.
As for the performance of the overall decision making, this work focuses on the individual error probabilities: $\forall\,\theta\in\left[M\right]\setminus\{m\},\forall\,m\in\left[M\right]$,
$\pi_{m|\theta}\triangleq\mathbb{P}_{\theta}\left(\phi=m\right),$ (1)
the probability of mistaking $\theta$ for $m$. Note that there are in total $M(M-1)$ individual error probabilities.
II-C Budget constraints and error exponents
For describing the budget constraints, we first define a set $\mathcal{C}_{B}$ as follows:
$\mathcal{C}_{B}\triangleq\left\{\mathbf{x}\triangleq\left\{x_{1},x_{2},\ldots,x_{n}\right\}\in\mathbb{R}^{n}_{+}\,\middle|\,b_{i}(\mathbf{x})\leq U_{i}\ \forall\,i\in\mathcal{I}\right\},$ (2)
where $\mathbb{R}_{+}$ denotes the set of non-negative reals, $\mathcal{I}$ is an index set, each $b_{i}(\cdot)$ is a linear function with bounded coefficients, $b_{i}:\mathbb{R}^{n}_{+}\rightarrow\mathbb{R}_{+}$, satisfying $b_{i}(\mathbf{0})=0$, and each $U_{i}$ is a positive real number, representing the budget. Note that $\mathcal{I}$ is allowed to be countable or uncountable and it is given to the decision maker along with $b_{i}(\cdot)$ and $U_{i}\ \forall\,i\in\mathcal{I}$. Then, the budget constraints including expected stopping time constraints are given below: for some $T\geq 1$ and a given set $\mathcal{C}_{B}$:
$\mathbb{E}_{\theta}\left[\tau\right]\leq T\quad\text{and}\quad\mathbf{B}^{\theta}\triangleq\left\{B^{\theta}_{1},B^{\theta}_{2},\ldots,B^{\theta}_{n}\right\}\in\mathcal{C}_{B}\quad\forall\,\theta\in\left[M\right],$ (3)
where $\mathcal{C}_{B}$ is defined in (2) and each $B^{\theta}_{j}\triangleq\mathbb{E}_{\theta}\left[\sum_{t=1}^{\tau}\mathbbm{1}{\left\{j\in A_{t}\cap Z_{t}\right\}}\right]$ denotes the expected total number of selecting source $j\in\left[n\right]$ under hypothesis $\theta\in\left[M\right]$.
Since in the definition of $\mathcal{C}_{B}$, $b_{i}$’s are linear functions, the constraints in (3) are also called linear expected budget constraints. Note that the budget constraints considered in the literature [5, 6, 8] are special cases of (3). In [5, 6], individual expected budget constraints were introduced, where =$\mathcal{I}=\left\{1,\ldots,n\right\}$, $b_{i}(\mathbf{x})=x_{i}$, $\forall\,i\in\mathcal{I}$, $\forall\,\mathbf{x}\in\mathbb{R}^{n}_{+}$. [8] considered total expected budget constraints, where $|\mathcal{I}|=1$ and $b_{i}(\mathbf{x})=\sum_{j=1}^{n}x_{j},i\in\mathcal{I}$, $\forall\,\mathbf{x}\in\mathbb{R}^{n}_{+}$.
Our focus is on the fundamental limits of the performance in an asymptotic regime as $U_{i}\ \forall\,i\in\mathcal{I}$ and $T$ all tend to infinity with fixed ratios
$r_{i}\triangleq\frac{U_{i}}{T}>0\quad\forall\,i\in\mathcal{I}.$
The achievable individual error exponent $e_{m|\theta},\forall\,\theta\in\left[M\right]\setminus\{m\},\forall\,m\in\left[M\right]$ is defined as the follows.
Definition 2 (Error exponents)
A given $M(M-1)$-tuple
$\Big{\{}e_{m|\theta}\,\Big{|}\,m\in\left[M\right],\theta\in\left[M\right]\setminus\{m\}\Big{\}}$ (4)
is achievable if there exists an active sequential test (comprising action taking, stopping time, and inference) satisfying the constraints (3) such that each of the $M(M-1)$ individual error probabilities vanishes and satisfies
$\liminf_{T\to\infty}\frac{1}{T}\log\frac{1}{\pi_{m|\theta}}\geq e_{m|\theta}\quad\forall\,\theta\in\left[M\right]\setminus\{m\},\forall\,m\in\left[M\right].$ (5)
Our goal is to characterize the optimal error exponent region, the collection of all achievable tuples.
III Main Results
In this section, we first establish the optimal region of the achievable individual error exponents. Then we specialize our result under Chernoff’s setting [1, 2, 3] to clearly demonstrate that there may exist tradeoffs among the individual error exponents. Finally, we show that there is no tradeoff in the asymptotic regime when it comes to the performance metrics considered in prior works [1, 2, 3], such as the decision risks.
Notations
×$\bigtimes_{m=0}^{M-1}A_{m}$ denotes the direct product over sets, i.e., $A_{0}\times A_{2}\times...\times A_{M-1}$. $\text{Conv}(A)$ and $\text{extr}(A)$ denote as the convex hull of set $A$ and the collection of extreme points of set $A$, respectively. $\mathbb{R}_{+}$ denotes the set of non-negative reals. When the context is clear, we drop the indices and use $\{e_{m|\theta}\}$ to denote the $M(M-1)$-tuple in (4). $\{x_{j}\}_{j=1}^{k}$ denotes the $k$-tuple $\{x_{1},x_{2},...,x_{k}\}$ where $x_{j}$ is any value indexed by $j\in\{1,2,...,k\}$ for any integer $k\geq 2$.
III-A Main theorem
For the characterization of the optimal error exponents, the following two technical assumptions are imposed.
Assumption 1
There exists $L<\infty$ such that
$\sup_{m\in\left[M\right]}\sup_{\theta\in\left[M\right]\setminus\{m\}}\sup_{\mathbf{x}\in\mathcal{X}_{\theta}}\left\lvert\log\frac{P_{X_{1},\ldots,X_{n};\theta}(\mathbf{x})}{P_{X_{1},\ldots,X_{n};m}(\mathbf{x})}\right\rvert\leq L.$ (6)
where $\log\frac{0}{0}=0$ by convention and $\mathcal{X}_{\theta}$ denotes the support of the joint distribution $P_{X_{1},\ldots,X_{n};\theta}\quad\forall\,\theta\in\left[M\right]$.
Assumption 2
For any two hypotheses $m,\theta\in\left[M\right]$, $\theta\neq m$, $\exists z\in\mathcal{Z}$, $\exists a\in\mathcal{A}$, such that ${D}\left(P^{a,z}_{m}\,\middle\|P^{a,z}_{\theta}\right)>0$.
Assumption 1 is used in our achievability and converse proof and commonly made in the literature [2, 1, 3, 7, 8]. Assumption 2 is needed for proving the achievability, and further discussion will be made in Section IV. Roughly speaking, it implies that for any two different hypotheses, there exists a feasible set of data sources that can distinguish one from the other and vice versa.
We are ready to state our main result.
Theorem 1 (The optimal error exponent region)
Suppose both Assumption 1 and 2 hold. Let
$\bm{\beta}^{m}=\Big{\{}\beta^{m}_{a,z}\,\Big{|}\,a\in\mathcal{A},z\in\mathcal{Z}\Big{\}}\in\mathbb{R}^{|\mathcal{A}||\mathcal{Z}|}_{+}$
denote a tuple of coefficients, indexed by the hypothesis $m$, the action $a$, and the available subset $z$. Define
$\zeta\triangleq\bigtimes\limits_{m=0}^{M-1}\,\left\{\bigcup\limits_{\bm{\beta}^{m}\in\mathcal{C}}\zeta_{m}(\bm{\beta}^{m})\right\}$ (7)
where
$\mathcal{C}$$\triangleq\left\{\bm{\beta}\in\mathbb{R}^{|\mathcal{A}||\mathcal{Z}|}_{+}\,\middle|\,b_{i}\left(\bm{\omega}(\bm{\beta})\right)\leq r_{i},\,\sum_{a\in\mathcal{A}}\beta_{a,z}\leq\alpha_{z}\triangleq P_{Z}(z),\ \forall\,i\in\mathcal{I},\ \forall\,z\in\mathcal{Z}\right\},$
$\bm{\omega}(\bm{\beta})$$\triangleq\left\{\omega_{1}(\bm{\beta}),\ldots,\omega_{n}(\bm{\beta})\right\},$
$\omega_{j}(\bm{\beta})$ $\triangleq\sum_{a\in\mathcal{A}}\sum_{z\in\mathcal{Z}}\beta_{a,z}\mathbbm{1}{\left\{j\in a\cap z\right\}}\quad\forall\,j\in\left[n\right],$
and for each $m\in\left[M\right]$,
$\zeta_{m}(\bm{\beta}^{m})$$\triangleq\left\{\big{\{}e_{m|\theta}\,\big{|}\,\theta\in\left[M\right]\setminus\{m\}\big{\}}\in\mathbb{R}_{+}^{M-1}\,\middle|\,e_{m|\theta}\leq e^{*}_{m|\theta}(\bm{\beta}^{m}),\,\forall\,\theta\in\left[M\right]\setminus\{m\}\right\},$
$e^{*}_{m|\theta}(\bm{\beta}^{m})$ $\triangleq\sum_{a\in\mathcal{A}}\sum_{z\in\mathcal{Z}}\beta^{m}_{a,z}{D}\left(P^{a,z}_{m}\,\middle\|P^{a,z}_{\theta}\right)\quad\forall\,m\in\left[M\right],\,\forall\,\theta\in\left[M\right]\setminus\{m\}.$
Then, any $M(M-1)$-tuple $\{e_{m|\theta}\}$ inside $\zeta$ can be achieved by our proposed test with a corresponding adaptive selection policy. Conversely, an achievable $M(M-1)$-tuple $\{e_{m|\theta}\}$ must lie inside the region $\zeta$.
Proof:
The proposed test and the achievability proof are provided in Section IV. The proof of the converse part of Theorem 1 primarily relies on the data processing inequality and Doob’s Optional Stopping Theorem, similar to that in [8]. Details can be found in Section V.∎
Before we proceed, let us make a few remarks based on the characterization in our main theorem.
Remark 1 (Tradeoffs among error exponents)
Theorem 1 characterizes the optimal error exponent region as a direct product of $M$ components as in (7). It can be immediately seen that the tradeoff only appears among the $(M-1)$-tuples in the component $\bigcup_{\bm{\beta}^{m}\in\mathcal{C}}\zeta_{m}(\bm{\beta}^{m})$ for each $m\in\left[M\right]$. In words, in each region $\zeta_{m}(\bm{\beta}^{m})$, one may not find an optimal $\bm{\beta}^{m}\in\mathcal{C}$ to maximize all bounds $\left\{e^{*}_{m|\theta}(\bm{\beta}^{m})\,\middle|\,\theta\in\left[M\right]\setminus\{m\}\right\}$ simultaneously. Later in Section III-B, the tradeoff is illustrated in a simple yet classical setting originally considered by Chernoff [1].
Remark 2 (No tradeoff when $M=2$)
[…]
Remark 3 (Convexity of the optimal region $\zeta$)
The convexity of the region $\zeta$ is a straightforward result due to the direct product form in $\zeta$ and the convexity of the region $\bigcup\limits_{\bm{\beta}^{m}\in\mathcal{C}}\zeta_{m}(\bm{\beta}^{m})$, $\forall\,m\in\left[M\right]$. The later property is obtained by the definition of convexity and the fact that $\mathcal{C}$ is a convex set for any index set $\mathcal{I}$ and each $\zeta_{m}(\bm{\beta}^{m})$ is a hyperrectangle where each of its cardinalities is upper bounded by a linear mapping of $\bm{\beta}^{m}$, i.e., $e^{*}_{m|\theta}(\bm{\beta}^{m})$ for each $m\in\left[M\right]$, $\theta\in\left[M\right]\setminus\{m\}$. Moreover, in the case of $|\mathcal{I}|<\infty$, $\mathcal{C}$ is composed of finitely many intersections of linear inequalities and is bounded, leading to the convex polytope representation. Therefore, for each $m\in\left[M\right]$, after including the property of $\zeta_{m}(\bm{\beta}^{m})$ mentioned above, there are only finitely many extreme points of the set $\bigcup\limits_{\bm{\beta}^{m}\in\mathcal{C}}\zeta_{m}(\bm{\beta}^{m})$, namely, a convex polytope. Note that the definition of extreme points is the common one defined on a convex set. Consequently, the optimal region $\zeta$ is also a convex polytope due to its direct product form.
III-B Specialization
To highlight the tradeoffs among individual error exponents, we specialize Theorem 1 under Chernoff’s setting [1, 2, 3], where all data sources are available at each time instance ($\mathcal{Z}=\left\{\left[n\right]\right\}$), but only one data source can be chosen per time ($\mathcal{A}=\left\{\{1\},\{2\},...\{n\}\right\}$) without the budget constraints ($\mathcal{C}_{B}=\mathbb{R}^{n}_{+}$). The following corollary summarizes the error exponent region in this case.
Corollary 1 (The optimal error exponent region in Chernoff’s setting)
The optimal error exponent region for individual error probabilities in Chernoff’s setting [1] is given by
$\zeta=\bigtimes\limits_{m=0}^{M-1}\,\left\{\bigcup\limits_{\bm{\beta}^{m}\in\mathcal{P}^{n}}\zeta_{m}(\bm{\beta}^{m})\right\}$ (8)
where $\mathcal{P}^{n}$ is the $n$-dimensional probability simplex, and for each $m\in\left[M\right]$, $\zeta_{m}(\bm{\beta}^{m})$ is defined as in Theorem 1 with the following specialization:
$e^{*}_{m|\theta}(\bm{\beta}^{m})=\sum_{j=1}^{n}\beta^{m}_{j}{D}\left(P_{m,j}\,\middle\|P_{\theta,j}\right)\quad\forall\,\theta\in\left[M\right]\setminus\{m\},$
$\beta^{m}_{j}\triangleq\beta^{m}_{j,z}\quad\forall\,j\in\mathcal{A}=\left\{\{1\},\{2\},...\{n\}\right\},\,z\in\mathcal{Z}=\left\{\left[n\right]\right\},$
$P_{\theta,j}(\cdot)\triangleq P_{X_{j};\theta}(\cdot)\quad\forall\,j\in\left[n\right],\,\theta\in\left[M\right].$
To illustrate the tradeoffs, let us further consider a special case when there are just three hypotheses and two data sources in Corollary 1. In this case, for any $\theta,m\in[3]$, $\theta\neq m$, the bound $e^{*}_{m|\theta}(\bm{\beta}^{m})$ is simplified as
$e^{*}_{m|\theta}(\bm{\beta}^{m})=\beta^{m}_{1}{D}\left(P_{m,1}\,\middle\|P_{\theta,1}\right)+\beta^{m}_{2}{D}\left(P_{m,2}\,\middle\|P_{\theta,2}\right).$
Now let us focus on a particular $m$, say $m=0$, since tradeoffs only happen within each component in the direct product in (7). Specifically, let us look at the tradeoff between $e^{*}_{0|1}(\bm{\beta}^{0})$ and $e^{*}_{0|2}(\bm{\beta}^{0})$, where
$e^{*}_{0|1}(\bm{\beta}^{0})=\beta^{0}_{1}{D}\left(P_{0,1}\,\middle\|P_{1,1}\right)+\beta^{0}_{2}{D}\left(P_{0,2}\,\middle\|P_{1,2}\right),$
$e^{*}_{0|2}(\bm{\beta}^{0})=\beta^{0}_{1}{D}\left(P_{0,1}\,\middle\|P_{2,1}\right)+\beta^{0}_{2}{D}\left(P_{0,2}\,\middle\|P_{2,2}\right).$
With $\bm{\beta}^{0}\in\mathcal{P}^{2}$, i.e., $\beta^{0}_{1}+\beta^{0}_{2}=1$ and $\beta^{0}_{1},\beta^{0}_{2}\geq 0$, the tradeoff emerges when each source is the most informative in distinguishing hypothesis 00 from exactly one hypothesis, for example,
${D}\left(P_{0,1}\,\middle\|P_{1,1}\right)>{D}\left(P_{0,2}\,\middle\|P_{1,2}\right)$
but ${D}\left(P_{0,1}\,\middle\|P_{2,1}\right)<{D}\left(P_{0,2}\,\middle\|P_{2,2}\right)$
As a result, $e^{*}_{0|1}(\bm{\beta}^{0})$ and $e^{*}_{0|2}(\bm{\beta}^{0})$ cannot be maximized simultaneously, resulting the tradeoff. Conversely, if one of the sources (say source $1$) is the most informative in distinguishing hypothesis 00 from both $1$ and $2$, one can set $\beta^{0}_{1}=1$ to maximize $e^{*}_{0|1}(\bm{\beta}^{0})$ and $e^{*}_{0|2}(\bm{\beta}^{0})$ simultaneously.
From the above illustrative example, it can be seen that in Chernoff’s setting, the tradeoffs among the individual error exponents are due to the heterogeneous level of informativeness across data sources in distinguishing a declared hypothesis against other ground truth hypotheses. This very insight carries to the general case considered in our main result. As for why the tradeoff only appears when $M\geq 3$, from a high-level viewpoint, for each $m\in\left[M\right]$, $\bm{\beta}^{m}$ for $\zeta_{m}(\bm{\beta}^{m})$ can be interpreted as a “guidance of selection” for distinguishing hypotheses, and when $M\geq 3$, $\bm{\beta}^{m}$ is responsible for distinguishing hypothesis $m$ from more than just one hypotheses (i.e., $\lvert\left[M\right]\setminus\{m\}\rvert\geq 2$), making it challenging to find a guidance simultaneously optimal under the heterogeneous distinguishability among data sources. However, when $M=2$, each $\bm{\beta}^{m}$ for $\zeta_{m}(\bm{\beta}^{m})$ just needs to take care of distinguishing one hypothesis from the other, and hence there is no tension in picking the guidance of selection.
III-C Performance metrics without tradeoff
In the literature of active sequential hypothesis testing, tradeoffs in the asymptotic regime have not been observed. The reason is that prior studies have not delved into individual error probabilities for $M\geq 3$. For example, [1, 3, 2] focused on a coarser performance metric, the decision risk of declaring the decision as hypothesis $m$
$\mathrm{R}_{m}\triangleq\sum_{\theta\in\left[M\right]\setminus\{m\}}\rho_{\theta}\pi_{m|\theta}\quad\forall\,m\in\left[M\right].$
Here $\rho_{\theta}$ is the prior probability for hypothesis $\theta\in\left[M\right]$.
Fundamental limits of the exponentially decaying rates of the decision risks, denoted as $\{\gamma_{m}\mid m\in\left[M\right]\}$, can then be characterized from our main result as a simple corollary.
Corollary 2 (Exponents of decision risks)
The optimal exponent region consists of non-negative $(\gamma_{1},...,\gamma_{M})$ satisfying
$\gamma_{m}\leq\max_{\bm{\beta}\in\mathcal{P}^{n}}\min_{\theta\in\left[M\right]\setminus\{m\}}e^{*}_{m|\theta}(\bm{\beta})\quad\forall\,m\in\left[M\right].$
Note that Corollary 2 can also be derived from the respective results in [1, 3, 2]. Clearly there are no tradeoffs among these $\gamma_{m}$’s, $m\in\left[M\right]$ due to the product form shown in (8).
Meanwhile, suppose we take the error probabilities of individual hypotheses
$\pi_{\theta}\triangleq\sum_{m\in\left[M\right]\setminus\{\theta\}}\pi_{m|\theta}\quad\forall\,\theta\in\left[M\right]$
as performance metrics. In contrast to the decision risks, the tradeoffs among the corresponding error exponents exist. The asymptote of error probabilities of individual hypotheses, however, was not studied in [1, 3, 2].
IV Proof of Achievability
In this section, we first provide an overview of the proposed asymptotically optimal test and discuss the necessity of “exploration” in the test. Then, the details of the proposed test are given, followed by the analysis of achievability.
IV-A Overview
For the achievability part of Theorem 1, from the expression in (7), the goal is to show that for any
$\bm{\beta}^{0},\bm{\beta}^{1},\ldots,\bm{\beta}^{M-1}\in\mathcal{C},$
the tuple of $M(M-1)$ individual error exponents $\Big{\{}e_{m|\theta}\,\Big{|}\,m\in\left[M\right],\theta\in\left[M\right]\setminus\{m\}\Big{\}}$ is achievable if
$\big{\{}e_{m|\theta}\,\big{|}\,\theta\in\left[M\right]\setminus\{m\}\big{\}}\in\zeta_{m}(\bm{\beta}^{m})\quad\forall\,m\in\left[M\right].$
Recall that each $\bm{\beta}^{m},m\in\left[M\right]$ comprises coefficients indexed by the action $a$ and the available subset $z$, hinting the probability of taking action $a$ when the available subset is $z$ under a randomized policy.
As for the strategy of taking actions, note that $\bm{\beta}^{\theta}$ (here we change the indexing convention from $m$ to $\theta$ for notational convenience) corresponds to the action taking policy under hypothesis $\theta$, and they are in general different across all $\theta\in\left[M\right]$. One strategy that exploits the information accumulated so far is to first find a maximum likelihood estimate (MLE) $\hat{\theta}_{\text{MLE}}\in\left[M\right]$ and employ the policy corresponding to $\bm{\beta}^{\hat{\theta}_{\text{MLE}}}$ [1]. In our design, in addition to exploitation, we further mix it with a small (vanishing-in-$T$) probability of exploration inspired by [7], which is a common idea in the literature of sequential decision making problems such as bandits. This probability of exploration and the rate of uniformly exploring all sources are carefully chosen to meet the constraints and achieve the target error exponents.
As for the stopping rule and the inference strategy, it follows those in the multiple-hypothesis sequential probability ratio tests (MSPRT) [9] with thresholds designed judiciously to achieve the target error exponents and meet the constraints. A key statistic here is the cumulative log-likelihood ratio
$S_{t,\theta,m}\triangleq\log\frac{\mathsf{P}_{\theta}^{(t)}(F_{t})}{\mathsf{P}_{m}^{(t)}(F_{t})},$
where $\mathsf{P}_{\theta}^{(t)}(\cdot)$ represents the probability distribution of $F_{t}$, the accumulated information up to the end of time instance $t$ (see Definition 1) under hypothesis $\theta$. The explicit form of $S_{t,\theta,m}$ is provided in Lemma 3 in the Appendix.
To this end, let us sketch the proposed test, with more details left in Section IV-B. For a given $(\bm{\beta}^{0},\bm{\beta}^{1},\ldots,\bm{\beta}^{M-1})$, at each time instance $t\geq 1$, upon $Z_{t}$ being revealed, the decision maker carries out the following:
  1. 1.
    Action taking policy and sample collection:
    At the end of time slot $t-1$, find the MLE $\hat{\theta}_{\text{MLE}}$, that is, $\theta$ such that $\min_{m\neq\theta}S_{t-1,\theta,m}\geq 0$. Based on this $\hat{\theta}_{\text{MLE}}$ and the corresponding $\bm{\beta}^{\hat{\theta}_{\text{MLE}}}$ together with a judiciously designed exploration rate, we then set up a conditional pmf $P_{A_{t}|Z_{t},F_{t-1}}(\cdot|z,f_{t-1})$ for randomized action taking. How to design this conditional pmf is elaborated later. Subsequently, a random action $a$ is drawn from the conditional pmf, and the decision maker collects a set of samples $\mathbf{X}_{t}$ accordingly.
  2. 2.
    Stopping and inference rule:
    If there exists a hypothesis $\theta$ such that $S_{t,\theta,m}$ exceeds a set of designed thresholds $\forall\,m\neq\theta$, then stop and declare such a $\theta$ as the inferred result.
As for the conditional pmf used for action taking, if hypothesis $\theta$ has the largest cumulative log-likelihood ratio $(\hat{\theta}_{\text{MLE}}=\theta)$ at the end of time slot $t-1$, then we set
$P_{A_{t}|Z_{t},F_{t-1}}(a|z,f_{t-1})=\begin{cases}(1-\epsilon(T))\frac{\beta^{\theta}_{a,z}}{\alpha_{z}}+\epsilon(T)\frac{1}{h\alpha_{z}}&\text{if }a\in\mathcal{A}\setminus\{\emptyset\}\\[4.30554pt] 1-\sum_{a^{\prime}\in\mathcal{A}\setminus\{\emptyset\}}P_{A_{t}|Z_{t},F_{t-1}}(a^{\prime}|z,f_{t-1})&\text{if }a=\emptyset\end{cases},$
where for notational convenience, $\alpha_{z}\triangleq P_{Z}(z)$, the probability of set $z$ being available, $\forall\,z\in\mathcal{Z}$. The probability of conducting exploration $\epsilon(T)$ is judiciously designed to attain the asymptotic optimality of the test, and it vanishes as $T\rightarrow\infty$. The exploration rate $\frac{1}{h}>0$ is chosen to be bounded from above as
$\frac{1}{h}\leq\min\left(\ \frac{\min_{z\in\mathcal{Z}}\alpha_{z}}{\lvert\mathcal{A}\rvert}\ ,\ \inf_{i\in\mathcal{I}}\frac{r_{i}}{b_{i}\left(\bm{\omega}(\mathbf{1})\right)}\ \right),$ (9)
where each $b_{i}(\cdot)$ is defined in (2) and $\mathbf{1}=\{1,\ldots,1\}\in\mathbb{R}^{|\mathcal{A}||\mathcal{Z}|}_{+}$. Such a choice is to satisfy the expected budget constraints and that the total probability of action taking sums up to be $1$.
Before we move on to the details of the proposed test and the analysis, we make the following remark on the exploration part in the action taking policy.
Remark 4 (On the necessity of exploration)
If we consider “pure-exploitation”, that is, $\epsilon(T)=0$, the result presented in Theorem 1 remains the same except that for the achievability part, a stronger assumption than Assumption 2 is needed, namely, $\forall\,z\in\mathcal{Z}$, $\forall\,a\in\mathcal{A}\setminus\{\emptyset\}$,
${D}\left(P^{a,z}_{\theta}\,\middle\|P^{a,z}_{m}\right)>0\quad\forall\theta,m\in\left[M\right],\theta\neq m.$
This stronger assumption can be traced back to Chernoff [1]. Some prior works [2, 3] also provide modified tests (with a cold-start phase for example in [2]) for the achievability under the weaker assumption (Assumption 2), which may serve as alternatives to our proposed scheme.
IV-B Details of the proposed test
Notations
The summation $\sum_{a,z}$ without its index set, is a short-hand notation for $\sum_{a\in\mathcal{A}\setminus\{\emptyset\}}\sum_{z\in\mathcal{Z}}$. Note that we exclude the empty set because when $a=\emptyset$, ${D}\left(P^{a,z}_{\theta}\,\middle\|P^{a,z}_{m}\right)$ is defined as zero by convention for any $z\in\mathcal{Z}$, $\theta,m\in\left[M\right]$. The notations used in the proposed test are summarized in Table I.
Notation Expression
$F_{t}$$\left\{\left\{\mathbf{X}_{s}\right\}_{s=1}^{t},\left\{A_{s}\right\}_{s=1}^{t},\left\{Z_{s}\right\}_{s=1}^{t}\right\}$
$S_{t,\theta,m}$ $\log\frac{\mathsf{P}_{\theta}^{(t)}(F_{t})}{\mathsf{P}_{m}^{(t)}(F_{t})}$
$l$ $\frac{1}{T^{1/6}}\left(\frac{\min_{\theta\in\left[M\right]}\min_{m\neq\theta}A_{\theta,m}}{\min_{\theta\in\left[M\right]}\min_{m\neq\theta}\tilde{A}_{\theta,m}}\right)^{2}$
$\epsilon(T)$ $\left(\log T\right)^{-\frac{1}{4}}$
$\tilde{A}_{\theta,m}$ $\frac{1}{h}\sum_{a,z}{D}\left(P^{a,z}_{\theta}\,\middle\|P^{a,z}_{m}\right)$
$A_{\theta,m}$ $\frac{\left(1-\epsilon(T)\right)e^{*}_{\theta|m}\left(\bm{\beta}^{\theta}\right)+\epsilon(T)\tilde{A}_{\theta,m}}{1+l}$
$I(T)$ $\epsilon(T)\min_{\theta\in\left[M\right]}\min_{m\neq\theta}\tilde{A}_{\theta,m}$
$b(l)$ $\frac{l^{3}}{4(1+l)^{3}}\left(\frac{I(T)}{4L}\right)^{2}$
$B(l)$ $2+\frac{2(1+l)^{2}}{l^{2}}\left(\frac{4L}{I(T)}\right)^{2}$
$q(l,n,I(T))$ $1+\frac{1}{b(l)}\Big{(}1+\log\left(MB(l)\left(1+b(l)\right)\right)\Big{)}$
$\Delta$ $-MB(l)\left(1+b(l)\right)e^{-b(l)(T-1)}$
$K(x)$ $\sqrt{xe+1}-1$ with domain:$[-e^{-1},0]$ and range:$[-1,0]$
$\epsilon_{\theta}$ $A^{\text{max}}_{\theta}\left(1-\frac{K\left(\Delta\right)}{b(l)}\right)$
$A^{\text{max}}_{\theta}$ $\max_{m\neq\theta}A_{\theta,m}$
TABLE I: Summary of notations in the proposed test.
Let $e$ be Euler’s number ($e\cong 2.71828...$), $q\left(l,n,I(T)\right)$ be defined in Table I, $n$ be the total number of data sources, $l$ and $I(T)$ be two specific sequences (both indexed by $T$) defined in Table I. Depending on the expected stopping time constraint $T$, our proposed test comprises two regimes:
  • Regime 1: $T<\max\{e,q\left(l,n,I(T)\right)\}$. The decision maker stops at time instance $t=1$, does not select any sources, and randomly guesses a hypothesis.
  • Regime 2: $T\geq\max\{e,q\left(l,n,I(T)\right)\}$. At each time instance $t\geq 1$, the decision maker carry out the test proposed below.
One may concern that the random guess in the first regime affects the error exponent. Nevertheless, since we only consider the asymptotic error performance (when $T\to\infty$) and by Lemma 2 in the Appendix, $q\left(l,n,I(T)\right)$ is sublinear in $T$, the random guess does not dominate the asymptotes of error probabilities. Note that the test in the first regime also naturally satisfies the constraints stated in (3).
Let us now provide the procedure of the test in the second regime as follows. Given any $\bm{\beta}^{0},\bm{\beta}^{1},\ldots,\bm{\beta}^{M-1}\in\mathcal{C}$, at each time instance $t\geq 1$:
  1. 1.
    Firstly, the decision maker conducts the randomized action taking following the pmf $P_{A_{t}|Z_{t},F_{t-1}}(\cdot|z,f_{t-1})$ with the MLE $\hat{\theta}_{\text{MLE}}$ at time slot $t-1$, under a revealed $Z_{t}=z\in\mathcal{Z}$. Specifically, if $\hat{\theta}_{\text{MLE}}=\theta$ at time slot $t-1$,
    $P_{A_{t}|Z_{t},F_{t-1}}(a|z,f_{t-1})=\begin{cases}(1-\epsilon(T))\frac{\beta^{\theta}_{a,z}}{\alpha_{z}}+\epsilon(T)\frac{1}{h\alpha_{z}},&a\in\mathcal{A}\setminus\{\emptyset\}\\[4.30554pt] 1-\sum_{a^{\prime}\in\mathcal{A}\setminus\{\emptyset\}}P_{A_{t}|Z_{t},F_{t-1}}(a^{\prime}|z,f_{t-1}),&a=\emptyset\end{cases}$
    where the exploration rate $\frac{1}{h}>0$ satisfies (9). Then, the decision maker collects a set of samples drawn from the selected data sources, namely,
    $\mathbf{X}_{t}\triangleq\{X_{t,j}\,|\,j\in A_{t}\cap Z_{t}\}.$
  2. 2.
    Secondly, after collecting $\mathbf{X}_{t}$, the decision maker decides to stop when there exists $\theta\in\left[M\right]$ such that $\forall\,m\in\left[M\right]\setminus\{\theta\}$, the cumulative log likelihood ratio
    $S_{t,\theta,m}\geq\left(TA_{\theta,m}-\epsilon_{\theta}\right),$ (10)
    and declares that the inferred hypothesis is the $\theta$ that satisfies (10). Note that $A_{\theta,m}$ and $\epsilon_{\theta}$, the parameters related to thresholding the log likelihood ratio, are defined in Table I. Otherwise, it will proceed to the next time slot. In words, the stopping time
    $\tau\triangleq\inf\left\{t\geq 1:\exists\,\theta\in\left[M\right],\min_{m\neq\theta}\left\{S_{t,\theta,m}-\left(TA_{\theta,m}-\epsilon_{\theta}\right)\right\}\geq 0\right\}.$
IV-C Analysis of the proposed test
The proof of achievability comprises three parts:
  1. 1.
    Demonstrating that our proposed test satisfies the expected stopping time constraints with the non-asymptotic analysis slightly modified from the achievability proofs of [2, Proposition 2] and [7, Theorem 4], which originated from [1, Theorem 1].
  2. 2.
    Showing that our proposed test satisfies the expected budget constraints, i.e., $\forall\,\theta\in\left[M\right]$, $\mathbf{B}^{\theta}\in\mathcal{C}_{B}$ ($\mathbf{B}^{\theta}$ and $\mathcal{C}_{B}$ are defined in Section II-C).
  3. 3.
    Verifying that, with a given $\{\bm{\beta}^{i}\,|\,\bm{\beta}^{i}\in\mathcal{C},\,i\in\left[M\right]\}$, our proposed test achieves the tuples of $M(M-1)$ individual error exponents $\big{\{}e_{m|\theta}\,\big{|}\,m\in\left[M\right],\theta\in\left[M\right]\setminus\{m\}\big{\}}$ satisfying
    $\big{\{}e_{m|\theta}\,\big{|}\,\theta\in\left[M\right]\setminus\{m\}\big{\}}\in\zeta_{m}(\bm{\beta}^{m})\quad\forall\,m\in\left[M\right].$
Note that Part 2) and 3) are similar to the achievability proof of [8, Theorem 1].
Throughout the proof of achievability, we are given a set of $\bm{\beta}^{0},\bm{\beta}^{1},\ldots,\bm{\beta}^{M-1}\in\mathcal{C}$.
IV-C1 Expected stopping time constraints
In this part, the goal is to show that for $T\geq 1$,
$\mathbb{E}_{\theta}[\tau]\leq T\quad\forall\,\theta\in\left[M\right].$ (11)
It is straightforward that the above is satisfied in Regime 1 of the proposed test. Hence in the following, we focus on Regime 2, that is, $T\geq\max\{e,q\left(l,n,I(T){}\right)\}$.
Let us first define a special stopping time $\tau_{\theta}$ for any $\theta\in\left[M\right]$ as follows:
$\tau_{\theta}\triangleq\inf\left\{t\,:\,\min_{m\neq\theta}\left\{S_{t,\theta,m}-TA_{\theta,m}+\epsilon_{\theta}\right\}\geq 0\right\}$
Therefore, by the definition of the stopping time $\tau$ in our proposed test, we have $\tau=\min_{\theta\in\left[M\right]}\tau_{\theta}$, leading to $\mathbb{E}_{\theta}\left[\tau\right]\leq\mathbb{E}_{\theta}\left[\tau_{\theta}\right]$, $\forall\,\theta\in\left[M\right]$. Hence, it suffices to show that in Regime 2 of our proposed test,
$\mathbb{E}_{\theta}\left[\tau_{\theta}\right]\leq T\quad\forall\,\theta\in\left[M\right].$ (12)
To proceed, for each $\theta\in\left[M\right]$, we decompose $\mathbb{E}_{\theta}\left[\tau_{\theta}\right]$ into two terms and upper bound them as
$\mathbb{E}_{\theta}\left[\tau_{\theta}\right]\leq\kappa_{\theta,T,l}+\sum_{k=\kappa_{\theta,T,l}}^{\infty}\mathbb{P}_{\theta}\left(\tau_{\theta}>k\right),$
where
$\kappa_{\theta,T,l}\triangleq\left\lceil T-\frac{\epsilon_{\theta}}{A^{\text{max}}_{\theta}}\right\rceil,\quad A^{\text{max}}_{\theta}\triangleq\max_{m\neq\theta}A_{\theta,m},$
and $\epsilon_{\theta}$ and $A_{\theta,m}$ are defined in Table I. In the following, we are going to first give a key lemma stating that the probability $\mathbb{P}_{\theta}\left(\tau_{\theta}>k\right)$ decays exponentially as $k\rightarrow\infty$. With the lemma, it is then simple to show that $\sum_{k=\kappa_{\theta,T,l}}^{\infty}\mathbb{P}_{\theta}\left(\tau_{\theta}>k\right)\leq\frac{\epsilon_{\theta}}{A^{\text{max}}_{\theta}}-1$ if $T\geq\max\{e,q\left(l,n,I(T)\right)\}$, leading to the conclusion that (12) is met, and consequently (11).
Lemma 1
For any $k\geq\kappa_{\theta,T,l}$ and $\theta\in\left[M\right]$,
$\mathbb{P}_{\theta}\left(\tau_{\theta}>k\right)\leq MB(l)e^{-kb(l)}$
where $l$, $b(l)$, and $B(l)$ are defined in Table I.
The proof of Lemma 1 is given in Section VI.
By Lemma 1 and the fact that $b(l)$ and $B(l)$ are positive, we upper bound $\sum_{k=\kappa_{\theta,T,l}}^{\infty}\mathbb{P}_{\theta}\left(\tau_{\theta}>k\right)$ as follows:
$\sum_{k=\kappa_{\theta,T,l}}^{\infty}\mathbb{P}_{\theta}\left(\tau_{\theta}>k\right)$ $\leq MB(l)\frac{e^{-\kappa_{\theta,T,l}b(l)}}{1-e^{-b(l)}}$
$\leq MB(l)\left(1+\frac{1}{b(l)}\right)e^{-\kappa_{\theta,T,l}b(l)}$
$=MB(l)\left(1+\frac{1}{b(l)}\right)e^{-\left\lceil T-\frac{\epsilon_{\theta}}{A^{\text{max}}_{\theta}}\right\rceil b(l)}$
$=MB(l)\left(1+\frac{1}{b(l)}\right)e^{-\left\lceil T-\left(1-\frac{K(\Delta)}{b(l)}\right)\right\rceil b(l)}$
$\leq MB(l)\left(1+\frac{1}{b(l)}\right)e^{-\left(T-\left(1-\frac{K(\Delta)}{b(l)}\right)\right)b(l)}$
$=\frac{1}{b(l)}MB(l)\left(b(l)+1\right)e^{-b(l)\left(T-1\right)}e^{-K(\Delta)}$
$=-\frac{1}{b(l)}\Delta e^{-K(\Delta)}$
$=-\frac{1}{b(l)}W_{0}(\Delta)e^{W_{0}(\Delta)}e^{-K(\Delta)}$ (13)
where $W_{0}(\cdot)$ is the principal branch of the Lambert function. Recall that $\Delta\triangleq-MB(l)\left(b(l)+1\right)e^{-b(l)\left(T-1\right)}$ as defined in Table I. Also note that for $T\geq\max\{e,q\left(l,n,I(T)\right)\}$, $\Delta$ satisfies $-e^{-1}\leq\Delta<0$ and hence falls in the domain of $W_{0}(\cdot)$.
Finally, we show that (13)$\leq\frac{\epsilon_{\theta}}{A^{\text{max}}_{\theta}}-1=-\frac{K(\Delta)}{b(l)}$ to complete the proof. By [11, Theorem 3.2],
$K(x)\leq W_{0}(x)\quad\text{for}\ -e^{-1}\leq x<0,$
and hence we have $W_{0}(\Delta)\geq K(\Delta)\geq-1$. Due to the fact $xe^{x}$ is increasing when $x\geq-1$, it is straightforward that $W_{0}(\Delta)e^{W_{0}(\Delta)}\geq K(\Delta)e^{K(\Delta)}$, completing the proof of the first part.
IV-C2 Linear expected budget constraints
To show that the proposed test satisfies the expected budget constraints, that is, $\mathbf{B}^{\theta}\in\mathcal{C}_{B}$ for every $\theta\in\left[M\right]$, or equivalently,
$b_{i}\left(\mathbf{B}^{\theta}\right)\leq U_{i}\equiv r_{i}T\quad\forall\,i\in\mathcal{I},\forall\,\theta\in\left[M\right],$
let us first introduce a stochastic process $\left\{M_{t,i}\,\middle|\,t\geq 0\right\}$ for every linear constraint function $b_{i}(\cdot)$ as follows:
$M_{t,i}\triangleq b_{i}\left(\mathbf{B}_{t}\right)-r_{i}t,$
where $\mathbf{B}_{t}=\left\{B_{t,1},B_{t,2},\ldots,B_{t,n}\right\}$ and $B_{t,j}\triangleq\sum_{s=1}^{t}\mathbbm{1}{\left\{j\in A_{t}\cap Z_{t}\right\}}$ denotes the number of times that source $j\in\left[n\right]$ is selected up to time instance $t\geq 1$. We then show that this stochastic process is a super martingale adapted to the natural filtration $\{\mathcal{F}_{t}\triangleq\sigma(F_{t})\}_{t\geq 0}$ under any given hypothesis $\theta\in\left[M\right]$. Specifically, we have
$\mathbb{E}_{\theta}\left[M_{t+1,i}\,\middle|\,\mathcal{F}_{t}\right]$ $=\mathbb{E}_{\theta}\left[b_{i}\left(\mathbf{B}_{t+1}\right)-\left(t+1\right)r_{i}\,\middle|\,\mathcal{F}_{t}\right]$
$=b_{i}(\mathbf{B}_{t})-r_{i}t+\mathbb{E}_{\theta}\left[b_{i}\left(\left\{\mathbbm{1}{\left\{j\in A_{t+1}\cap Z_{t+1}\right\}}\right\}_{j=1}^{n}\right)\,\middle|\,\mathcal{F}_{t}\right]-r_{i}$
$=M_{t,i}+\mathbb{E}_{\theta}\left[b_{i}\left(\left\{\mathbbm{1}{\left\{j\in A_{t+1}\cap Z_{t+1}\right\}}\right\}_{j=1}^{n}\right)\,\middle|\,\mathcal{F}_{t}\right]-r_{i}$
$\leq M_{t,i}$ (14)
where (14) holds because
$\mathbb{E}_{\theta}\left[b_{i}\left(\left\{\mathbbm{1}{\left\{j\in A_{t+1}\cap Z_{t+1}\right\}}\right\}_{j=1}^{n}\right)\,\middle|\,\mathcal{F}_{t}\right]$
$=b_{i}\left(\left\{\mathbb{E}_{\theta}\left[\mathbbm{1}{\left\{j\in A_{t+1}\cap Z_{t+1}\right\}}\,\middle|\,\mathcal{F}_{t}\right]\right\}_{j=1}^{n}\right)$
$=b_{i}\left(\left\{\sum_{a,z}P_{A_{t+1}|Z_{t+1},F_{t}}\left(a|z,f_{t}\right)\alpha_{z}\mathbbm{1}{\left\{j\in a\cap z\right\}}\right\}_{j=1}^{n}\right)$
$\leq\max_{\theta\in\left[M\right]}b_{i}\left(\left\{\sum_{a,z}\left(\left(1-\epsilon\left(T\right)\right)\beta^{\theta}_{a,z}+\epsilon(T)\frac{1}{h}\right)\mathbbm{1}{\left\{j\in a\cap z\right\}}\right\}_{j=1}^{n}\right)$ (15)
$=\max_{\theta\in\left[M\right]}\left(1-\epsilon(T)\right)b_{i}\left(\bm{\omega}(\bm{\beta}^{\theta})\right)+\epsilon(T)\frac{1}{h}b_{i}\left(\bm{\omega}(\mathbf{1})\right)$
$\leq(1-\epsilon(T))r_{i}+\epsilon(T)r_{i}$ (16)
$=r_{i}.$
Note that (15) is due to the action taking policy defined in our proposed test, and (16) is due to $\bm{\beta}^{\theta}\in\mathcal{C}$ for all $\theta\in\left[M\right]$ and the design of $\frac{1}{h}$ in our proposed test. Recall that $\bm{\omega}(\cdot)$ is defined in Theorem 1.
Consequently, by the fact that $\mathbb{E}_{\theta}[\tau]\leq T$ as shown Part 1) and the Optional Stopping Theorem, we have
$0=\mathbb{E}_{\theta}\left[M_{0,j}\right]\geq\mathbb{E}_{\theta}\left[M_{\tau,j}\right]\geq\mathbb{E}_{\theta}\left[b_{i}\left(\mathbf{B}_{\tau}\right)\right]-r_{i}T=b_{i}\left(\left\{\mathbb{E}_{\theta}\left[B_{\tau,j}\right]\right\}_{j=1}^{n}\right)-r_{i}T$
Plugging the definition of $\mathbf{B}^{\theta}$ (see Section II-C), we have
$b_{i}\left(\mathbf{B}^{\theta}\right)=b_{i}\left(\left\{\mathbb{E}_{\theta}\left[B_{\tau,j}\right]\right\}_{j=1}^{n}\right)\leq r_{i}T\quad\forall\,i\in\mathcal{I},\forall\,\theta\in\left[M\right],$
completing the proof of this part.
IV-C3 Achievable error exponents
Recall that in our proposed test, there are two regimes depending on the value of $T$: $T<\max\{e,q\left(l,n,I(T)\right)\}$ (regime 1) and $T\geq\max\{e,q\left(l,n,I(T)\right)\}$ (regime 2). As shown in the following lemma, $q\left(l,n,I(T)\right)\in o(T)$, and hence for $T$ sufficiently large, the test operates in the second regime. As a result, for deriving the achievable error exponent, we only need to consider the second regime.
Lemma 2
$q(l,n,I(T))$ is sublinear in $T$, that is, $q(l,n,I(T))\in o(T)$
The proof of Lemma 2 is given in Appendix -A.
Hence, given any hypotheses $\theta$, $\forall\,m\neq\theta$, when $T$ is sufficiently large, we have
$\pi_{m|\theta}$ $=\mathbb{P}_{\theta}\left(\min_{i\neq m}\left\{S_{\tau,m,i}-TA_{m,i}+\epsilon_{m}\right\}\geq 0\right)$
$\leq\mathbb{P}_{\theta}\left(S_{\tau,m,\theta}-TA_{m,\theta}+\epsilon_{m}\geq 0\right)$
$=\mathbb{P}_{\theta}\left(e^{S_{\tau,m,\theta}}\geq e^{TA_{m,\theta}-\epsilon_{m}}\right)$
$\leq\frac{\mathbb{E}_{\theta}\left[e^{S_{\tau,m,\theta}}\right]}{e^{TA_{m,\theta}-\epsilon_{m}}}.$
To proceed, it is shown in the next lemma that $\left\{e^{S_{t,m,\theta}}\,\middle|\,t\geq 0\right\}$ is a martingale
Lemma 3
The stochastic process $\left\{e^{S_{t,m,\theta}}\,\middle|\,t\geq 0\right\}$ is a martingale adapted to the natural filtration $\left\{\mathcal{F}_{t}\right\}_{t\geq 0}$ under hypothesis $\theta$, $\forall\,\theta,m\in\left[M\right]$, $\theta\neq m$.
The proof of Lemma 3 is given in Appendix -B.
Finally, by the Optional Stopping Theorem, $\mathbb{E}_{\theta}\left[e^{S_{\tau,m,\theta}}\right]=\mathbb{E}_{\theta}\left[e^{S_{0,m,\theta}}\right]=1$, and hence
$\pi_{m|\theta}\leq e^{-T\left(A_{m,\theta}-\frac{\epsilon_{m}}{T}\right)}$
for $T$ sufficiently large. Therefore, $\forall\,\theta,m\in\left[M\right]$, $\theta\neq m$, we can conclude that
$\liminf_{T\to\infty}\left\{\frac{1}{T}\log\frac{1}{\pi_{m|\theta}}\right\}\geq\liminf_{T\to\infty}\left\{A_{m,\theta}-\frac{\epsilon_{m}}{T}\right\}=e^{*}_{m|\theta}(\bm{\beta}^{m})$
where $\epsilon_{m}\in o(T)$ as defined in Table I for any $m\in\left[M\right]$. The proof of this part is complete.
V Proof of Converse
Let us first denote the binary KL divergence and the binary entropy function as follows: for $p,q\in[0,1]$,
${d}\left(p\,\middle\|q\right)$ $\textstyle\triangleq p\log\frac{p}{q}+(1-p)\log\frac{1-p}{1-q}={D}\left(\mathrm{Ber}(p)\,\middle\|\mathrm{Ber}(q)\right),$
$h(p)$ $\textstyle\triangleq p\log\frac{1}{p}+(1-p)\log\frac{1}{1-p}=H(\mathrm{Ber}(p)).$
To prove the converse part of Theorem 1, it suffices to show that, for any achievable $M(M-1)$ types of error exponents $\Big{\{}e_{m|\theta}\,\Big{|}\,m\in\left[M\right],\theta\in\left[M\right]\setminus\{m\}\Big{\}}$, there exist $\bm{\beta}^{0},\bm{\beta}^{1},...,\bm{\beta}^{M-1}\in\mathcal{C}$ such that $\forall\,m\in\left[M\right]$ $\forall\,\theta\in\left[M\right]\setminus\{m\}$,
$e_{m|\theta}\leq\sum_{a\in\mathcal{A}}\sum_{z\in\mathcal{Z}}\beta^{m}_{a,z}{D}\left(P^{a,z}_{m}\,\middle\|P^{a,z}_{\theta}\right),$
thereby the $\Big{\{}e_{m|\theta}\,\Big{|}\,m\in\left[M\right],\theta\in\left[M\right]\setminus\{m\}\Big{\}}\in\zeta$, the optimal error exponent region defined in Theorem 1.
The proof of converse is outlined as follows. First, we utilize the data processing inequality and Optional Stopping Theorem to find the upper bound of the $\frac{1}{T}\log\frac{1}{\pi_{m|\theta}}$ for any $m\in\left[M\right]$, $\theta\in\left[M\right]\setminus\{m\}$, $\forall\,T\geq 1$, where the action taking frequency that appears in the upper bound plays a key role. Then, we derive the necessary conditions for the action taking frequency to meet the linear expected budget constraints including expected stopping time constraints. We then observe that the collection of action taking frequencies satisfying the necessary conditions is the constraint set $\mathcal{C}$ defined in Theorem 1. Finally, we leverage the fact that $\mathcal{C}$ is a compact set to complete the proof.
Let us start the proof with an useful decomposition of $\log\frac{1}{\pi_{m|\theta}}$: $\forall\,m\in\left[M\right],\ \forall\,\theta\in\left[M\right]\setminus\{m\}$,
$\log\frac{1}{\pi_{m|\theta}}=\frac{1}{\pi_{m|m}}\left({d}\left(\pi_{m|m}\,\middle\|\pi_{m|\theta}\right)+h\left(\pi_{m|m}\right)+\mathbb{P}_{m}\left(\phi\neq m\right)\log\mathbb{P}_{\theta}\left(\phi\neq m\right)\right).$ (17)
The derivation of the above decomposition goes as follows:
${d}\left(\pi_{m|m}\,\middle\|\pi_{m|\theta}\right)$ $=\mathbb{P}_{m}\left(\phi=m\right)\log{\left(\frac{\mathbb{P}_{m}\left(\phi=m\right)}{\mathbb{P}_{\theta}\left(\phi=m\right)}\right)}+\mathbb{P}_{m}\left(\phi\neq m\right)\log{\left(\frac{\mathbb{P}_{m}\left(\phi\neq m\right)}{\mathbb{P}_{\theta}\left(\phi\neq m\right)}\right)}$
$=-\mathbb{P}_{m}\left(\phi=m\right)\log\frac{1}{\mathbb{P}_{m}\left(\phi=m\right)}-\mathbb{P}_{m}\left(\phi\neq m\right)\log\frac{1}{\mathbb{P}_{m}\left(\phi\neq m\right)}$
$\quad-\mathbb{P}_{m}\left(\phi=m\right)\log\mathbb{P}_{\theta}\left(\phi=m\right)-\mathbb{P}_{m}\left(\phi\neq m\right)\log\mathbb{P}_{\theta}\left(\phi\neq m\right)$
$=-h\left(\pi_{m|m}\right)-\pi_{m|m}\log\pi_{m|\theta}-\mathbb{P}_{m}\left(\phi\neq m\right)\log\mathbb{P}_{\theta}\left(\phi\neq m\right).$
Rearranging terms leads to (17).
Next, by the data processing inequality,
${d}\left(\pi_{m|m}\,\middle\|\pi_{m|\theta}\right)\leq{D}\left(\mathsf{P}_{m}^{(\tau)}\,\middle\|\mathsf{P}_{\theta}^{(\tau)}\right)=\mathbb{E}_{m}\left[S_{\tau,m,\theta}\right],$ (18)
where
$S_{\tau,m,\theta}$ $=\sum_{s=1}^{\tau}\log\frac{P^{A_{s},Z_{s}}_{m}(\mathbf{X}_{s})}{P^{A_{s},Z_{s}}_{\theta}(\mathbf{X}_{s})}$
$=\sum_{s=1}^{\tau}\sum_{a\in\mathcal{A}}\sum_{z\in\mathcal{Z}}\mathbbm{1}\{A_{s}=a,Z_{s}=z\}\log\frac{P^{a,z}_{m}(\mathbf{X}_{s})}{P^{a,z}_{\theta}(\mathbf{X}_{s})}.$
Note that this explicit form of the cumulative log-likelihood ratio was derived in Lemma 3. Then, by Optional Stopping Theorem, we have
$\mathbb{E}_{m}\left[S_{\tau,m,\theta}\right]=\sum_{a\in\mathcal{A}}\sum_{z\in\mathcal{Z}}\mathbb{E}_{m}\left[\sum_{s=1}^{\tau}\mathbbm{1}{\left\{A_{s}=a,Z_{s}=z\right\}}\right]{D}\left(P^{a,z}_{m}\,\middle\|P^{a,z}_{\theta}\right).$ (19)
To this end, let us define the frequency for any $a\in\mathcal{A}$ and $z\in\mathcal{Z}$ as follows:
$\beta^{m,(T)}_{a,z}\triangleq\frac{\mathbb{E}_{m}\left[\sum_{s=1}^{\tau}\mathbbm{1}{\left\{A_{s}=a,Z_{s}=z\right\}}\right]}{T}\quad\forall\,T\geq 1.$ (20)
Let $\bm{\beta}^{m,(T)}\triangleq\left\{\beta^{m,(T)}_{a,z}\,\middle|\,a\in\mathcal{A},z\in\mathcal{Z}\right\}$. In the following, we show that
$\bm{\beta}^{m,(T)}\in\mathcal{C}\quad\forall\,T\geq 1,\,\forall\,m\in\left[M\right]$ (21)
are a set of necessary conditions that $\bm{\beta}^{m,(T)},\ m\in\left[M\right]$ should satisfy for the constraints in (3) to be met. Recall that $\mathcal{C}$ is defined in Theorem 1.
  1. 1.
    As for the linear expected budget constraints, plugging the definition of $B^{m}_{j}$ (see Section II-C), we have $\forall\,j\in\left[n\right]$,
    $B^{m}_{j}$ $=\sum_{t=1}^{\infty}\mathbb{E}_{m}\left[\mathbbm{1}{\left\{j\in A_{t}\cap Z_{t}\right\}}\mathbbm{1}{\left\{t\leq\tau\right\}}\right]$
    $=\sum_{t=1}^{\infty}\sum_{a,z}\mathbb{P}_{m}\left(A_{t}=a,Z_{t}=z\right)\mathbbm{1}{\left\{j\in a\cap z\right\}}\mathbb{E}_{m}\left[\mathbbm{1}{\left\{t\leq\tau\right\}}\,\middle|\,A_{t}=a,Z_{t}=z\right]$
    $=\sum_{a,z}\mathbbm{1}{\left\{j\in a\cap z\right\}}\sum_{t=1}^{\infty}\mathbb{P}_{m}\left(A_{t}=a,Z_{t}=z,t\leq\tau\right)$
    $=\sum_{a,z}\mathbbm{1}{\left\{j\in a\cap z\right\}}\mathbb{E}_{m}\left[\sum_{t=1}^{\tau}\mathbbm{1}{\left\{A_{t}=a,Z_{t}=z\right\}}\right].$
    Since $\mathbf{B}^{m}\in\mathcal{C}_{B}\ \forall\,m\in\left[M\right]$, we have $b_{i}(\mathbf{B}^{m})\leq U_{i}\ \forall\,i\in\mathcal{I}$, and by the definition of $\bm{\beta}^{m,(T)}$ in (20), we have
    $b_{i}\left(\bm{\omega}\left(\bm{\beta}^{m,(T)}\right)\right)\leq\frac{U_{i}}{T}=r_{i}\quad\forall\,i\in\mathcal{I},$
    which is obtained by dividing both sides of the inequality $b_{i}(\mathbf{B}^{m})\leq U_{i}\ \forall\,i\in\mathcal{I}$ by $T$. Recall that $\bm{\omega}(\cdot)$ is defined in Theorem 1 and each $b_{i}(\cdot)$ is a linear function defined in Section II-C.
  2. 2.
    For the expected stopping time constraint, since $\mathbb{E}_{m}\left[\tau\right]\leq T$, we have: $\forall\,z\in\mathcal{Z}$,
    $\sum_{a\in\mathcal{A}}\beta^{m,(T)}_{a,z}=\frac{\mathbb{E}_{m}\left[\sum_{s=1}^{\tau}\mathbbm{1}{\left\{Z_{s}=z\right\}}\right]}{T}=\frac{\mathbb{E}_{m}\left[\tau\right]\alpha_{z}}{T}\leq\alpha_{z}.$
With the necessary conditions (21), we combine them along with (17), (18), and (19) to arrive at: $\forall T\geq 1$, $\forall\,m\in\left[M\right]$, $\forall\,\theta\in\left[M\right]\setminus\{m\}$,
$\frac{1}{T}\log\frac{1}{\pi_{m|\theta}}$ $\leq\frac{1}{\pi_{m|m}}\left(\sum_{a\in\mathcal{A}}\sum_{z\in\mathcal{Z}}\beta^{m,(T)}_{a,z}{D}\left(P^{a,z}_{m}\,\middle\|P^{a,z}_{\theta}\right)+\frac{h\left(\pi_{m|m}\right)}{T}+\frac{\mathbb{P}_{m}\left(\phi\neq m\right)\log\mathbb{P}_{\theta}\left(\phi\neq m\right)}{T}\right)$
$\leq\frac{1}{\pi_{m|m}}\left(\sum_{a\in\mathcal{A}}\sum_{z\in\mathcal{Z}}\beta^{m,(T)}_{a,z}{D}\left(P^{a,z}_{m}\,\middle\|P^{a,z}_{\theta}\right)+\frac{h\left(\pi_{m|m}\right)}{T}\right),$ (22)
where (22) holds since $\mathbb{P}_{\theta}\left(\phi\neq m\right)\leq 1$. We then take liminf on both sides of the inequality (22) and get
$\liminf_{T\to\infty}\left\{\frac{1}{T}\log\frac{1}{\pi_{m|\theta}}\right\}\leq\liminf_{T\to\infty}\left\{\sum_{a\in\mathcal{A}}\sum_{z\in\mathcal{Z}}\beta^{m,(T)}_{a,z}{D}\left(P^{a,z}_{m}\,\middle\|P^{a,z}_{\theta}\right)\right\}$
since $\pi_{m|m}\rightarrow 1$ as $T\rightarrow\infty$ and $h(\pi_{m|m})\leq\log(2)$.
Now, since $\mathcal{C}$ is a compact set for any index set $\mathcal{I}$, by the Bolzano-Weierstrass Theorem, there exists a convergent subsequence of $\left\{\bm{\beta}^{m,(T)}\,\middle|\,T\geq 1\right\}$ indexed by some $\left\{T_{k}\right\}_{k\in{\mathbb{N}_{+}}}$, i.e., $\left\{\bm{\beta}^{m,(T_{k})}\,\middle|\,k\in{\mathbb{N}_{+}}\right\}$, converging to a $\bm{\beta}^{m,*}\in\mathcal{C}$. Furthermore, because the liminf over a sequence is upper bounded by the liminf over any of its subsequences, we have $\forall\,m\in\left[M\right]$, $\forall\,\theta\in\left[M\right]\setminus\{m\}$,
$\liminf_{T\to\infty}\left\{\frac{1}{T}\log\frac{1}{\pi_{m|\theta}}\right\}$ $\leq\liminf_{k\to\infty}\left\{\sum_{a\in\mathcal{A}}\sum_{z\in\mathcal{Z}}\beta^{m,(T_{k})}_{a,z}{D}\left(P^{a,z}_{m}\,\middle\|P^{a,z}_{\theta}\right)\right\}$
$=\sum_{a\in\mathcal{A}}\sum_{z\in\mathcal{Z}}\beta^{m,*}_{a,z}{D}\left(P^{a,z}_{m}\,\middle\|P^{a,z}_{\theta}\right),$
where the equality holds true since $\bm{\beta}^{m,(T_{k})}$ converges to $\bm{\beta}^{m,*}$, and those KL-divergences are constant.
Consequently, for any achievable $M(M-1)$ types of error exponents $\Big{\{}e_{m|\theta}\,\Big{|}\,m\in\left[M\right],\theta\in\left[M\right]\setminus\{m\}\Big{\}}$, we have shown that $\forall\,m\in\left[M\right]$, there exists a $\bm{\beta}^{m}\in\mathcal{C}$ ($\bm{\beta}^{m,*}$ here) such that
$e_{m|\theta}\leq\sum_{a\in\mathcal{A}}\sum_{z\in\mathcal{Z}}\beta^{m}_{a,z}{D}\left(P^{a,z}_{m}\,\middle\|P^{a,z}_{\theta}\right)\quad\forall\,\theta\in\left[M\right]\setminus\{m\},$
thereby completing the proof. ∎
VI Proof of Lemma 1
Without loss of generality, let us focus on the case where hypothesis $\theta$ is the ground truth. Define the following two sets for any time slot $t\geq 1$ as in the proof of [2, Proposition 2]:
$\mathbf{Q}_{\theta}(t)\triangleq\left\{F_{t}\,:\,\min_{m\neq\theta}\left\{S_{t,\theta,m}-\left(TA_{\theta,m}-\epsilon_{\theta}\right)\right\}<0\right\}$
$\tilde{\mathbf{Q}}_{\theta}(t)\triangleq\left\{F_{t}\,:\,\min_{m\neq\theta}S_{t,\theta,m}<0\right\}$
where $F_{t}$ given in Definition 1 denotes the information accumulated up to the end to time $t$. Furthermore, define $\tilde{\tau}_{\theta}$ as follows:
$\tilde{\tau}_{\theta}\triangleq\inf\left\{t\,:\,\min_{m\neq\theta}S_{t^{\prime},\theta,m}\geq 0\,,\forall\,t^{\prime}\geq t\right\}.$
Note that for any time slot $t\geq\tilde{\tau}_{\theta}$, the maximum likelihood estimate $\hat{\theta}_{\text{MLE}}=\theta$.
We are ready to upper bound $\mathbb{P}_{\theta}(\tau_{\theta}>k)$. By the definitions of the stopping time $\tau_{\theta}$ and the region $\mathbf{Q}_{\theta}(k)$ we have: $\forall\,k\geq\kappa_{\theta,T,l}$,
$\mathbb{P}_{\theta}\left(\tau_{\theta}>k\right)$ $\leq\mathbb{P}_{\theta}\left(\mathbf{Q}_{\theta}(k)\right)$
$=\mathbb{P}_{\theta}\left(\mathbf{Q}_{\theta}(k)\cap\left\{\tilde{\tau}_{\theta}\leq g_{k,l}\right\}\right)+\mathbb{P}_{\theta}\left(\mathbf{Q}_{\theta}(k)\cap\left\{\tilde{\tau}_{\theta}>g_{k,l}\right\}\right)$
$\leq\mathbb{P}_{\theta}\left(\mathbf{Q}_{\theta}(k)\cap\left\{\tilde{\tau}_{\theta}\leq g_{k,l}\right\}\right)+\mathbb{P}_{\theta}\left(\tilde{\tau}_{\theta}>g_{k,l}\right)$
$\leq\underbrace{\mathbb{P}_{\theta}\left(\mathbf{Q}_{\theta}(k)\cap\left\{\tilde{\tau}_{\theta}\leq g_{k,l}\right\}\right)}_{(i)}+\sum_{t^{\prime}=g_{k,l}}^{\infty}\underbrace{\mathbb{P}_{\theta}\left(\tilde{\mathbf{Q}}_{\theta}(t^{\prime})\right)}_{(ii)}$
where $g_{k,l}\triangleq k\left(\frac{l/2}{1+l}\right),\forall\,k\geq 1$. Recall that $l$ is a specific sequence defined in Table I.
In the following, we further derive upper bounds for the term $(i)$ and the term $(ii)$ respectively.
VI-A Upper bounding the term (i)
For the upper bound of the term $(i)$, we slightly modify the achievability proof in [7, Theorem 4]. Specifically, by union bound and the definition of $\mathbf{Q}_{\theta}(k)$, we have
$(i)$ $=\mathbb{P}_{\theta}\left(\left\{\min_{m\neq\theta}\left\{S_{k,\theta,m}-\left(TA_{\theta,m}-\epsilon_{\theta}\right)\right\}<0\right\}\cap\left\{\tilde{\tau}_{\theta}\leq g_{k,l}\right\}\right)$
$\leq\sum_{m\neq\theta}\mathbb{P}_{\theta}\left(\left\{S_{k,\theta,m}<TA_{\theta,m}-\epsilon_{\theta}\right\}\cap\left\{\tilde{\tau}_{\theta}\leq g_{k,l}\right\}\right)$
$=\sum_{m\neq\theta}\mathbb{P}_{\theta}\left(\left\{U_{k}<TA_{\theta,m}-\epsilon_{\theta}-\sum_{s=1}^{k}\mathbb{E}_{\theta}\left[\text{LLR}^{A_{s},Z_{s}}_{\theta,m}(\mathbf{X}_{s})\,\middle|\,\mathcal{F}_{s-1}\right]\right\}\cap\left\{\tilde{\tau}_{\theta}\leq g_{k,l}\right\}\right)$
where
$\text{LLR}^{A_{s},Z_{s}}_{\theta,m}(\mathbf{X}_{s})\triangleq\log\frac{P^{A_{s},Z_{s}}_{\theta}(\mathbf{X}_{s})}{P^{A_{s},Z_{s}}_{m}(\mathbf{X}_{s})}\quad\forall\,s\geq 1,$
$U_{k}\triangleq S_{k,\theta,m}-\sum_{s=1}^{k}\mathbb{E}_{\theta}\left[\text{LLR}^{A_{s},Z_{s}}_{\theta,m}(\mathbf{X}_{s})\,\middle|\,\mathcal{F}_{s-1}\right].$ (23)
To further bound the term $(i)$, we proceed by deriving a lower bound for $\sum_{s=1}^{k}\mathbb{E}_{\theta}\left[\text{LLR}_{\theta,m}(s)\,\middle|\,\mathcal{F}_{s-1}\right]$,
$\sum_{s=1}^{k}\mathbb{E}_{\theta}\left[\text{LLR}^{A_{s},Z_{s}}_{\theta,m}(\mathbf{X}_{s})\,\middle|\,\mathcal{F}_{s-1}\right]$ $=\sum_{s=1}^{\tilde{\tau}_{\theta}}\mathbb{E}_{\theta}\left[\text{LLR}^{A_{s},Z_{s}}_{\theta,m}(\mathbf{X}_{s})\,\middle|\,\mathcal{F}_{s-1}\right]+\sum_{s=\tilde{\tau}_{\theta}+1}^{k}\mathbb{E}_{\theta}\left[\text{LLR}^{A_{s},Z_{s}}_{\theta,m}(\mathbf{X}_{s})\,\middle|\,\mathcal{F}_{s-1}\right]$
$\geq\sum_{s=\tilde{\tau}_{\theta}+1}^{k}\mathbb{E}_{\theta}\left[\text{LLR}^{A_{s},Z_{s}}_{\theta,m}(\mathbf{X}_{s})\,\middle|\,\mathcal{F}_{s-1}\right]$ (24)
$=\sum_{s=\tilde{\tau}_{\theta}+1}^{k}\sum_{a,z}P_{A_{s},Z_{s}|\mathcal{F}_{s-1}}\left(a,z\,\middle|\,\mathcal{F}_{s-1}\right){D}\left(P^{a,z}_{\theta}\,\middle\|P^{a,z}_{m}\right)$ (25)
$=\sum_{s=\tilde{\tau}_{\theta}+1}^{k}\sum_{a,z}P_{A_{s}|Z_{s},F_{s-1}}\left(a\,\middle|\,z,f_{s-1}\right)\alpha_{z}{D}\left(P^{a,z}_{\theta}\,\middle\|P^{a,z}_{m}\right)$ (26)
$=\sum_{s=\tilde{\tau}_{\theta}+1}^{k}\sum_{a,z}\left(\left(1-\epsilon(T)\right)\beta^{\theta}_{a,z}+\epsilon(T)\frac{1}{h}\right){D}\left(P^{a,z}_{\theta}\,\middle\|P^{a,z}_{m}\right).$ (27)
Here (24) and (25) hold because $\forall\,s\geq 1$, we have
$\mathbb{E}_{\theta}\left[\text{LLR}^{A_{s},Z_{s}}_{\theta,m}(\mathbf{X}_{s})\,\middle|\,\mathcal{F}_{s-1}\right]$ $=\sum_{a,z}P_{A_{s},Z_{s}|\mathcal{F}_{s-1}}\left(a,z\,\middle|\,\mathcal{F}_{s-1}\right)\mathbb{E}_{\theta}\left[\log\frac{P^{a,z}_{\theta}(\mathbf{X}_{s})}{P^{a,z}_{m}(\mathbf{X}_{s})}\,\middle|\,A_{s}=a,Z_{s}=z,\mathcal{F}_{s-1}\right]$
$=\sum_{a,z}P_{A_{s},Z_{s}|\mathcal{F}_{s-1}}\left(a,z\,\middle|\,\mathcal{F}_{s-1}\right){D}\left(P^{a,z}_{\theta}\,\middle\|P^{a,z}_{m}\right)$
$\geq 0.$
(26) holds since the current available set $Z_{s}$ and the past information $F_{s-1}$ are independent and we switch from conditioning on $\mathcal{F}_{s-1}$ to merely conditioning on $F_{s-1}$ by Definition 1. (27) is due to our proposed test in the case of $\hat{\theta}_{\text{MLE}}=\theta$ and by the definition of $\tilde{\tau}_{\theta}$.
By the definitions of $A_{\theta,m}$ and $\tilde{A}_{\theta,m}$ in Table I, (27) can be expressed in an alternative form as follows:
(27)$=A_{\theta,m}\left(1+l\right)\left(k-\tilde{\tau}_{\theta}\right),$
where $l$ is a specific sequence also defined in Table I.
Consequently, the upper bound of the term $(i)$ becomes
$(i)$ $=\mathbb{P}_{\theta}\left(\mathbf{Q}_{\theta}(k)\cap\left\{\tilde{\tau}_{\theta}\leq g_{k,l}\right\}\right)$
$\leq\sum_{m\neq\theta}\mathbb{P}_{\theta}\left(\left\{U_{k}<TA_{\theta,m}-\epsilon_{\theta}-A_{\theta,m}\left(1+l\right)\left(k-\tilde{\tau}_{\theta}\right)\right\}\cap\left\{\tilde{\tau}_{\theta}\leq g_{k,l}\right\}\right)$
$\leq\sum_{m\neq\theta}\mathbb{P}_{\theta}\left(\left\{U_{k}<TA_{\theta,m}-\epsilon_{\theta}-k\left(1+\frac{l}{2}\right)A_{\theta,m}\right\}\cap\left\{\tilde{\tau}_{\theta}\leq g_{k,l}\right\}\right)$ (28)
$\leq\sum_{m\neq\theta}\mathbb{P}_{\theta}\Bigg{(}U_{k}<A_{\theta,m}\underbrace{\left(T-\frac{\epsilon_{\theta}}{A^{\text{max}}_{\theta}}-k\left(1+\frac{l}{2}\right)\right)}_{(*)}\Bigg{)},$ (29)
where (28) is because $\tilde{\tau}_{\theta}\leq g_{k,l}=k\left(\frac{l/2}{1+l}\right)$ and $A^{\text{max}}_{\theta}\triangleq\max_{m\neq\theta}A_{\theta,m}$ (see Table I).
To further bound (29), we leverage Azuma’s inequality (see the fact below) to bound each term in the summation.
Fact 1 (Azuma’s inequality)
Let $\{V_{i}\}_{i\geq 0}\triangleq\left\{V_{0},V_{1},...\right\}$ be a martingale. Furthermore, for any $i\geq 1$, if $\underline{V}_{i}\leq V_{i}-V_{i-1}\leq\overline{V}_{i}$ and there exists a finite sequence $\{C_{i}\}_{i\geq 1}\triangleq\left\{C_{1},C_{2},...\right\}$ such that $\overline{V}_{i}-\underline{V}_{i}\leq C_{i}$ for any $i\geq 1$, then we have for any $v>0$,
$P\left(V_{i}-V_{0}\leq-v\right)\leq\exp{\left\{\frac{-2v^{2}}{\sum_{s=1}^{i}C^{2}_{s}}\right\}}.$
In the following, we shall first show that $\{U_{i}\}_{i\geq 0}$ as defined in (23) is a martingale. Then we identify the parameters $\{C_{i}\}_{i\geq 1}$ and $v$ in Azuma’s inequality (Fact 1) to establish an upper bound of
$\mathbb{P}_{\theta}\left(U_{k}<A_{\theta,m}\left(T-\frac{\epsilon_{\theta}}{A^{\text{max}}_{\theta}}-k\left(1+\frac{l}{2}\right)\right)\right).$
  1. 1.
    To show that $\{U_{i}\}_{i\geq 0}$ is a martingale, note that by plugging the expression of $S_{k,\theta,m}$ shown in Lemma 3 into the definition of $U_{i}$ in (23), we have
    $U_{i}=\begin{cases}\sum_{s=1}^{i}\left(\text{LLR}^{A_{s},Z_{s}}_{\theta,m}(\mathbf{X}_{s})-\mathbb{E}_{\theta}\left[\text{LLR}^{A_{s},Z_{s}}_{\theta,m}(\mathbf{X}_{s})\,\middle|\,\mathcal{F}_{s-1}\right]\right),&i\geq 1\\[4.30554pt] 0,&i=0\end{cases}.$ (30)
    The martingale property of the stochastic process $\{U_{i}\}_{i\geq 0}$ can then be shown as follows:
    $\mathbb{E}_{\theta}\left[U_{i}-U_{i-1}\,\middle|\,\mathcal{F}_{i-1}\right]&=\mathbb{E}_{\theta}\left[\text{LLR}^{A_{i},Z_{i}}_{\theta,m}(\mathbf{X}_{i})-\mathbb{E}_{\theta}\left[\text{LLR}^{A_{i},Z_{i}}_{\theta,m}(\mathbf{X}_{i})\,\middle|\,\mathcal{F}_{i-1}\right]\,\middle|\,\mathcal{F}_{i-1}\right]\\ &=\mathbb{E}_{\theta}\left[\text{LLR}^{A_{i},Z_{i}}_{\theta,m}(\mathbf{X}_{i})\,\middle|\,\mathcal{F}_{i-1}\right]-\mathbb{E}_{\theta}\left[\mathbb{E}_{\theta}\left[\text{LLR}^{A_{i},Z_{i}}_{\theta,m}(\mathbf{X}_{i})\,\middle|\,\mathcal{F}_{i-1}\right]\,\middle|\,\mathcal{F}_{i-1}\right]\\ &=\mathbb{E}_{\theta}\left[\text{LLR}^{A_{i},Z_{i}}_{\theta,m}(\mathbf{X}_{i})\,\middle|\,\mathcal{F}_{i-1}\right]-\mathbb{E}_{\theta}\left[\text{LLR}^{A_{i},Z_{i}}_{\theta,m}(\mathbf{X}_{i})\,\middle|\,\mathcal{F}_{i-1}\right]\\ &=0\end{aligned}.$
  2. 2.
    To identify the parameters $\{C_{i}\}_{i\geq 1}$, first note that by Assumption 1, it is straightforward that
    $|\text{LLR}^{a,z}_{\theta,m}(\mathbf{x})|=\left\lvert\log\frac{P^{a,z}_{\theta}(\mathbf{x})}{P^{a,z}_{m}(\mathbf{x})}\right\rvert\leq L$
    for all $a\in\mathcal{A}$, $z\in\mathcal{Z}$, $\theta,m\in\left[M\right]$, $\theta\neq m$, $\mathbf{x}\in\mathcal{X}^{a,z}_{\theta}$, where $\mathcal{X}^{a,z}_{\theta}$ stands for the support of $P^{a,z}_{\theta}$. Hence, we have $\forall\,i\geq 1$,
    $-2L\leq U_{i}-U_{i-1}\leq 2L$
    where $U_{i}-U_{i-1}=\text{LLR}^{A_{i},Z_{i}}_{\theta,m}(\mathbf{X}_{i})-\mathbb{E}_{\theta}\left[\text{LLR}^{A_{i},Z_{i}}_{\theta,m}(\mathbf{X}_{i})\,\middle|\,\mathcal{F}_{i-1}\right]$. As a result, by assigning $\overline{V}_{i}=2L$ and $\underline{V}_{i}=-2L$, we have $\overline{V}_{i}-\underline{V}_{i}\leq 4L$ and hence choose $C_{i}=4L$, $\forall\,i\geq 1$.
  3. 3.
    We aim to identify the parameter $v$ as
    $v=A_{\theta,m}\left(k\left(1+\frac{l}{2}\right)+\frac{\epsilon_{\theta}}{A^{\text{max}}_{\theta}}-T\right).$
    To do so, we just need to show that the term ()(*)( ∗ ) defined in (29) times $A_{\theta,m}$ is smaller than zero.
    First note that by Assumption 2, we have $\tilde{A}_{\theta,m}>0$, namely,
    $A_{\theta,m}>0\quad\forall\,\theta,m\in\left[M\right],\theta\neq m.$
    It remains to show that the term $(*)<0$. When $k\geq\kappa_{\theta,T,l}=\lceil T-\frac{\epsilon_{\theta}}{A^{\text{max}}_{\theta}}\rceil$, we have $\forall\,T\geq q\left(l,n,I(T)\right)$,
    ((*)( ∗ ) $\leq T-\frac{\epsilon_{\theta}}{A^{\text{max}}_{\theta}}-\left(1+\frac{l}{2}\right)\kappa_{\theta,T,l}$
    $\leq T-\frac{\epsilon_{\theta}}{A^{\text{max}}_{\theta}}-\left(1+\frac{l}{2}\right)\left(T-\frac{\epsilon_{\theta}}{A^{\text{max}}_{\theta}}\right)$
    $=-\frac{l}{2}\left(T-\frac{\epsilon_{\theta}}{A^{\text{max}}_{\theta}}\right)<0$ (31)
    where (31) holds because $l\geq 0$ and by the following derivation:
    $T-\frac{\epsilon_{\theta}}{A^{\text{max}}_{\theta}}$ $=T-(1-\frac{K(\Delta)}{b(l)})$
    $\geq T-1-\frac{1}{b(l)}$ (32)
    $\geq q\left(l,n,I(T)\right)-1-\frac{1}{b(l)}$
    $=\frac{\log\left(MB(l)(1+b(l))\right)}{b(l)}>0\quad\forall\,T\geq q\left(l,n,I(T)\right).$
    Note that in the derivation above, we plug in the definitions of $\epsilon_{\theta}$, $q\left(l,n,I(T)\right)$, $\Delta$, $b(l)$, $B(l)$, and $K(\cdot)$ as shown in Table I. More specifically, (32) holds because as $T\geq q(l,n,I(T))$, we have $\Delta\geq-e^{-1}$ and hence $-1\leq K(\Delta)\leq 0$ by the range of the function $K(\cdot)$.
In summary, we apply Fact 1 with $C_{i}=4L\ \forall\,i\geq 1$ and $v=A_{\theta,m}\left(k\left(1+\frac{l}{2}\right)+\frac{\epsilon_{\theta}}{A^{\text{max}}_{\theta}}-T\right)$ and get
$\mathbb{P}_{\theta}\left(U_{k}<A_{\theta,m}\left(T-\frac{\epsilon_{\theta}}{A^{\text{max}}_{\theta}}-k\left(1+\frac{l}{2}\right)\right)\right)\leq\exp{\left\{\frac{-2A^{2}_{\theta,m}\left(k\left(1+\frac{l}{2}\right)+\frac{\epsilon_{\theta}}{A^{\text{max}}_{\theta}}-T\right)^{2}}{k\left(4L\right)^{2}}\right\}}.$
Consequently,
$(i)$ $\leq\sum_{m\neq\theta}\exp{\left\{\frac{-2A^{2}_{\theta,m}\left(k\left(1+\frac{l}{2}\right)+\frac{\epsilon_{\theta}}{A^{\text{max}}_{\theta}}-T\right)^{2}}{k\left(4L\right)^{2}}\right\}}$
$\leq\sum_{m\neq\theta}\exp{\left\{\frac{-A^{2}_{\theta,m}k\left(\left(1+\frac{l}{2}\right)+\frac{1}{k}\left(\frac{\epsilon_{\theta}}{A^{\text{max}}_{\theta}}-T\right)\right)^{2}}{8L^{2}}\right\}}$
$\leq\sum_{m\neq\theta}e^{-\frac{A^{2}_{\theta,m}k(\frac{l}{2})^{2}}{8L^{2}}}$ (33)
$\leq Me^{-k\frac{l^{2}}{2}(\frac{\min_{m\neq\theta}A_{\theta,m}}{4L})^{2}}.$ (34)
where (33) holds because $\forall\,k\geq\kappa_{\theta,T,l}\triangleq\left\lceil T-\frac{\epsilon_{\theta}}{A^{\text{max}}_{\theta}}\right\rceil$, $1+\frac{1}{k}\left(\frac{\epsilon_{\theta}}{A^{\text{max}}_{\theta}}-T\right)\geq 0$.
VI-B Upper bounding the term $(ii)$
To derive the upper bound for the term $(ii)$, let us begin with
$(ii)$ $=\mathbb{P}_{\theta}\left(\tilde{\mathbf{Q}}_{\theta}(t^{\prime})\right)\leq\sum_{m\neq\theta}\mathbb{P}_{\theta}\left(S_{t^{\prime},\theta,m}<0\right)$ (35)
$=\sum_{m\neq\theta}\mathbb{P}_{\theta}\left(U_{t^{\prime}}<-\sum_{s=1}^{t^{\prime}}\mathbb{E}_{\theta}\left[\text{LLR}^{A_{s},Z_{s}}_{\theta,m}(\mathbf{X}_{s})\,\middle|\,\mathcal{F}_{s-1}\right]\right).$ (36)
Here (35) holds true by union bound and the definition of $\tilde{\mathbf{Q}}_{\theta}(t^{\prime})$. In (36), recall the form of $U_{t^{\prime}}$ in (30).
To proceed, we have $\forall\,s\geq 1$,
$\mathbb{E}_{\theta}\left[\text{LLR}^{A_{s},Z_{s}}_{\theta,m}(\mathbf{X}_{s})\,\middle|\,\mathcal{F}_{s-1}\right]$ $=\sum_{a,z}P_{A_{s}|Z_{s},F_{s-1}}\left(a\,\middle|\,z,f_{s-1}\right)\alpha_{z}{D}\left(P^{a,z}_{\theta}\,\middle\|P^{a,z}_{m}\right)$
$\geq\sum_{a,z}\epsilon(T)\frac{1}{h}{D}\left(P^{a,z}_{\theta}\,\middle\|P^{a,z}_{m}\right)$ (37)
$\geq\epsilon(T)\min_{\theta\in\left[M\right]}\min_{m\neq\theta}\frac{1}{h}\sum_{a,z}{D}\left(P^{a,z}_{\theta}\,\middle\|P^{a,z}_{m}\right)$
$=\epsilon(T)\min_{\theta\in\left[M\right]}\min_{m\neq\theta}\tilde{A}_{\theta,m}=I(T),$
where the last line follows from the definition of $\tilde{A}_{\theta,m}$ and $I(T)$ in Table I, and (37) follows directly from our proposed test: $\forall\,a\in\mathcal{A}\setminus\{\emptyset\}$, $\forall\,z\in\mathcal{Z}$, $\forall\,f_{s-1}$, $P_{A_{s}|Z_{s},F_{s-1}}\left(a\,\middle|\,z,f_{s-1}\right)\geq\epsilon(T)\frac{1}{h\alpha_{z}}$.
Hence, (36) is further upper bounded as
(36)$\leq\sum_{m\neq\theta}\mathbb{P}_{\theta}\left(U_{t^{\prime}}<-t^{\prime}I(T)\right).$ ) .
Again, applying Azuma’s inequality (Fact 1) with $v=t^{\prime}I(T)$ and $C_{i}=4L\ \forall\,i\geq 1$, we have
$(ii)$ $\leq Me^{-t^{\prime}\frac{I^{2}(T)}{8L^{2}}}$
$\leq Me^{-t^{\prime}\frac{l^{2}}{2}(\frac{I(T)/(1+l/2)}{4L})^{2}},$ (38)
where (38) holds because $\frac{I^{2}\left(T\right)}{8L^{2}}\geq\frac{I^{2}\left(T\right)}{8L^{2}}\left(\frac{l}{2\left(1+\frac{l}{2}\right)}\right)^{2}=\frac{l^{2}}{2}\left(\frac{I\left(T\right)/\left(1+l/2\right)}{4L}\right)^{2}$.
VI-C Combining the upper bounds of the term $(i)$ and the term $(ii)$
Finally, the upper bound of $\mathbb{P}_{\theta}(\tau_{\theta}>k)$ can be obtained by combining the upper bound of the term $(i)$ in (34) and that of the term $(ii)$ in (38) as follows,
$\mathbb{P}_{\theta}\left(\tau_{\theta}>k\right)$ $\leq Me^{-k\frac{l^{2}}{2}(\frac{\min_{m\neq\theta}A_{\theta,m}}{4L})^{2}}+\sum_{t^{\prime}=g_{k,l}}^{\infty}Me^{-t^{\prime}\frac{l^{2}}{2}(\frac{I(T)/(1+l/2)}{4L})^{2}}$
$=Me^{-k\frac{l^{2}}{2}(\frac{\min_{m\neq\theta}A_{\theta,m}}{4L})^{2}}+M\frac{e^{-k\frac{l^{3}}{4(1+l)}(\frac{I(T)/(1+l/2)}{4L})^{2}}}{1-e^{-\frac{l^{2}}{2}(\frac{I(T)/(1+l/2)}{4L})^{2}}}$
$\leq Me^{-k\frac{l^{2}}{2(1+l)^{2}}(\frac{D}{4L})^{2}}+M\frac{e^{-k\frac{l^{3}}{4(1+l)^{3}}(\frac{I(T)}{4L})^{2}}}{1-e^{-\frac{l^{2}}{2(1+l)^{2}}(\frac{I(T)}{4L})^{2}}},$ (39)
where $D\triangleq\min_{\theta\in\left[M\right]}\min_{m\neq\theta}A_{\theta,m}$.
Subsequently, by the definitions of $l$, $I(T)$, and $\epsilon(T)$ (see Table I), we have $\forall\,T\geq e$,
$D^{2}$ $=lT^{\frac{1}{6}}\left(\frac{I(T)}{\epsilon(T)}\right)^{2}=T^{\frac{1}{6}}l\left(\log(T)\right)^{\frac{1}{2}}I^{2}(T)>I^{2}(T)l>I^{2}(T)\frac{l}{2(1+l)},$
where $T^{1/6}\left(\log\left(T\right)\right)^{1/2}>1$, $\forall\,T\geq e$. Hence, it is straightforward that
$k\frac{l^{2}}{2\left(1+l\right)^{2}}\left(\frac{D}{4L}\right)^{2}>k\frac{l^{3}}{4(1+l)^{3}}\left(\frac{I(T)}{4L}\right)^{2}.$ (40)
Then, plugging (40) back into (39) and by the definitions of $b(l)$ and $B(l)$ (see Table I), we have
$\mathbb{P}_{\theta}\left(\tau_{\theta}>k\right)$ $\leq Me^{-k\frac{l^{3}}{4(1+l)^{3}}\left(\frac{I(T)}{4L}\right)^{2}}\left(1+\frac{1}{1-e^{-\frac{l^{2}}{2(1+l)^{2}}\left(\frac{I(T)}{4L}\right)^{2}}}\right)$
$\leq Me^{-kb(l)}\left(2+\frac{1}{\frac{l^{2}}{2(1+l)^{2}}\left(\frac{I(T)}{4L}\right)^{2}}\right)$ (41)
$=Me^{-kb(l)}B(l),$
where (41) holds since $\frac{1}{1-e^{-x}}\leq 1+\frac{1}{x},\forall\,x>0$. This completes the proof of Lemma 1. ∎
VII Conclusion
In this work, we characterize the optimal individual error exponent region in active sequential multiple-hypothesis testing problems with temporally available data sources under linear expected budget constraints (including the expected stopping time constraint). From the characterization of the region, it is found that the tradeoffs among the individual error exponents only appear when there are more than two hypotheses to be distinguished. Such tradeoffs are due to the heterogeneous level of informativeness across data sources in distinguishing a declared hypothesis against other ground truth hypotheses, and they appear even in the most basic setting in [1]. Our asymptotically optimal test strikes a balance between exploitation and exploration and enjoys a weaker assumption while achieving the optimal error exponents.
-A Proof of Lemma 2
Proof:
Recall the definition of $q(l,n,I(T))$ as well as the relevant terms in Table I:
$q(l,n,I(T))$ $\triangleq 1+\frac{1}{b(l)}\Big{(}1+\log\left(MB(l)\left(1+b(l)\right)\right)\Big{)},$
$l$ $\triangleq\frac{1}{T^{1/6}}\left(\frac{\min_{\theta\in\left[M\right]}\min_{m\neq\theta}A_{\theta,m}}{,}{\min_{\theta\in\left[M\right]}\min_{m\neq\theta}\tilde{A}_{\theta,m}}\right)^{2},$
$b(l)$ $\triangleq\frac{l^{3}}{4(1+l)^{3}}\left(\frac{I(T)}{4L}\right)^{2},$
$B(l)$ $\triangleq 2+\frac{2(1+l)^{2}}{l^{2}}\left(\frac{4L}{I(T)}\right)^{2},$
$I(T)$ $\triangleq\epsilon(T)\min_{\theta\in\left[M\right]}\min_{m\neq\theta}\tilde{A}_{\theta,m},$
$\epsilon(T)$ $\triangleq\left(\log T\right)^{-\frac{1}{4}}.$
It suffices to show that $\frac{1}{b(l)}\log\left(MB(l)\left(1+b(l)\right)\right)\in o(T)$.
Note the following:
  1. 1.
    $\tilde{A}_{\theta,m}$, $A_{\theta,m}$ are finite for any two different hypotheses $\theta,m\in\left[M\right]$ due to Assumption 1.
  2. 2.
    $M$, $L$ are also finite.
By plugging the definitions of the respective terms, including $l$, $b(l)$, $B(l)$, $I(T)$, and $\epsilon(T)$, straightforward calculations lead to
$\frac{1}{b(l)}\log\left(MB(l)\left(1+b(l)\right)\right)\in O\left(T^{\frac{5}{6}}\left(\log(T)\right)^{\frac{1}{2}}\right)\subseteq o(T),$
and the proof is complete. ∎
-B Proof of Lemma 3
Proof:
We first derive the explicit form of $\mathsf{P}_{m}^{(t)}$ for any $m\in\left[M\right]$ and any $t\geq 1$, demonstrating that $S_{t,m,\theta}$ can be expressed as the sum of log-likelihood ratios. Subsequently, we show the martingale property by leveraging the additivity of log-likelihood ratios.
Firstly, let us derive the explicit form of $S_{t,m,\theta}$. We start with the following expression: $\forall\,t\geq 1$, $\forall\,f_{t}$, $\forall\,m\in\left[M\right]$,
$\mathsf{P}_{m}^{(t)}\left(F_{t}=f_{t}\right)$
$=\mathbb{P}_{m}\left(\mathbf{X}_{1}=\mathbf{x}_{1},Z_{1}=z_{1},A_{1}=a_{1}\right)\prod_{s=2}^{t}\mathbb{P}_{m}\left(\mathbf{X}_{s}=\mathbf{x}_{s},Z_{s}=z_{s},A_{s}=a_{s}\,\middle|\,F_{s-1}=f_{s-1}\right)$
$=\mathbb{P}_{m}\left(\mathbf{X}_{1}=\mathbf{x}_{1},Z_{1}=z_{1},A_{1}=a_{1}\right)\prod_{s=2}^{t}P_{Z}\left(z_{s}\right)\mathbb{P}_{m}\left(\mathbf{X}_{s}=\mathbf{x}_{s},A_{s}=a_{s}\,\middle|\,F_{s-1}=f_{s-1},Z_{s}=z_{s}\right)$ (42)
$=\mathbb{P}_{m}\left(\mathbf{X}_{1}=\mathbf{x}_{1},Z_{1}=z_{1},A_{1}=a_{1}\right)$
$\quad\cdot\prod_{s=2}^{t}P_{Z}\left(z_{s}\right)P_{A_{s}|Z_{s},F_{s-1}}\left(a_{s}\,\middle|\,z_{s},f_{s-1}\right)\mathbb{P}_{m}\left(\mathbf{X}_{s}=\mathbf{x}_{s}\,\middle|\,F_{s-1}=f_{s-1},Z_{s}=z_{s},A_{s}=a_{s}\right)$
$=p_{\text{init}}\prod_{s=2}^{t}P_{Z}\left(z_{s}\right)P_{A_{s}|Z_{s},F_{s-1}}\left(a_{s}\,\middle|\,z_{s},f_{s-1}\right)\prod_{s=1}^{t}P^{a_{s},z_{s}}_{m}\left(\mathbf{x}_{s}\right)$ (43)
where $p_{\text{init}}\triangleq P_{Z}(z_{1})P_{A_{1}|Z_{1}}(a_{1}\,|\,z_{1})$, $\mathbf{x}_{t}\triangleq\{x_{t,j}\,|\,j\in a_{t}\cap z_{t}\}$, and by convention, $\mathsf{P}_{m}^{(0)}\left(f_{0}\right)\triangleq 0$. Note that (42) holds true since $Z_{t}$ and $F_{t-1}$ are independent $\forall\,t\geq 1$, and (43) is due to the conditional independence of samples given the current actions across time slots and the distribution of $\mathbf{X}_{s}$ is $P^{a_{s},z_{s}}_{m}$ given $A_{s}=a_{s},Z_{s}=z_{s}$ under hypothesis $m$, $\forall\,s\geq 1$.
Hence, we have the explicit expression of $S_{t,m,\theta}$: $\forall\,m,\theta\in\left[M\right]$, $m\neq\theta$, $\forall\,t\geq 1$,
$S_{t,m,\theta}=\log{\frac{\mathsf{P}_{m}^{(t)}(F_{t})}{\mathsf{P}_{\theta}^{(t)}(F_{t})}}=\sum_{s=1}^{t}\log{\frac{P^{A_{s},Z_{s}}_{m}(\mathbf{X}_{s})}{P^{A_{s},Z_{s}}_{\theta}(\mathbf{X}_{s})}}$
where $S_{0,m,\theta}\triangleq 0$ by convention.
Then, we are ready to show the martingale property. To begin with, $\forall\,t\geq 0$, by the expression of $S_{t,m,\theta}$, we have
$\mathbb{E}_{\theta}\left[e^{S_{t+1,m,\theta}}\,\middle|\,\mathcal{F}_{t}\right]=e^{S_{t,m,\theta}}\mathbb{E}_{\theta}\left[e^{\log{\frac{P^{A_{t+1},Z_{t+1}}_{m}(\mathbf{X}_{t+1})}{P^{A_{t+1},Z_{t+1}}_{\theta}(\mathbf{X}_{t+1})}}}\,\middle|\,\mathcal{F}_{t}\right]$
where $\left\{e^{S_{t,m,\theta}}\,\middle|\,t\geq 0\right\}$ is adapted to the natural filtration $\left\{\mathcal{F}_{t}\right\}_{t\geq 0}$. The goal is turned to show for any $t\geq 0$,
$\mathbb{E}_{\theta}\left[e^{\log{\frac{P^{A_{t+1},Z_{t+1}}_{m}(\mathbf{X}_{t+1})}{P^{A_{t+1},Z_{t+1}}_{\theta}(\mathbf{X}_{t+1})}}}\,\middle|\,\mathcal{F}_{t}\right]=1.$
Since the distribution of action taking at time slot $t+1$ is a function of $F_{t}$ and $Z_{t+1}$ (see Definition 1), we simply switch from conditioning on $\mathcal{F}_{t}$ to merely conditioning on $F_{t}$ without loss of generality in the following steps. So, we have, $\forall\,t\geq 0$,
$\mathbb{E}_{\theta}\left[e^{\log{\frac{P^{A_{t+1},Z_{t+1}}_{m}(\mathbf{X}_{t+1})}{P^{A_{t+1},Z_{t+1}}_{\theta}(\mathbf{X}_{t+1})}}}\,\middle|\,\mathcal{F}_{t}\right]$
$=\sum_{a\in\mathcal{A}}\sum_{z\in\mathcal{Z}}P_{A_{t+1},Z_{t+1}|F_{t}}\left(a,z\,\middle|\,f_{t}\right)\mathbb{E}_{\theta}\left[e^{\log{\frac{P^{a,z}_{m}(\mathbf{X}_{t+1})}{P^{a,z}_{\theta}(\mathbf{X}_{t+1})}}}\,\middle|\,A_{t+1}=a,Z_{t+1}=z,F_{t}=f_{t}\right]$
$=\sum_{a\in\mathcal{A}}\sum_{z\in\mathcal{Z}}P_{A_{t+1},Z_{t+1}|F_{t}}\left(a,z\,\middle|\,f_{t}\right)\mathbb{E}_{\theta}\left[\frac{P^{a,z}_{m}(\mathbf{X}_{t+1})}{P^{a,z}_{\theta}(\mathbf{X}_{t+1})}\,\middle|\,A_{t+1}=a,Z_{t+1}=z,F_{t}=f_{t}\right].$
Subsequently, since the random sample set $\mathbf{X}_{t+1}$ follows the marginal distribution $P^{a,z}_{\theta}$ under a given $A_{t+1}=a$ and $Z_{z+1}=z$, and the samples are conditionally independent given the current action across time slots, we have
$\mathbb{E}_{\theta}\left[\frac{P^{a,z}_{m}(\mathbf{X}_{t+1})}{P^{a,z}_{\theta}(\mathbf{X}_{t+1})}\,\middle|\,A_{t+1}=a,Z_{t+1}=z,F_{t}=f_{t}\right]=1$
by Assumption 1.
To complete the proof, we can conclude, $\forall\,t\geq 0$,
$\mathbb{E}_{\theta}\left[e^{\log{\frac{P^{A_{t+1},Z_{t+1}}_{m}(\mathbf{X}_{t+1})}{P^{A_{t+1},Z_{t+1}}_{\theta}(\mathbf{X}_{t+1})}}}\,\middle|\,\mathcal{F}_{t}\right]$ $=\sum_{a\in\mathcal{A}}\sum_{z\in\mathcal{Z}}P_{A_{t+1},Z_{t+1}|F_{t}}\left(a,z\,\middle|\,f_{t}\right)$
$=\sum_{z\in\mathcal{Z}}\alpha_{z}\sum_{a\in\mathcal{A}}P_{A_{t+1}|Z_{t+1},F_{t}}\left(a\,\middle|\,z,f_{t}\right)$
$=1$
where the last equality holds by Definition 1 and the stochastic model of temporally available data sources in Section II-A. ∎